A Complete Bibliography of Applied Statistics: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 March 2017 Version 4.05

Title word cross-reference

(x′x)^{−1} [367]. 2 × 2 [159, 376]. 2 × 2 × K [667]. 2 × J × K [667]. ^{(R)} [656]. χ^{2} [112, 412, 416, 417, 713]. F [635, 663, 456, 619, 129, 285]. G [635, 663, 456]. I [604]. K [194, 30]. L_{1} [562, 686]. m [370]. M/E_{T} /m [606]. n [211]. p [278]. {P (l, k)} [134]. R × C [505, 135, 413, 660]. S_{B} [435]. T [436, 107, 263, 769, 437, 2, 619, 27, 8, 709, 203, 31, 542, 10]. U [342]. W [225, 239, 311, 299, 529]. x [278]. X^{2} [607].

*Jr [182, 381].

-Distribution [437, 2]. -Means [194, 30].

-Projections [604]. -Related [129]. -sphere [370]. -Test [203]. -The [321].

1.1 [632]. 100 [13, 138]. 106 [417, 14]. 108 [26, 562, 15]. 112 [17, 504, 139]. 122 [18, 191]. 125 [193, 20]. 126 [21, 220]. 127 [22, 242]. 128 [23, 636]. 129 [24, 66, 140]. 131 [770, 25]. 135 [29, 338]. 136 [194, 30]. 139 [32, 90, 137]. 141 [195, 33]. 146 [62]. 147 [741, 94, 63, 484]. 148 [64, 507]. 149 [85]. 150 [86]. 151 [87]. 152 [268, 88, 712]. 153 [89, 160, 416, 417]. 154 [111]. 155 [112, 417]. 156 [113]. 157 [133]. 158 [134]. 159 [135]. 16 [564, 5, 16]. 160 [136]. 161 [159]. 162 [361, 161]. 163 [162]. 164 [690, 163, 196]. 165 [185]. 166 [186]. 167 [187]. 168 [188]. 169 [418, 270, 189]. 170 [190]. 171 [214]. 172 [215]. 173 [216]. 174 [217]. 175 [218]. 176 [360, 530, 219]. 177 [310, 235]. 178 [236, 740]. 179 [237]. 180 [238]. 181 [239, 311, 529]. 182 [240]. 183 [459, 241, 362, 506]. 184 [263]. 185 [264]. 187 [266]. 188 [306]. 189 [307]. 190 [308, 440, 563]. 191 [309]. 192 [334]. 193 [438, 335]. 194 [336]. 195 [356, 439]. 196 [357]. 197 [358]. 1979 [158]. 198 [359]. 199 [544, 386]. 2 [656]. 2.0 [735]. 2.6 [707]. 200 [387]. 201 [388]. 202 [389]. 203 [411]. 204 [412]. 205 [505, 413]. 206 [414, 543, 565]. 207 [415]. 208 [433]. 209 [434]. 210 [435]. 211 [635, 663, 456]. 212 [457]. 213 [458]. 214 [480]. 215 [481, 711]. 216 [482]. 217 [483]. 218 [503]. 219 [527]. 220 [528]. 221 [541]. 222 [662, 560]. 223 [561]. 224 [581]. 225 [582]. 226 [774, 775, 583]. 227 [584]. 228 [604]. 229 [605]. 230 [606]. 231 [607, 713]. 232 [633]. 233 [634]. 234 [658]. 235 [659]. 236 [660]. 237 [661]. 238 [686]. 239 [687]. 240 [688]. 241 [689]. 242 [708]. 243 [709]. 244 [710]. 245 [737]. 246 [738]. 247 [739]. 248 [765]. 249 [766]. 250 [767]. 251 [768]. 2nd [181, 154, 520, 624, 180, 455, 287, 109, 131, 682]. 3 [706]. 32 [7, 741, 484]. 3rd [522, 61]. 4 [1, 8]. 5 [437, 2, 8]. 5th [354]. 6 [6, 28, 11, 267, 3]. 60 [504]. 61 [506]. 62 [507]. 63 [529]. 64 [530]. 65 [542]. 66 [543]. 67 [544]. 7 [6, 267, 4]. 73 [9, 192]. 75 [269]. 76 [769, 436, 27, 31, 542, 10]. 99 [12, 138]. A. [754, 58, 734, 679, 432, 56, 526, 157, 353, 755, 756, 402, 327, 59]. Absolute [26, 562, 15]. Abundance [52]. acceptance [422]. Accommodation [450]. Accounting [670]. Accurate [107]. Action [614, 654]. Adaptive [644]. Addendum [557, 558]. Additional [701]. Additive [611, 349]. Adjusted [666]. Adjustment [683, 153]. Adjustments [346, 673]. Adjusts [676]. against [278]. Agricultural [73, 384]. Aitchison [601]. Akaike [642]. Alan [476]. Alcohol [540]. Algebra [385]. Algorithm [162, 336, 635, 663, 448, 544, 418, 337, 194, 456, 265, 111, 270, 30, 438, 335, 737, 305, 309, 358, 285, 634, 386, 189, 185, 187, 599, 411, 26, 29, 738, 433, 435, 17, 387, 660, 21, 28, 237, 7, 18, 263, 528, 769, 414, 480, 24, 66, 140, 86, 87, 436, 437, 162, 236, 336, 635, 663, 1, 2, 85, 134, 112, 23, 334, 765, 581, 541, 527, 505, 770, 768, 662, 604, 503, 11, 89, 160, 412, 607, 686, 713, 139, 456, 774, 9, 214, 111, 710, 133, 338, 415, 560, 30, 483, 240]. Algorithm [4, 3, 238, 22, 62, 12, 13, 27, 438, 564, 306, 335, 481, 688, 113, 360, 33, 605, 658, 357, 20, 775, 94, 63, 218, 25, 561, 583, 709, 766, 264, 458, 88, 308, 440, 136, 216, 388, 737, 711, 661, 584, 64, 159, 309, 459, 358, 266, 739, 565, 15, 190, 215, 767, 633, 135, 507, 359, 634, 740, 386, 235, 239, 311, 434, 563, 659, 690, 413, 356, 439, 530, 217, 687, 636, 708, 712, 14, 219, 161, 307, 606, 163, 188, 189, 196, 457, 482, 185, 5, 16, 389, 31, 542, 241, 362, 689, 582, 32]. Algorithm [90, 137, 10, 506, 186, 187, 220]. Algorithms [411, 6, 197, 198, 26, 29, 738, 433, 435, 17, 387, 660, 21, 28, 237, 7, 18, 191, 263, 528, 769, 414, 480, 195, 24, 66, 140, 86, 87, 436, 437, 162, 236, 336, 635, 663, 544, 1, 2, 85, 134, 741, 112, 23, 334, 765, 581, 541, 418, 527, 505, 504, 770, 337, 768, 662, 604, 194, 503, 11, 89, 160, 412, 416, 417, 607, 562, 686, 713, 139, 456, 265, 267, 774, 9, 192, 193, 214, 111, 710, 269, 133, 338, 415, 560, 270, 30, 483, 240, 4, 3, 238, 22, 62, 12]. Algorithms [13, 27, 138, 438, 564, 8, 306, 335, 481, 688, 113, 360, 33, 605, 310, 658, 357, 361, 20, 775, 63, 218, 25, 561, 583, 709, 766, 268, 264, 458, 88, 308, 440, 136, 216, 388, 737, 711, 661, 584, 64, 159, 309, 459, 358, 266, 739, 543, 565, 15, 190, 215, 767, 633, 135, 507, 359, 484, 634, 740, 386, 235, 239, 311, 434, 529, 559, 563, 659, 690, 413, 356, 439, 530, 217, 687, 636, 708, 712, 14, 219, 161, 307, 606, 163, 188, 189, 196, 457, 482, 185, 5, 16, 242, 389, 31, 542, 241, 362]. Algorithms [689, 582, 32, 90, 137, 10, 506, 186, 187, 220, 129, 502, 501]. Allocation [466]. Alternative [134, 78, 297]. Alternatives [173]. Alvan [625]. Alvey [326]. Amalgamation [85]. America [521]. analyse [117]. Analyses [587, 153, 679]. Analysing [668, 523, 158]. Analysis [91, 716, 699, 611, 684, 738, 396, 499, 734, 204, 474, 276, 725, 375, 316, 355, 676, 445, 154, 126, 290, 373, 592, 652, 500, 471, 223, 357, 361, 371, 381, 34, 612, 280, 536, 161, 352, 601, 246, 297, 651, 555, 720, 717, 97, 36, 202, 556, 423, 367, 123, 462, 147, 150, 77, 117, 754, 55, 292, 734, 629, 155, 752, 353, 333, 678, 332]. Analysis-Methods [474]. Analytic [320]. Analyzing [55]. Anderson [478, 158, 410]. Andrew [51]. Angular [274, 467]. animal [374, 171]. anisotropic [423]. Annotated [81]. Ant [322]. Application [492, 471, 643, 597, 228, 424, 648, 121, 104, 590, 669]. Applications [382, 383, 731, 644, 247, 333, 577, 555, 367, 172, 700, 474, 728]. Applied [253, 259, 317, 753, 61, 183, 500, 354, 428, 502, 478, 279, 455, 332]. Approach [110, 128, 297, 717, 422, 127, 401]. approaches [746]. appropriate [175]. Approximate [396, 334, 377, 310, 309, 235]. Approximating [387, 23, 658, 697, 636]. Approximation [619, 606, 173, 177]. Approximations [224]. Architecture [625]. area [175]. ARIMA [240]. ARMA [336, 309]. Armigera [146]. Arnold [519, 500]. Arsenic [597]. Ascher [498]. Aspects [375]. Aspiration [409]. ASR [544, 504, 543, 507, 529, 530, 542, 506].

Assay [443]. Assessing [695, 490, 469, 299]. Assessment [671, 394, 723, 255]. Assisted [524]. Association [495, 136, 295]. Assumption [427]. Assumptions [317]. Asymptotic [256, 570]. Athletic [620]. attained [254]. Auditing [449]. Authors [314]. Authorship [524]. AUTOBOX [735]. AUTOBOX-PLUS [735]. Autocorrelated [617]. Autocorrelations [398]. Automatic [264, 735]. autoregression [169]. Autoregressive [738, 301, 111, 661, 358, 708, 122, 569, 578, 669]. Autoregressive-Moving [111, 358, 569]. Auxiliary [1, 8, 482]. Average [111, 464, 661, 358, 708, 511, 569, 578].

B [260, 623, 522, 454, 261, 626, 83, 109, 131]. B. [654, 519, 629, 291, 352]. Back [67, 95, 114, 141, 164, 199, 221, 243, 271, 293, 312, 339, 363, 390, 419, 441, 460, 485, 508, 531, 545, 566, 585, 608, 637, 664, 692, 714, 742, 771, 162]. Baker [50]. Balakrishnan [630]. Balance [683]. Balanced [527, 216]. Bands [209]. Bank [248]. Banks [650]. Bargaining [409]. Barry [262]. Bartholomew [752, 51]. Based [348, 761, 100, 299, 389, 618, 538, 172, 150]. Basford [731]. Basic [522, 269, 327, 681, 57]. Basics [333]. Bayesian [652, 499, 496, 147, 449, 510, 77, 598, 246]. Beat [454]. Beet [536]. Before [319]. Behavioral [157, 156]. Bell [648]. Belson [622]. Berenson [753]. Bergstrom [380]. Bernard [651]. best [254]. best-fitting [254]. Beta [396, 774, 775, 583, 307]. Beta-Binomial [396]. Between [611, 350, 495, 388, 295, 675]. Bias [666, 620]. Bibby [155]. BIBD [223]. Bibliography [81]. Bimodal [208]. Binary [445, 576, 79, 584, 205, 98, 284]. Bingham [570]. Binomial [396, 19, 24, 66, 140, 643, 549, 307, 229, 65, 143]. Biochemistry [540].

Biological [52, 210, 148]. Biomass [670]. Biostatistics [291, 682, 653]. Biothermal [146]. Biplot [428]. Bird [228]. Births [315]. Bivariate [769, 87, 436, 322, 27, 31, 542, 10]. Blanchflower [651]. Block [755, 429, 424, 284]. Blocking [581]. Book [677, 55, 50, 754, 727, 260, 684, 382, 329, 473, 289, 680, 52, 58, 654, 623, 499, 132, 326, 181, 522, 292, 404, 734, 179, 521, 262, 60, 259, 679, 759, 385, 383, 650, 474, 432, 331, 477, 431, 524, 603, 54, 154, 631, 126, 127, 728, 290, 753, 61, 519, 629, 56, 155, 731, 733, 453, 525, 291, 454, 652, 130, 182, 261, 622, 84, 288, 183, 430, 757, 655, 475, 500, 409, 328, 451, 752, 526, 110, 128, 354, 380, 627, 129, 157, 452, 626, 381, 407, 602, 406, 156, 625, 683, 401, 653, 330, 761, 520]. Book [681, 403, 758, 353, 624, 405, 502, 478, 180, 630, 755, 756, 408, 455, 729, 82, 158, 57, 287, 333, 402, 83, 523, 327, 59, 51, 352, 109, 131, 81, 628, 732, 601, 501, 384, 410, 651, 682, 498, 476, 678, 760, 53, 730, 332, 656, 262]. Bootstrap [648, 555]. Bootstrapping [593]. Both [323]. Bound [544, 634, 386, 356, 439]. Bounds [254, 449, 511]. Bowerman [627]. Bowman [727]. Box [153, 579]. Bozdogan [754]. Branch [544, 634, 386]. Breast [373]. Bridge [554]. Brillinger [451]. Britain [651]. Brite [179]. Broemeling [499]. Brook [500]. Brown [291]. Bruce [627]. Burden [288]. Burnham [410]. Business [179, 655, 627, 455]. Buying [245].

C [677, 181, 292, 759, 126, 519, 183, 500, 526, 653, 681, 333, 402, 83, 59, 384, 410, 476, 760, 730]. C. [154, 183, 128]. Calculated [398]. Calculating [769, 436, 619, 342, 27, 645, 31, 542, 517, 10]. Calculation [480, 134, 309]. Calculations [261]. calculator [697]. Calibration [283, 71, 535, 516, 590]. Campbell [626]. Cancer [373, 430, 381, 295, 428]. Canonical [91, 36, 202]. Capture [410]. Capture-Recapture [410]. carcinogenic [568]. Carcinogenicity [427]. Carlo [480, 249, 226]. CARP [763, 764]. Carter [381]. Case [85, 366, 490, 357, 361, 696, 59, 161, 103, 721, 76, 575]. Case-Control [357, 361, 161, 103, 76]. Case-Weighted [696]. Casebook [291]. Cases [345, 574]. Categorical [55, 69, 185, 120, 166]. Catherine [290]. Cause [699, 640]. Cause-of-Death [699, 640]. Causes [498]. Cells [543]. Censored [503, 193, 564, 20, 628, 5, 16, 32, 90, 137, 74]. Census [228]. Central [263, 2, 416, 774, 27, 8, 775, 709, 190, 285, 31, 542, 10]. Centrality [190]. Chain [277]. Chain-Deferred [277]. Challenge [444]. Change [167, 523, 174, 147, 77]. Change-Measurement [523]. change-point [77]. changing [669]. Channel [158]. Characteristics [528]. Chart [369, 758, 100, 233, 392]. Charts [368, 518, 511]. Chatfield [154]. Chemical [452]. Chemotherapy [381]. Chi [190, 687, 745]. Chi-square [745]. Chi-Squared [190, 687]. Child [395]. Children [514, 490, 720]. Choice [488, 318, 719]. choosing [569]. Chris [110]. Circle [274, 467]. Circular [93, 42, 324]. Clarke [180, 327, 730]. Classes [761]. Classification [746]. Cleveland [602]. Clinical [625]. Clinimetrics [756]. Closed [667, 619, 410]. Clustering [194, 30, 746, 731]. Clusters [706]. Code [770, 25, 584]. Coefficient [44, 246]. coefficients [177]. Cohort [596]. Collapsibility [710, 667]. Collecting [81]. collinearities [702]. Collinearity [236, 740]. Column [113, 135, 705]. Combination [112, 412, 416, 417]. Combined [388, 717, 103, 392]. Combining [40, 581, 113]. come [748]. Comment [741]. Common [44, 228, 425, 101]. Communities [52]. Comparative [156]. Comparing [393, 19, 24, 66, 140, 251, 281, 591, 65, 748]. Comparison [395, 465, 671, 516, 320, 577, 555, 642, 118, 598]. Comparisons [227]. competing [462]. Competition [620]. complex [74]. Component [581, 113, 371]. Components [514, 319, 257, 704, 551, 593]. Composite [145]. Compositional [446, 601]. Compounds [614]. Computation [483, 190, 633, 247]. Computational [375]. Computed [395]. Computer [524, 261, 627, 320, 109, 131, 370]. Computer-Assisted [524]. Computerized [166]. Computing [774, 605, 775, 583, 285, 333, 618, 182, 327]. concentration [232]. Concepts [761]. Concerning [44]. Concordance [300]. concurrence [284]. Conditional [357, 361, 161, 143]. Conditions [388]. Conference [158]. Confidence [480, 209, 618, 538, 256]. Configurations [739]. Confined [32, 90, 137]. Confounder [255]. Confounding [517]. Conover [287]. consistent [667]. constancy [72]. constant [41]. Constrained [562, 582]. Constraints [604, 690, 163, 196]. Construct [185]. Construction [62, 559, 705, 422, 166, 494]. Consultant [654]. Consumption [540]. contamination [748]. contiguous [705]. Contingency [48, 264, 136, 159, 739, 206, 376, 717]. Continuity [733]. Continuous [733, 205, 144]. Contract [554]. Contrasts [186]. Contributions [408]. Control [368, 518, 357, 361, 100, 226, 161, 103, 392, 76]. Controlled [302]. controls [76]. Convenient [305]. Cooke [327]. Cooper [353]. Cope [404]. Cornell [679, 430]. Corner [661]. Cornish [658]. Cornish-Fisher [658]. Correction [48, 65, 66, 210, 505, 355, 160, 713, 269, 565, 311, 196, 362]. Corrections [440, 439]. Correlated [646]. Correlation [301, 44, 768, 458, 668, 633]. Correlations [395]. Correspondence [717]. Corrigenda [6, 91, 92, 93, 94]. Corrigendum [518, 691].

Corry [651]. Counting [86, 87, 584].

Course [253, 321, 453, 753]. courses [590]. Covariance [346, 23, 766, 613, 636, 97, 150, 117]. Covariates [193, 596, 20, 534, 539, 673]. Coverage [512]. Cox [448, 579, 102]. Cramér [218]. Cramér-Wold [218]. Craven [327]. Crayton [476]. Criminal [345, 574]. criteria [642]. Critical [159]. Critique [176]. Crop [730]. Cross [9, 192, 576, 79, 152, 118]. Cross-Over [576, 118]. Cross-Spectrum [9, 192]. Cross-Validatory [79]. Crossover [445]. Crow [728]. Cubic [443]. Cuming [473]. Cumulative [709, 268, 88, 100, 712, 224, 173, 177, 392]. Curve [492, 276, 457, 38]. Curves [435, 276, 334, 167, 12, 138, 209]. Customers [606]. Cusum [369, 233]. Cusums [368, 518]. Cyert [652].

D [499, 259, 183, 407, 653, 761, 502, 158]. D. [260, 654, 292, 404, 753, 733, 525, 328, 451, 752, 405, 763, 764, 327, 732, 410, 53, 332]. D’Agostino [623]. Daily [315]. Damped [210, 148]. Daniel [81]. Data [446, 55, 754, 716, 699, 684, 303, 467, 396, 734, 87, 251, 162, 93, 42, 210, 447, 316, 355, 395, 503, 265, 193, 290, 596, 169, 515, 587, 629, 703, 564, 643, 552, 69, 576, 464, 551, 20, 640, 381, 641, 613, 491, 443, 353, 279, 372, 34, 594, 333, 402, 81, 304, 185, 5, 16, 246, 389, 205, 324, 32, 90, 137, 74, 572, 152, 98, 148, 101, 462, 534, 104, 150, 77, 117, 166, 748, 102, 651, 602, 523, 352, 601]. Data-Based [389]. David [333, 51, 651]. Davies [452]. Davis [55]. Death [699, 640, 295]. Debugging [612]. Decision [127, 409, 678]. decisive [422]. Decomposability [710]. Decomposition [28, 11, 267, 3]. decontaminants [748]. Deferred [277]. Defined [704]. Definite [267, 4]. Definition [212]. Degree [464]. Degrees [607, 713]. DeGroot [624, 652]. deletions [575]. Demand [176]. Demonstrations [761]. Density [446, 629, 360, 530, 224, 219, 494]. departures [647]. Dependence [695, 306, 171]. Dependent [596, 534, 539, 144]. Depths [597]. Derivatives [266]. Derived [17, 504, 139]. Description [559]. Design [666, 431, 581, 126, 288, 113, 216, 180, 536, 670, 646, 284, 500, 59]. Design-Adjusted [666]. designed [117]. Designs [679, 581, 211, 113, 465, 424, 186, 187, 705, 478, 755]. Desirability [346]. Destructive [350]. Detecting [178, 572]. Detection [236, 450, 740, 534]. Determining [544, 386]. Development [146]. deviance [575]. Deviation [238]. Device [298]. Dey [755]. Diagnosis [79]. Diagnostic [614]. Diagnostics [744, 575]. Diagonalization [635, 663, 456]. Diapausing [146]. Dice [750]. Dichotomous [668]. Dickinson [520]. Differences [274, 388, 675]. Different [350]. Differential [677, 79]. Dilon [474]. Dimensional [280]. Dip [483]. Direct [216, 215]. Directed [513]. directional [423]. Directory [526]. Discordancy [323, 168]. Discovering [331]. Discrete [54, 265, 320, 101]. Discriminant [91, 185, 36, 556, 166]. Discriminating [611]. Discrimination [615, 318, 205, 120]. disease [103]. Dispersion [399, 688, 767]. Display [429]. Distance [279]. Distances [582]. distributed [170]. Distribution [433, 492, 40, 437, 1, 2, 146, 112, 765, 541, 503, 412, 416, 417, 607, 713, 214, 587, 43, 8, 481, 570, 450, 597, 709, 766, 711, 491, 190, 285, 434, 14, 224, 245, 307, 606, 511, 689, 208, 173, 168, 232, 149, 177, 150, 598, 727]. Distribution-Free [40, 150]. Distributions [411, 17, 19, 24, 66, 140, 319, 296, 504, 770, 139, 728, 249, 619, 342, 129, 626, 25, 74, 101, 722, 247, 370, 65, 477, 519, 733]. Distributions-Regression [626]. Diversity [52]. Divestiture [648]. DNA [352]. DO [215]. DO-LOOPS [215]. Domain [451]. Donald [454, 381]. Dose [587, 209, 38, 748]. Dose-Response [587, 209, 748]. Douglas [181]. Down [133]. Drawing [305, 599]. Drinking [295]. dually [667]. Dunn [701]. Duration [703]. Durbin [89, 160, 416]. Durbin-Watson [89, 160, 416]. During [282]. Dyke [730].

E. [292, 728, 109, 131]. Early [725, 288]. Eaton [401]. Econometrics [382]. Economic [521, 652, 409, 646, 648, 719]. Economics [655]. ed [354]. Edition [656, 181, 522, 154, 61, 520, 180, 455, 287, 109, 131]. Editorial [314]. Editors [580, 213, 325, 621, 108, 726, 600, 479, 379, 125, 400, 258, 80, 286, 234, 49, 649, 184, 351]. Edn [682, 624]. EDP [276]. Educated [404]. Educational [757]. Edwards [262]. Effect [315, 103, 118]. Effects [70, 255, 534]. efficiency [645]. Efficient [342, 170, 584, 459, 135, 255, 241, 362, 506, 247, 599]. Efficiently [512]. Efron [528, 291]. Efron-Morris [528]. Ehrenberg [402]. Eigenvalues [458]. Elementary [525]. Elements [503, 602, 349]. Elliott [526]. Elston [653]. Emergence [146]. Emphasis [52]. Empirical [765]. Employment [651]. Encyclopedia [329, 626, 758]. Engen [52]. Engineering [181]. Engineers [56]. Ensemble [476]. Entropy [733]. Enumeration [660, 237, 505, 413]. Epidemiologic [468]. Epidemiology [625, 83, 555]. Equal [211, 378]. Equality [767, 143]. Equating [328]. Equations [469, 677]. Error [356, 439]. Errors [26, 562, 15, 670, 722, 571]. Essays [408]. Essentials [653]. Establishment [651]. Estimate [86, 579, 747]. Estimated [514]. Estimates [29, 40, 87, 338, 372, 512, 745]. Estimating [350, 274, 568, 670, 398, 38, 124]. Estimation [411, 446, 397, 472, 369, 541, 366, 562, 193, 111, 629, 43, 548, 733, 564, 306, 481, 718, 496, 360, 380, 20, 613, 711, 491, 681, 228, 121, 226, 530, 219, 307, 5, 16, 389, 489, 32, 90, 137, 152, 122, 97, 232, 374, 645, 721]. Estimator [238]. Estimators [91, 727, 301, 528, 279, 470, 425, 36, 598]. EV [764]. Evaluating [367]. Evaluation [660, 554, 574, 766, 426, 672]. Event [54]. Everitt [654]. Evidence [345, 574]. Exact [17, 660, 19, 24, 66, 140, 504, 139, 214, 111, 310, 358, 206, 235, 708, 65]. Examination [464, 384]. Examine [295]. Example [540, 533, 73]. Examples [230, 231, 290]. exciting [169]. Exercise [302]. Expansions [570, 658, 426]. expectations [105]. Expected [310, 235]. Experience [492, 683]. Experiences [510, 82]. Experiment [227, 384, 568]. Experimental [431, 500, 180, 187]. Experimentation [730]. Experiments [699, 640, 465, 216, 536, 577, 429, 99, 117, 748, 124, 126]. Explain [680]. Explanation [523]. explanatory [175]. Exploitable [670]. Exploiting [733]. Exploratory [734, 204, 69, 333]. Exploring [515]. Exponential [193, 252, 20, 491, 457, 74, 149]. Exponentially [210, 148]. Extension [225]. Extra [347, 549, 229]. Extra-Binomial [549, 229]. Extra-Poisson [347]. Extraneous [549]. Extreme [503, 365, 481, 711, 74]. Extreme-Value [365, 481, 711].

F [404, 759, 383, 333, 51, 760]. F. [684, 61, 407, 83, 682]. Fabric [39]. Factor [734, 752, 73]. Factorial [465, 216, 478, 577, 186, 99]. Factorization [218]. Factors [581, 103]. Fadil [455]. Failure [296, 471]. Failure-Time [296]. familial [747]. Family [296, 395, 78, 626]. Farlow [501]. Fast [662, 265, 560, 360, 358, 530, 219]. Feature [544, 634, 386]. Feingold [498]. Feinstein [625, 756]. Fibre [574]. Fiebig [733]. Field [384]. fields [104]. Fildes [53]. filter [153].

Filtering [111]. Financial [398]. Finding [604]. Findings [288]. Finite [604, 9, 192, 240]. First [132, 753]. Fisher [503, 168, 232, 214, 450, 658]. Fishery [670]. Fishman [54]. Fit [623, 336, 765, 490, 723, 105, 463]. Fitting [433, 435, 553, 296, 686, 101, 415, 12, 138, 552, 45, 739, 539, 594, 457, 482, 304, 102, 74, 254]. Five [435, 321]. Fixed [237, 380]. Fixed-Point [380]. Fleming [759, 760]. Fluctuations [228]. Fluoride [295]. Follow [491, 255]. Follow-Up [491, 255]. footrule [212]. Forbes [51]. Force [470]. Fore [321]. Forecasting [47, 735, 82, 53, 722]. Form [176, 581, 619, 723, 667, 113]. Formatting [188]. Forms [14, 367]. Formulae [466]. FORTRAN [166, 215, 185]. Foundations [732]. Fourier [662, 265, 9, 192, 560, 360, 426, 530, 219]. Fox [475]. Fractional [211, 478]. France [521]. Frankel [288]. Free [40, 587, 150]. Freedom [607, 713]. French [289]. Frequency [770, 451, 589, 25]. Front [68, 96, 115, 142, 165, 200, 222, 244, 272, 294, 313, 340, 364, 391, 420, 442, 461, 486, 509, 532, 546, 567, 586, 609, 638, 665, 693, 715, 743, 772]. Fuller [763, 764]. Function [301, 19, 24, 66, 140, 1, 765, 43, 8, 709, 737, 434, 224, 185, 349, 173, 65, 121, 166]. Functional [176, 723]. Functions [207, 145]. Fundamentals [157].

G [733, 291, 430, 683, 732]. G.

[383, 731, 500, 763, 764, 82, 327, 410, 678, 730]. Galambos [603]. Galen [381]. Galwey [326]. Game [554]. Games [204]. Gamma [727, 7, 741, 323, 275, 94, 63, 737, 266, 484, 687]. Gard [677]. Gauss [236, 740]. Gaussian [399, 245, 489]. Gaussian-Poisson [245]. Gelfand [476]. General [468, 296, 249, 61, 415, 320, 304]. General-Purpose [320]. Generalization [203]. Generalized [92, 344, 342, 481, 471, 711, 37, 245, 555, 575, 145]. Generated [298]. Generating [135]. Generation [399, 725, 249, 22, 589, 458, 216, 584, 320, 242, 186, 170, 724, 370]. Generator [459, 241, 362, 506]. Genesis [524]. Genetic [431, 747]. Genstat [35, 326]. Gentle [182]. Geoffrey [180]. Geometric [511]. George [54]. Germany [521]. Gibbons [520]. Gilchrist [403]. Giri [126]. Given [399, 544, 135, 386, 256, 284]. Givens [162]. GLIM [74, 50, 448, 403, 102, 595, 296]. global [463]. GMDH [501]. Going [162]. Goldsmith [452]. Goldstein [474, 84, 757]. Golf [321]. Goodman [55]. Goodness [623, 92, 336, 765, 37, 105, 463]. Goodness-of-Fit [623, 765, 105, 463]. GPSS [130]. GPSS-Simulation [130]. graph [556]. Graphical [429, 572, 98, 534]. Graphing [602]. Gray [526, 584]. Great [651]. Greer [56]. Gregory [678]. Grid [45]. Griffiths [502]. Groeneveld [57]. Group [371, 744]. Grouped [564, 491, 5, 16]. Grouping [156]. Groups [281]. Growth [276, 167, 330]. Guernsey [158]. Guessing [404]. Guide [473, 290, 656]. guided [702]. Guidelines [314]. Gupta [754].

H [650, 331, 61, 525, 652, 381, 624, 455, 327]. H. [754, 733, 84, 757, 526, 156, 763, 764]. Haber [157]. Haematology [540]. Haemolytic [490]. Hakim [290]. Hand [554, 654]. Handbook [354, 684, 631]. Harold [498]. Harry [498]. Hastings [173]. Hausch [454]. Hazard [207, 389, 425]. Hazards [373, 696, 463]. Hebden [382]. Height [527]. Heiler [406]. Held [158]. Heliothis [146]. Helmut [628]. Hennes [259]. Herman [380]. Heterogeneity [595]. Heteroscedastic [424]. highest [494]. Hill [502]. Hines [181]. Hinkelmann [431]. Hitting [557, 493, 558]. Hoaglin [333]. Hogg [477]. Hokstad [606]. Holland [328]. Homogeneity [159, 647]. Honour [408]. Horse [365]. Human [330]. Hypergeometric [268, 88, 712].

I. [502]. Ian [523]. Identification

[723, 376, 578]. Identifying [661]. II [202]. Ill [704]. Ill-Defined [704]. IMF [683]. immunoassay [121]. Implants [597]. Improved [418, 270, 377, 634, 189]. Improvement [550]. Inc. [707, 735]. incidence [284]. Incomplete [7, 251, 741, 94, 63, 266, 484, 687, 186, 284]. Incorporated [736]. Incorporating [467, 70]. Independent [393, 414, 543, 565, 724, 147]. Index [197, 60]. Indexing [321]. Indices [615, 748]. Individual [317]. Individuals [466]. Indonesian [720]. Industry [510, 627, 452]. Inequality [604, 582]. Inference [492, 18, 191, 174, 731, 348, 520, 359, 498, 476, 667, 630]. Inferences [44]. Influence [696, 617]. Information [116, 201, 273, 341, 421, 487, 547, 610, 694, 773, 467, 503, 109, 131]. Infrared [283]. Insights [701]. Inspection [277, 719]. Instat [657]. Instrument [283]. Instruments [516]. Integral [21, 7, 769, 436, 437, 2, 741, 27, 8, 94, 63, 266, 484, 687, 224, 31, 542, 10, 220]. Integrals [1, 768, 8]. Integrated [681, 279]. intensities [721]. Interaction [73]. Intergroup [300]. International [158]. Interpret [640]. Interpretation [280, 349]. Interpreting [345]. Interval [43, 512, 256]. Intervals [480, 618, 538, 494]. Intraclass [668]. Introduction [677, 326, 292, 154, 110, 128, 57, 682]. Introductory [58, 603, 453, 402]. Inverse [301, 399, 688, 224, 245, 489, 173, 697]. Inversion [195, 267, 4, 33]. invertibility [569]. Investigating [69]. Investigation [317]. Ireland [695]. Islands [158]. Isotonic [414, 209, 543]. Israel [315]. Issues [648]. Istonic [565]. Italy [521]. Iterative [718, 550]. Iteratively [343].

J [50, 684, 654, 60, 331, 631, 127, 731, 182, 261, 500, 752, 129, 407, 353, 763, 764, 287, 51, 501, 682]. J. [55, 50, 679, 155, 182, 475, 156, 681]. Jackknife [64, 507]. James [528, 259]. Janos [603]. Japan [521]. Jean [520]. Jenkins [153, 82]. Joanne [288]. John [601]. Johnson [329, 435, 337, 12, 13, 138, 626, 653, 758, 332, 408]. Johnson-Normal [337, 13, 138]. Joint [62]. Jordan [236, 740]. Jr [291]. Judge [620]. Julia [382].

K. [727, 431, 728, 155, 731, 183, 129, 83, 410]. Kalman [111]. Kempson [730]. Kennedy [182, 763, 764]. Kent [155]. Kernal [360]. Kernel [446, 447, 530, 219]. Khuri [679]. Kickert [127]. Kimber [631]. Kinetics [432]. Kingdom [521, 759, 760]. Klugman [477]. Known [274]. Kotz [329, 404, 626, 758]. Krishnaiah [451]. Kurtosis [434, 46].

L [329, 179, 728, 753, 183, 526, 452, 626, 381, 401, 758, 478, 630, 410, 408]. L. [55, 727, 680, 291, 84, 109, 131]. Labour [470]. Lachenbruch [685]. Lacking [640]. Lamb [643]. Lambrecht [432]. Lamps [204]. Lane [326]. Large [269, 515, 69, 225, 476]. Large-Scale [476]. Latent [55, 752]. law [171]. Leach [110, 128]. Least [663, 690, 163, 196, 343, 457, 72]. Leech [83]. Length [511, 256]. Lethality [676, 640]. Letter [213, 726, 600, 479, 379, 80, 234]. Letters [580, 325, 621, 108, 125, 400, 258, 286, 49, 649, 184, 351]. Level [359, 41]. Levels [409, 38, 186, 187]. Leverage [617]. Levine [753]. Lewis [183]. Liar [750]. Liesner [521]. Life [248]. lifetimes [422]. Likelihood [411, 660, 472, 396, 541, 574, 193, 111, 43, 564, 481, 20, 711, 593, 309, 358, 708, 307, 304, 5, 16, 618, 349, 32, 90, 137, 746, 749, 598, 144]. Limits [260]. Line [686]. Linear [55, 26, 29, 369, 368, 518, 347, 499, 92, 292, 162, 344, 106, 112, 604, 412, 416, 417, 562, 710, 269, 338, 415, 238, 475, 471, 576, 658, 516, 15, 279, 37, 690, 163, 196, 343, 482, 555, 229, 582, 32, 90, 137, 717, 749, 145, 575, 681]. lines [77]. Linguistics [524]. Link [92, 344, 37, 145]. List [198]. Literature [300]. Location [318]. Log [55, 303, 347, 710, 415, 576, 717]. Log-Linear [55, 347, 710, 415, 576, 717]. Log-Logistic [303]. Logarithm [737]. logarithmically [170]. Logistic [433, 303, 588, 490, 357, 361, 744, 723, 443, 161, 38, 229]. Lognormal [728, 43]. Long [695]. Long-Memory [695]. Longitudinal [523]. Look [132]. LOOPS [215]. Loren [631]. Lorenz [492]. Loss [477, 224]. Lothar [354]. Lyle [499].

M [58, 259, 432, 127, 753, 155, 652, 130, 381, 180, 82, 327, 730]. M. [623, 759, 126, 753, 407, 330, 681, 624, 630, 760]. MacSpin [632]. Made [130]. Magidson [55]. Magnitude [211]. Mahalanobis [582]. Making [127, 409]. man [747]. Management [181, 631]. Manager [473]. Manipulation [109, 131]. Mann [342]. Manpower [51, 41]. Manual [261, 730, 50]. Many [252, 101]. marathon [590]. Mardia [155, 129]. Margaret [683]. Marginal [136, 46]. market [669]. Markov [647, 721]. Marks [464, 671]. MARMA [633]. Martin [758]. Matched [361, 161, 255, 76, 673]. Maths [262]. Matrices [22, 458, 767, 633, 242]. Matrix [28, 195, 385, 23, 503, 11, 267, 4, 3, 688, 33, 613, 216, 636, 367, 284]. Matter [67, 95, 114, 68, 96, 115, 141, 164, 199, 142, 165, 200, 221, 243, 271, 222, 244, 272, 293, 312, 339, 294, 313, 340, 363, 390, 419, 364, 391, 420, 441, 460, 485, 442, 461, 486, 508, 531, 545, 509, 532, 546, 566, 585, 608, 567, 586, 609, 637, 664, 692, 638, 665, 693, 714, 742, 771, 715, 743, 772]. Matthew [474]. Maunder [759, 760]. Max [29, 338]. Maximum [411, 263, 541, 193, 111, 733, 564, 481, 20, 711, 307, 304, 5, 16, 32, 90, 137, 746, 645, 598, 144]. Maximum-Likelihood [481, 711]. McCullagh [729]. McLachlan [731]. McLean [478]. Mean [399, 496, 766, 613, 171]. Means [393, 40, 85, 194, 111, 30]. Measurement [288, 469, 597, 523, 120]. Measurements [738, 591]. Measures [696, 747]. Measuring [60, 590]. Memory [695]. Method [433, 40, 447, 84, 591, 661, 378, 135, 618, 572, 374, 534, 247, 150, 144]. Methods [474, 290, 430, 452, 394, 516, 320, 510, 729, 280, 577, 501, 410, 99, 667, 473, 475, 156]. Mezzich [156]. Michael [473, 650]. Mick [706]. Microcomputer [526, 761]. Microcomputer-Based [761]. Migration [330]. Miller [291]. Mills [176]. Min [29, 338]. Min-Max [29, 338]. Minimal [342]. Minimal-Storage [342]. Minimizing [582]. Minimum [26, 562, 15]. Miscellanea [107, 152, 497, 176, 230, 557, 153, 211, 232, 256, 377, 257, 450, 579, 177, 305, 378, 233, 285, 751, 559, 46, 178, 324, 124]. Misconceptions [498]. Missing [344, 718, 613, 117, 124]. Misspecified [47]. Misuse [283]. Mixed [375, 643, 318, 491, 205, 117]. mixed-up [117]. Mixed-Variable [318]. Mixing [541, 366]. Mixture [472, 488, 319, 731, 593, 746, 462]. Mixtures [411]. Model [207, 639, 162, 448, 146, 375, 317, 538, 415, 365, 490, 469, 576, 318, 264, 443, 612, 708, 246, 349, 32, 90, 137, 174, 172, 642, 462, 463, 104, 700, 669, 102, 575]. Model-based [538]. Modeling [754, 432, 627, 498, 476]. Modelling [595, 596, 515, 695, 549, 82, 501]. Models [55, 48, 468, 472, 303, 467, 52, 347, 553, 92, 195, 336, 344, 106, 537, 431, 47, 70, 167, 111, 710, 373, 703, 731, 757, 475, 471, 643, 752, 33, 45, 661, 309, 358, 353, 739, 428, 279, 372, 37, 617, 691, 616, 343, 424, 482, 304, 555, 229, 120, 749, 569, 578, 539, 747, 145, 144, 499, 758]. Modification [233]. Modified [749].

Moments [433, 435, 12, 138]. Monetary [449]. Monetary-Unit [449]. Money [176]. Monte [480, 249, 226]. Montgomery [181, 292]. Moore [522]. Morris [528, 652, 401]. Mortality [643, 428, 720, 175]. Moses [680, 291]. Moss [84]. Moving [162, 111, 661, 358, 708, 511, 569, 578]. Multi [237, 124]. Multi-Sets [237]. multi-stratum [124]. Multidimensional [136, 739]. Multilevel [757]. Multinormal [766, 426]. multiparameter [745]. Multiple [26, 29, 263, 639, 562, 338, 703, 744, 15, 577, 572, 724]. Multiple- [263]. Multiple-Group [744]. Multiple-Spell [703]. Multiplicative [611]. Multipurpose [397, 538]. Multisample [217]. multistate [647]. Multivariable [191]. Multivariate [433, 18, 734, 319, 474, 768, 155, 357, 361, 551, 401, 394, 299, 356, 439, 217, 46, 161, 332, 572, 98, 97, 202, 151, 754]. Murray [261].

N [126, 762]. N. [326, 126, 758, 630]. Names [198]. Nash [681]. NBD [317]. Near [283]. Negative [14]. Neighbouring [70, 123]. Nelder [50]. Nested [215, 555]. Nests [322]. Netherlands [248]. Neuropathology [493]. Newbold [655]. Newell [383]. Nir [516]. Non [263, 2, 774, 373, 588, 27, 8, 110, 775, 709, 159, 190, 285, 217, 14, 670, 389, 31, 542, 555, 10, 681]. Non-Central [263, 2, 774, 27, 8, 775, 709, 190, 285, 31, 542, 10]. Non-Centrality [190]. Non-Linear [681]. Non-Negative [14]. Non-Nested [555]. Non-Parametric [588, 110, 217, 389]. Non-Proportional [373]. Non-Randomized [159]. Non-Sampling [670]. Noncentral [769, 436, 437, 607, 713, 583]. nonlinear [702]. Nonnegative [607, 713]. Nonparametric [639, 620, 644, 128, 721, 520, 287]. nonrandom [571]. Norm [686]. Normal [595, 260, 433, 387, 21, 472, 769, 319, 436, 537, 23, 337, 768, 238, 13, 27, 138, 564, 597, 310, 593, 235, 356, 439, 636, 14, 224, 628, 5, 16, 31, 542, 689, 32, 90, 137, 10, 220, 173, 177, 697, 147, 329]. Normal-Johnson [337, 13, 138]. Normality [641, 394, 225, 239, 311, 299, 529, 46]. normalizing [107]. Norman [408, 626]. North [104]. Note [257, 701, 450, 71, 98, 579, 494]. Notes [248]. null [149]. Number [315, 593, 459, 659, 241, 362, 506, 76]. Numbers [75, 237, 298, 724].

O [727]. O. [158]. Occupational [596]. Occurrence [39, 104]. O’Connell [627]. October [158]. Odds [588, 372]. Odeh [260]. O’Donovan [130]. oil [104]. Old [204]. Olympic [204]. One [18, 191, 162, 348, 751, 101]. One-pass [751]. One-Sample [348]. One-Sided [18, 191]. Operating [528]. Operational [522]. Operator [236, 740]. Optimal [488, 544, 386]. Optimization [670]. Optimum [561, 151]. Order [537, 23, 249, 306, 310, 359, 235, 636, 642, 118]. Ordered [540, 497, 134, 206, 120]. Ordering [85]. Ordinal [388]. Organising [501]. Organization [127]. Orthogonal [581, 22, 242]. Otis [410]. Outliers [230, 93, 42, 252, 450, 376, 572, 99, 231, 149]. overdispersion [745]. Overlap [70]. Owen [260, 452].

P [336, 157, 633, 410]. P. [326, 522, 60, 685, 328, 451, 129, 502, 408, 729]. P.P.S. [62]. Packages [553]. Page [109, 131]. Paired [465, 203]. pairs [673]. Pan [89, 160, 416]. Panel [453]. Paper [176]. Parameter [738, 435, 43, 561, 681, 190, 232, 101, 598]. Parameters [274, 481, 323, 711, 307, 482, 569, 747, 669]. Parametric [588, 110, 217, 389, 539]. Parent [395]. Parent-Child [395]. Pareto [519]. Park [763, 764]. Part [343, 482, 651]. Part-Time [651]. Partial [136, 633]. Partition [73]. pass [751]. Paths [552, 594]. Pattern [322, 280]. Patterns [584]. Paul [655, 333]. Payments [683]. PC [763]. Pearce [706, 384, 730]. Pearson [334, 105]. Peck [292]. Percentage [334, 619, 658, 38, 689, 105]. Performance [368, 518, 156]. Period [167]. Permutation [537, 265, 673]. Permutations [237, 297]. Perspective [734]. Perspectives [650]. Peter [452]. Pharmaceutical [510]. Pharmacologic [261]. Pharmacological [246]. Pharmacopoeia [672]. Philip [650]. Physical [298]. Physiologic [432]. Piece [597]. piecewise [144]. Plackett [626]. planning [41, 51, 53]. Plans [260, 277, 422, 719]. Players [573]. Plewis [523]. Plews [58]. plot [123]. Plots [418, 270, 189]. Plotting [278]. PLUS [735]. pocket [697]. Point [495, 447, 322, 380, 38, 174, 147, 77, 178]. Points [334, 619, 658, 689, 105]. Poisson [347, 214, 377, 227, 245, 720]. Polynomial [186]. Polynomials [210, 148]. Polytomous [540]. Population [458, 330, 228, 633, 374]. populations [171, 628, 410]. Portable [459, 241, 362, 506]. Positive [412, 267, 4]. Possible [230, 231]. post [722]. post-sample [722]. posterior [247, 494]. Posts [760]. potency [568]. Power [19, 24, 66, 140, 695, 550, 226, 227, 65, 171, 331]. Practical [716, 251, 444, 761, 34, 82, 287]. Practice [510, 432, 380, 384]. precision [147]. Predicting [146, 365]. Prediction [240]. Predictions [388]. predictive [422]. Pregnancy [282]. Presence [371, 330, 549, 517, 118]. present [702]. Preserve [551]. pressure [172]. Prevalence [316, 355, 676, 548, 427]. prices [669]. Primer [402]. Principal [319, 257, 704, 371, 551]. Principles [54]. probabilistic [700]. Probabilities [263, 769, 436, 134, 89, 160, 416, 774, 27, 775, 583, 268, 88, 308, 440, 378, 359, 563, 356, 439, 712, 31, 542, 10, 721, 426]. Probability [21, 181, 557, 493, 537, 603, 768, 453, 62, 624, 190, 57, 691, 616, 220, 558, 697, 751]. Problem [29, 493, 71, 151]. Problems [716, 269, 338, 280]. Procedure [48, 725, 89, 160, 416, 686, 342, 646]. Procedures [250, 395, 269, 644, 577, 97, 202]. Proceedings [158]. Process [495, 86, 248, 447, 280, 721, 719]. Processes [132, 87, 240, 647]. Product [768]. Production [41, 452]. Profile [618, 749]. Profile-Likelihood-Based [618]. Profiles [591]. prognosis [172]. Programs [261]. Projections [604]. Properties [512, 476, 425, 727]. Proportional [106, 373, 588, 372, 696, 463, 751]. Proportions [251, 698, 366, 143, 675]. Pseudo [75, 459, 241, 362, 506, 170]. Pseudo-Random [75, 459, 241, 362, 506, 170]. pseudorandom [724]. Psychological [650]. Published [279]. Purchasing [317, 701]. Purpose [320].

Q [336, 633]. Quadratic [14, 367]. Qualitative [55]. Qualitative/Categorical [55]. Quality [631, 392, 719]. Quangle [233]. Quantal [587, 443, 38]. Quantiles [605, 308, 440, 563]. Quantitative [473, 345, 186, 187]. Quantities [69]. quasilikelihood [645]. Questions [674]. Queueing [383]. Queues [606].

R [727, 385, 474, 328, 451, 625, 756, 410, 682]. R. [50, 623, 132, 432, 331, 291, 261, 430, 500, 409, 157, 403, 353, 57, 682, 53, 730, 332]. R12 [11]. R21 [16]. R25 [26]. R26 [27]. R27 [28]. R30 [31]. R31 [90]. R32 [137]. R33 [138]. R34 [139]. R35 [140]. R36 [191]. R37 [192]. R38 [193]. R39 [194]. R40 [195]. R41 [220]. R42 [242]. R44 [267]. R45 [268].

R46 [270]. R47 [310]. R48 [337]. R49 [338]. R50 [360]. R51 [361]. R52 [416]. R53 [417]. R54 [418]. R55 [436]. R56 [437]. R57 [438]. R58 [459]. R68 [562]. R69 [563]. R70 [564]. R71 [635]. R72 [636]. R73 [662]. R74 [663]. R75 [690]. R76 [711]. R77 [712]. R78 [740]. R79 [741]. R8 [8]. R80 [769]. R81 [770]. R84 [774]. R95 [775]. Races [365]. Racetrack [454]. Racquetball [573]. Radday [524]. Ralph [631]. Randall [407]. Random [75, 497, 132, 112, 412, 416, 417, 39, 22, 298, 701, 305, 459, 135, 320, 246, 242, 241, 362, 506, 170, 147, 599, 178]. Randomised [81]. Randomized [159, 429]. Randomly [569]. Randomness [324, 405]. Range [21, 308, 440, 563, 220]. Rank [250, 658]. Ranked [571, 150]. Ranking [577, 571]. Ranks [372]. Rapid [515, 589]. Rates [592, 295, 428, 175, 118]. Rating [573]. Ratio [40, 396, 574, 593, 672, 425, 489]. Rayleigh [324]. Raymond [631]. Read [626]. Readable [262]. Reader [701]. Reading [259, 533]. Recall [84]. Recapture [410]. Recovery [210, 148]. recruitment [41]. Recurrences [639]. Recursive [660, 686]. Reduced [250]. Reduced-Rank [250]. Reduction [402]. Rees [732]. Reference [452]. Reflectance [283]. reformulations [702]. region [569]. Regions [159]. Regression [595, 120, 207, 26, 29, 540, 303, 414, 513, 553, 195, 292, 666, 448, 250, 316, 355, 562, 283, 373, 338, 592, 703, 588, 444, 500, 257, 33, 605, 626, 381, 45, 744, 723, 543, 565, 15, 696, 175, 617, 246, 720, 367, 174, 462, 579, 72, 747, 702, 144, 102]. regressor [254]. Reinhold [380]. Related [263, 475, 129]. Relating [540]. Relationship [350]. relationships [72]. Relative [611, 468, 592, 450]. Release [50]. Reliability [716, 469, 498]. Reliable [737]. Remark [26, 28, 191, 769, 195, 140, 436, 437, 635, 544, 741, 418, 504, 770, 337, 662, 194, 11, 416, 417, 562, 139, 267, 774, 192, 193, 338, 270, 27, 138, 438, 564, 8, 360, 310, 361, 775, 268, 711, 459, 543, 507, 484, 740, 529, 563, 530, 636, 712, 16, 242, 31, 542, 90, 137, 506, 220, 26, 28, 191, 195, 140, 436, 437, 635, 663, 741, 418, 770, 337, 662, 194, 11, 416, 417, 562, 139, 267, 192, 193, 338, 270, 27, 138, 438, 564, 8, 360, 310, 361, 775, 268, 711, 459, 740, 563, 690, 636]. Remark [712, 16, 242, 31, 90, 137, 220]. Remarks [445, 690]. Removal [410, 374]. renal [77]. Renewal [674]. Repairable [498]. Repeat [245]. Repeat-Buying [245]. Repeated [738, 505, 591, 413]. Repetition [237]. Replacement [497, 378]. Replicate [211]. replicates [705]. Representation [109, 131, 98]. Representativeness [427]. Rescigno [432]. Research [522, 290, 452, 622, 757, 625]. resemblance [747]. Residuals [211, 72]. Resistant [662, 560]. Resource [695]. Respondent [288]. Response [679, 70, 445, 587, 209, 81, 38, 748]. Restricted [359]. Results [365, 429]. Retrospective [674]. Reviews [677, 55, 50, 754, 727, 260, 684, 382, 329, 473, 657, 289, 680, 52, 58, 654, 623, 499, 132, 326, 181, 522, 292, 404, 734, 762, 179, 521, 262, 60, 259, 679, 685, 759, 385, 383, 650, 474, 432, 331, 477, 431, 524, 603, 54, 706, 154, 631, 126, 127, 728, 290, 753, 61, 519, 629, 56, 155, 731, 733, 453, 525, 291, 454, 652, 130, 182, 261, 622, 632, 84, 288, 183, 430, 757, 655, 475, 707, 500, 409, 328, 451, 752, 736, 526, 110, 128, 354, 380, 627, 129, 157, 452, 626, 381, 407, 602, 406, 156]. Reviews [625, 683, 401, 653, 330, 735, 761, 520, 681, 403, 758, 353, 624, 405, 502, 478, 180, 630, 763, 764, 755, 756, 408, 455, 729, 82, 158, 57, 287, 333, 402, 83, 523, 327, 59, 51, 352, 109, 131, 81, 628, 732, 601, 501, 384, 410, 651, 682, 498, 476, 678, 760, 53, 730, 332, 656]. Revised [438, 335]. Reweighted [343]. Richard [652, 627]. Ridge [283, 444, 561]. Rindskopf [753]. Risk [468, 304]. Risks [611, 462]. Robert [260, 179, 477, 653, 478]. Robust [152, 97, 202, 99, 232, 496, 613, 737, 578, 394, 449, 630, 202]. Rodent [699, 568]. Rohlf [682]. Role [699]. Rosander [59]. Rotation [704, 594]. Routine [198]. Row [505, 113, 135, 413, 705]. Row-and-Column [113]. Rubin [328]. Run [511]. Running [204]. Runs [133]. Runs-Down [133]. Runs-Up [133]. Runyon [157].

S [654, 54, 183, 602, 405, 402, 352, 109, 131]. S. [52, 404, 406, 758, 501, 384, 730]. Sachs [354]. Salsburg [405]. Sample [492, 399, 240, 348, 377, 223, 252, 305, 633, 227, 59, 425, 517, 256, 149, 722, 599, 751]. Samples [397, 393, 238, 323, 275, 225, 628, 556, 168]. Sampling [260, 62, 722, 449, 378, 670, 410, 422, 151, 719, 571]. Samuel [329, 626]. Scale [188, 476, 178]. Scaling [671]. Scatter [418, 270, 189]. Scheduling [289]. Schnell [763, 764]. Schnieder [628]. School [453]. Schwarz [642]. Science [181, 156, 83]. Sciences [626, 758, 329, 110, 128]. score [392]. Scores [387, 100]. Scoring [691, 616]. Screening [548, 646, 187, 260]. Sea [104]. Search [513, 45]. Searle [385]. Seasonal [315, 309, 153]. Secondary [290]. Sectioning [87]. Seeds [536]. Segmented [45]. selecting [642]. Selection [497, 62, 551, 561, 264, 79, 634, 188, 38, 302, 172, 751, 571]. Selectivity [666]. Self [501, 169]. self-exciting [169]. Self-Organising [501]. Sellers [83]. Semi [267, 4]. Semi-Definite [267, 4]. Sen [408]. Sensitive [81]. Separating [319]. Sequence [352, 147, 178]. Sequences [306, 724]. Sequencing [289]. sequential [422, 143]. Series [154, 451, 94, 63, 633, 158, 398, 153, 426, 82]. Serum [540]. Set [604, 119, 571]. Sets [237, 770, 515, 25, 279, 152, 101, 571]. Several [581, 671, 675]. Shapiro [225, 299]. Shapiro-Wilk [299]. Shayle [385].

Shenton [727]. Shewhart [392]. Shewhart-cumulative [392]. shift [178]. Shimizu [728]. Shrunken [91, 36]. Sided [18, 191, 646]. Sidney [259]. Sign [348]. Significance [395, 388, 254]. Significance-Testing [395]. Signs [282]. Silicon [597]. Silverman [629]. Similar [614, 342]. Simon [289]. Simple [85, 686, 130, 658, 94, 63, 305, 285, 612, 119, 224, 177, 599]. Simply [134]. Simulating [512]. Simulation [301, 130, 215, 54]. Simultaneous [48]. single [572, 575]. Size [544, 386, 227, 256, 374, 751]. Sizes [517]. Skane [525]. Skewness [434, 46]. Small [476, 425, 105, 675]. Small-Scale [476]. Smallest [503]. Smith [681]. Smooth [552, 594]. Smoothing [447, 662, 9, 192, 560]. Social [60, 650, 290, 84, 757, 110, 128]. Softpower [736]. Software [657, 762, 685, 706, 632, 707, 736, 735, 763, 764, 526]. Soiling [248]. Sokal [682]. Solomon [156]. solution [151]. Some [445, 279, 299, 690, 246, 425, 104, 77]. Sources [759, 290, 407, 760]. Space [695, 401]. Space-Time [695]. Sparse [269]. Spatial [495, 322]. Spearman [212]. Special [52, 452]. Specified [458]. Spectra [553]. Spectral [615, 87, 438, 335, 280, 122]. Spectral-Temporal [615]. Spectrum [86, 9, 192]. Spell [703]. Sphere [450, 208, 168, 232, 370]. Spherical [552]. SPSS [656]. SPSS-X [656]. square [745]. Squared [190, 687]. Squares [387, 663, 698, 269, 690, 119, 163, 196, 343, 457, 72]. squash [700]. Stabilise [550]. Stablein [381]. Stage [515]. Stahl [259]. Stainton [183]. Standard [238]. Standards [535]. STAT [736]. state [721]. States [521]. Statgraphics [707]. stationarity [569]. Statistic [89, 160, 416, 483, 593, 105, 463]. Statistical [411, 6, 197, 198, 754, 716, 26, 29, 684, 329, 738, 433, 435, 657, 17, 387, 660, 21, 28, 237, 7, 654, 18, 191, 263, 528, 769, 414, 480, 195, 24, 66, 140, 762, 86, 87, 436, 437, 162, 236, 336, 635, 663, 685, 544, 759, 85, 134, 741, 431, 112, 524, 23, 334, 765, 581, 375, 541, 418, 527, 505, 504, 770, 337, 768, 662, 604, 706, 194, 503, 11, 89, 160, 412, 416, 417, 607, 562, 686, 713, 139, 456, 265, 267, 774, 9, 192, 193, 214, 111, 710, 269, 133, 338, 415]. Statistical [560, 270, 30, 483, 240, 238, 22, 62, 12, 13, 27, 182, 138, 438, 564, 632, 8, 183, 430, 306, 335, 481, 475, 688, 707, 113, 360, 736, 526, 33, 605, 310, 658, 357, 361, 20, 775, 452, 63, 218, 626, 25, 561, 583, 709, 766, 268, 264, 458, 407, 88, 308, 440, 136, 71, 216, 388, 737, 711, 661, 584, 64, 159, 309, 459, 358, 735, 761, 266, 520, 758, 353, 739, 543, 565, 516, 15, 190, 215, 767, 633, 135, 507, 359, 763, 764, 484, 634, 740, 386, 235, 239, 311, 434, 529, 563, 659, 690, 413, 356]. Statistical [439, 530, 217, 672, 687, 636, 708, 712, 14, 83, 219, 327, 51, 161, 307, 352, 606, 163, 188, 189, 196, 457, 482, 601, 185, 5, 16, 242, 389, 384, 31, 542, 760, 332, 241, 362, 689, 582, 32, 90, 137, 10, 506, 186, 187, 220, 590]. Statistics [181, 521, 262, 253, 331, 537, 23, 249, 753, 61, 629, 56, 453, 348, 655, 110, 128, 354, 310, 658, 295, 401, 624, 502, 180, 408, 455, 57, 287, 235, 402, 672, 636, 423, 453, 680, 58, 179, 259, 385, 431, 525, 157, 406, 729, 732]. STATLAB [761]. Stein [528]. Step [207]. Stephens [623]. Stepwise [513, 318]. Stereological [493, 423]. Stigler [152]. Stirling [761]. Stochastic [677, 495, 52, 104]. stock [669]. Stone [467]. Stoodley [183]. Storage [342]. Stores [317]. Straight [686, 77]. Strata [758]. strategies [700]. stratification [151]. Stratified [468, 357, 304]. Stratum [361, 161, 124]. Stratum-Matched [361, 161]. strict [667]. Stripes [39]. Stroup [404]. Structural [469]. Structure [55, 768, 371, 551, 330]. Structured [770, 25]. structures [423]. STSC [707]. Stuart [477]. Student [671, 107]. Studentized [263, 308, 440, 563]. Studies [468, 357, 361, 59, 161, 255, 76, 430]. Study [301, 524, 366, 596, 288, 490, 122, 103, 721]. Subject [604, 690, 163, 196]. Subjects [671]. Subset [513, 544, 634, 386]. Subtables [376]. success [118]. suddenly [669]. Sufficient [739]. Sugar [536]. Sullivan [763, 764]. Sum [26, 387, 698, 562, 15]. Sums [119]. Surface [70]. Surfaces [679]. Surrogate [720]. SURVCALC [685]. Survey [622, 223, 670]. surveys [538, 84]. Survival [611, 303, 676, 193, 373, 592, 471, 20, 381, 372, 555, 74, 534, 539, 748, 102]. Swamurthy [330]. Sweep [236, 740]. Symmetric [28, 195, 11, 267, 4, 3, 33]. Symmetrically [238]. Synapse [557, 493, 558]. Syndrome [490]. Syski [132]. System [50, 681, 648, 41]. Systematic [62]. Systems [127, 735, 691, 616, 498, 476]. Systems-Theoretical [127].

T [524, 155, 454]. T. [677, 331, 183]. Table [48, 738, 533, 717]. Tables [260, 17, 660, 505, 504, 139, 589, 129, 105, 264, 136, 159, 739, 135, 206, 413, 376, 667]. Tablets [672]. Tabulating [770, 25]. Tabulation [369]. Tactics [750]. Tail [89, 160, 416, 697]. Tallarida [261]. Tally [659]. Tan [630]. Taxonomy [156]. Taylor [171]. Teaching [253, 761]. Technique [60, 515, 69, 469]. Techniques [354, 299, 51, 623, 183, 407]. Telecommunications [760]. temperature [172]. Temporal [615]. Tennis [691, 616]. Tensor [729]. Test [48, 611, 19, 24, 66, 140, 698, 620, 214, 483, 438, 335, 377, 328, 388, 159, 593, 203, 225, 239, 311, 529, 119, 324, 65, 149, 143]. Testing [495, 336, 647, 395, 323, 641, 72, 46, 300]. Tests [75, 350, 396, 92, 251, 765, 168, 614, 133, 252, 427, 767, 206, 37, 217, 672, 227, 675, 673, 745]. Textbook [402]. Theil [733]. Their [498].

Thelma [521]. Theoretical [127]. Theory [432, 345, 728, 380, 755, 674, 384, 590, 383, 603, 652]. There [230, 231]. Think [680]. Thomas [130]. Three [281, 43, 320, 429, 598]. Three-Parameter [43, 598]. Three-Treatment [429]. threshold [169]. Tietz [409]. Tiku [630]. Timber [350]. Time [296, 154, 596, 695, 451, 633, 82, 158, 398, 153, 647, 534, 72, 147, 539, 178, 651]. time-dependent [534, 539]. Times [204, 146]. Tolerance [260, 587]. Tolerance-Distribution-Free [587]. Totals [505, 135, 413]. Tournaments [573]. Toxaemic [282]. Tracer [432]. training [556]. Transfer [345, 574]. Transform [662, 265, 9, 192, 560, 360, 530, 219]. Transformation [550, 169, 579]. Transformations [230, 337, 13, 138, 78, 107, 231]. Transformed [641, 749]. transition [721]. transplant [77]. Treatment [429]. Trees [527]. Trend [369, 368, 518]. Trends [406, 206, 60]. Trials [725, 445, 123, 118]. Triangular [28, 11, 267, 3]. Trimmed [238]. Trimming [275]. Truncated [597, 766, 628]. Tumorigenicity [699, 640, 568]. Tumour [316, 355, 676, 640]. Two [207, 17, 393, 414, 230, 231, 19, 24, 66, 140, 319, 251, 504, 167, 139, 515, 614, 348, 377, 471, 113, 597, 73, 543, 565, 280, 227, 646, 376, 65, 556, 101, 103, 721, 149, 143]. Two-Dimensional [280]. Two-Factor [73]. two-parameter [101]. Two-Period [167]. Two-Piece [597]. Two-Sample [348, 377, 227]. Two-Sided [646]. Two-Stage [515]. two-state [721]. Two-Step [207]. Two-Way [17, 504, 376, 139]. Type [501]. Types [471].

U.K. [176]. Ullah [650]. Unbalanced [397]. Uncertain [404]. Uncertainty [652]. Unconditional [159]. Unconstrained [562]. Understanding [259, 405].

Unemployment [492, 650]. Unequal [378]. Unimodality [483]. Unit [449]. United [521, 759, 760]. Units [70]. univariate [572]. Unknown [738, 323, 147, 178]. Unweighted [698]. Updating [466, 688, 739]. Upper [308, 440, 563, 149]. Uraemic [490]. Use [210, 596, 444, 223, 257, 469, 464, 640, 295, 226, 747, 153, 148, 123, 105, 697, 702, 117]. Useful [769, 436, 385, 27, 31, 542, 10]. User [656]. Using [595, 553, 319, 296, 448, 334, 662, 560, 360, 551, 491, 372, 57, 280, 530, 672, 219, 523, 606, 246, 646, 651, 376, 717, 74, 423, 72, 426, 102, 575].

V [477, 155, 129, 730]. validation [152, 172]. Validatory [79]. Validity [622]. Value [503, 365, 481, 718, 711, 74]. Values [613, 123, 117, 124]. Variable [607, 713, 752, 318, 646, 254, 144, 76]. Variables [414, 112, 770, 412, 416, 417, 551, 25, 79, 668, 320, 14, 517, 120, 170, 147, 175]. Variance [595, 738, 214, 549, 550, 579, 171, 121]. Variances [398]. Variate [91, 107, 36, 202]. Variates [399, 543, 565, 226]. Variation [514, 347, 549, 672, 229]. Variety [725, 123]. varying [719]. Vector [401, 708]. Vellenian [333]. Velocity [514]. Verma [706]. Version [762, 663, 706, 632, 707, 735]. Veterinary [83]. via [9, 192, 746, 43]. Vijay [706]. Virgil [478]. Vitamin [720]. Vol [759, 758, 760]. Volume [116, 201, 273, 341, 421, 487, 547, 610, 694, 773, 626]. Vries [683]. vs [548].

W [181, 629, 291, 328, 81, 332]. W. [759, 127, 182, 622, 526, 761, 630, 763, 764, 287, 760]. Walker [681, 476]. Walker-Smith [681]. Wall [684]. Wallis [60]. Walsh [631]. Walter [381]. Wampler [381]. Water [295, 759]. Watson [89, 160, 416]. Way [17, 504, 376, 139]. Wayne [81]. Weather [759]. Weekes [353]. Weibull [74, 598]. Weight [514, 672].

Weighted [269, 543, 696]. weighting [175]. Weights [492, 18, 191]. Weir [352]. White [410]. Whitney [342]. Wichern [332]. Wilk [225, 299]. William [181, 474, 454, 602, 653]. Wilson [109, 131]. Wind [695]. Window [301, 122]. Without [497, 162, 378]. Wold [380, 218]. Wolf [407]. Wonnacott [331]. Wood [53]. Woodward [526]. Working [535]. World [404]. Wrigley [701]. Wrong [626]. Wurster [631].

X [736, 656, 720]. X-STAT [736]. XVII [759]. XVIII [760].

Y. [524]. Yates [734]. Youth [650].

Zemroch [129]. Zero [543]. Ziemba [454]. ZPA [739]. Zuwaylif [61, 455]. Zyskind [758]. Zyskind-Martin [758].

References

Cooper:1968:AAA

[1] B. E. Cooper. Algorithms: Algorithm AS 4: An auxiliary function for distribution integrals. Applied Statistics, 17(2): 190–192, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                             tronic). URL http://lib.stat.cmu.

edu/apstat/4. See remarks [8].

Cooper:1968:AAIb

[2] B. E. Cooper. Algorithms: Algorithm AS 5: The integral of the non-central t-distribution. Applied Statistics, 17(2): 193–194, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                             tronic). URL http://lib.stat.cmu.

edu/apstat/5. See remarks [1, 8, 437].

Healy:1968:AAT

[3] M. J. R. Healy. Algorithms: Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 17 (2):195–197, June 1968. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/6. See corrigen-

dum [6] and remark [11, 267, 28].

Healy:1968:AAI

[4] M. J. R. Healy. Algorithms: Algorithm AS 7: Inversion of a positive semi-definite symmetric matrix. Applied Statistics, 17(2):198–199, June 1968. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/7.

See corrigendum [6] and remark [267].

Swan:1969:SAAa

[5] A. V. Swan. Statistical algorithms: Algorithm AS 16: Maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 18(1): 110–114, March 1969. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/16. See remarks

[564, 16].

Anonymous:1969:SAA

[6] Anonymous. Statistical algorithms: Algorithms AS 6 and AS 7: Corrigenda. Applied Statistics, 18(1):118, 1969. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [3, 4, 11, 28].

Bhattacharjee:1970:SAA

[7] G. P. Bhattacharjee. Statistical algorithms: Algorithm AS 32: The incomplete gamma integral. Applied Statistics, 19(3):285–287, September 1970. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/32. See re-

mark [484, 741].

Hitchin:1973:SAR

[8] D. Hitchin. Statistical algorithms: Remark AS R8: A remark on algorithms AS 4 and AS 5: An auxiliary function for distribution integrals, and the integral of the non-central t- distribution. Applied Statistics, 22(3):428, 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [1, 2].

Frome:1974:SAA

[9] E. L. Frome. Statistical algorithms: Algorithm AS 73: Cross-spectrum smoothing via the finite Fourier transform. Applied Statistics, 23(2):238– 244, June 1974. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/73. See remark [192].

Young:1974:SAA

[10] J. C. Young and Ch. E. Minder. Statistical algorithms: Algorithm AS 76: An integral useful in calculating noncentral t and bivariate normal probabilities. Applied Statistics, 23(3):455– 457, September 1974. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/76. See remarks

[27, 31, 436, 542, 769].

Farebrother:1974:SAR

[11] R. W. Farebrother and G. Berry. Statistical algorithms: Remark AS R12: A remark on Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 23(3):477, 1974. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/6. See [3, 6,

28].

Hill:1976:SAAa

[12] I. D. Hill, R. Hill, and R. L. Holder. Statistical algorithms: Algorithm AS 99: Fitting Johnson curves by moments. Applied Statistics, 25(2):180– 189, June 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/99. See remark [138].

Hill:1976:SAAb

[13] I. D. Hill. Statistical algorithms: Algorithm AS 100: Normal-Johnson and Johnson-normal transformations. Applied Statistics, 25(2):190–192, June 1976. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/100. See remark [138, 337].

Sheil:1977:SAA

[14] J. Sheil and I. O’Muircheartaigh. Statistical algorithms: Algorithm AS 106: The distribution of non-negative quadratic forms in normal variables. Applied Statistics, 26(1):92–98, March 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/106. See remark [417].

Narula:1977:SAA

[15] Subhash C. Narula and John F. Wellington. Statistical algorithms: Algorithm AS 108: Multiple linear regression with minimum sum of absolute errors. Applied Statistics, 26(1):106– 111, March 1977. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/108. See remark [26, 562].

Swan:1977:SAR

[16] A. V. Swan. Statistical algorithms: Remark AS R21: A remark on Algorithm AS 16 maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 26(1):122, 1977. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

                                        lib.stat.cmu.edu/apstat/16. See [5,

564].

Baker:1977:SAA

[17] R. J. Baker. Statistical algorithms: Algorithm AS 112: Exact distributions derived from two-way tables. Applied Statistics, 26(2):199–206, June 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/112. See remarks [139, 504].

Bohrer:1978:SAA

[18] Robert Bohrer and Winston Chow. Statistical algorithms: Algorithm AS 122: Weights for one-sided multivariate inference. Applied Statistics, 27(1):100– 104, March 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/122. See remark [191].

Casagrande:1978:PFE

[19] J. T. Casagrande, M. C. Pike, and P. G. Smith. The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 27(2):176–180, 1978. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic). See correction [65].

Lagakos:1978:SAA

[20] S. W. Lagakos and M. H. Kuhns. Statistical algorithms: Algorithm AS 125: Maximum likelihood estimation for censored exponential survival data with covariates. Applied Statistics, 27(2):190– 197, June 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/125. See remark [193].

Barnard:1978:SAA

[21] J. Barnard. Statistical algorithms: Algorithm AS 126: Probability integral of the normal range. Applied Statistics, 27(2): 197–198, June 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/126. See remark [220].

Heiberger:1978:SAA

[22] Richard M. Heiberger. Statistical algorithms: Algorithm AS 127: Generation of random orthogonal matrices. Applied Statistics, 27(2):199– 206, June 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/127. See remark [242].

Davis:1978:SAA

[23] C. S. Davis and M. A. Stephens. Statistical algorithms: Algorithm AS 128: Approximating the covariance matrix of normal order statistics. Applied Statistics, 27(2):206–212, June 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/128. See

remark [636].

Casagrande:1978:SAA

[24] J. T. Casagrande, M. C. Pike, and P. G. Smith. Statistical algorithms: Algorithm AS 129: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 27(2):212–219, June 1978. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/129. See correc-

tion [65] and remark [140].

Leathers:1978:SAA

[25] Burton L. Leathers. Statistical algorithms: Algorithm AS 131: Tabulating frequency distributions for variables with structured code sets. Applied Statistics, 27(3):359–362, September 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/131. See remark [770].

Armstrong:1978:SAR

[26] Ronald D. Armstrong and Edward L. Frome. Statistical algorithms: Remark AS R25: A remark on Algorithm AS 108: Multiple linear regression with minimum sum of absolute errors. Applied Statistics, 27(3):378, 1978. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

                                        /lib.stat.cmu.edu/apstat/108. See

[15, 562].

Hill:1978:SARa

[27] I. D. Hill. Statistical algorithms: Remark AS R26: A remark on Algorithm AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 27(3):379, 1978. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [10, 31, 436, 769, 542].

Barrett:1978:SAR

[28] J. C. Barrett and M. J. R. Healy. Statistical algorithms: Remark AS R27: A remark on Algorithm AS 6: Triangular decomposition of a symmetric matrix. Applied Statistics, 27(3): 379–380, 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/6. See [11, 3, 6].

Armstrong:1979:SAA

[29] Ronald D. Armstrong and David S. Kung. Statistical algorithms: Algorithm AS 135: Min-max estimates for a linear multiple regression problem. Applied Statistics, 28(1):93– 100, March 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/135. See remark [338].

Hartigan:1979:SAA

[30] J. A. Hartigan and M. A. Wong. Statistical algorithms: Algorithm AS 136: A K-means clustering algorithm. Applied Statistics, 28(1):100– 108, March 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/136. See remark [194].

Thomas:1979:SAR

[31] G. E. Thomas. Statistical algorithms: Remark AS R30: A remark on Algorithm AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 28(1):113, 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [27, 10, 769, 436, 542].

Wolynetz:1979:SAAb

[32] M. S. Wolynetz. Statistical algorithms: Algorithm AS 139: Maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 28(2):195– 206, June 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/139. See remarks [90, 137].

Kent:1979:SAA

[33] Philippe Kent. Statistical algorithms: Algorithm AS 141: Inversion of a symmetric matrix in regression models. Applied Statistics, 28(2):214– 217, June 1979. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/141. See remark [195].

Pike:1980:PDA

[34] Derek J. Pike. Practical data analysis. Applied Statistics, 29(1):1–2, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1980:G

[35] G. H. Freeman. Genstat. Applied Statistics, 29(1):2–4, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Campbell:1980:SED

[36] Norm A. Campbell. Shrunken estimators in discriminant and canonical variate analysis. Applied Statistics, 29(1):5– 14, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigenda [92].

Pregibon:1980:GLT

[37] Daryl Pregibon. Goodness of link tests for generalized linear models. Applied Statistics, 29(1):15–24, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigenda [91].

Tsutkawa:1980:SDL

[38] Robert K. Tsutkawa. Selection of dose levels for estimating a percentage point of a logistic quantal response curve. Applied Statistics, 29(1):25–33, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkes:1980:ROS

[39] Alan G. Hawkes. Random occurrence of stripes in a fabric. Applied Statistics, 29(1):34–38, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bennett:1980:DFM

[40] B. M. Bennett and R. Mease. On a distribution-free method for combining estimates of a ratio of means. Applied Statistics, 29(1):39–42, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Abodunde:1980:PPM

[41] T. T. Abodunde and S. I. McClean. Production planning for a manpower system with a constant level of recruitment. Applied Statistics, 29(1):43–49, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Collett:1980:OCD

[42] D. Collett. Outliers in circular data. Applied Statistics, 29(1):50–57, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigenda [93].

Griffiths:1980:IET

[43] David A. Griffiths. Interval estimation for the three-parameter lognormal distribution via the likelihood function. Applied Statistics, 29(1):58–68, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Donner:1980:ICC

[44] Allan Donner and Bernard Rosner. On inferences concerning a common correlation coefficient. Applied Statistics, 29 (1):69–76, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lerman:1980:FSR

[45] P. M. Lerman. Fitting segmented regression models by grid search. Applied Statistics, 29(1):77–84, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Small:1980:MMS

[46] N. J. H. Small. Miscellanea: Marginal skewness and kurtosis in testing multivariate normality. Applied Statistics, 29(1):85–87, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1980:FMM

[47] N. Davies and P. Newbold. Forecasting with misspecified models. Applied Statistics, 29(1):87–92, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitkin:1980:CST

[48] M. Aitkin. Correction: A simultaneous test procedure for contingency table models. Applied Statistics, 29(1):92, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stern:1980:LE

[49] G. J. A. Stern, P. D. Oldham, D. J. Newell, Douglas M. Hawkins, and P. T. Davies. Letters to the editors. Applied Statistics, 29(1):93–97, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altman:1980:BRG

[50] D. G. Altman. Book reviews: The GLIM System Release 3 Manual, by R. J. Baker and J. A. Nelder. Applied Statistics, 29(1):98, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1980:BRS

[51] A. R. Smith. Book reviews: Statistical Techniques for Manpower Planning, by David J. Bartholomew and Andrew F. Forbes. Applied Statistics, 29(1):99, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Biggins:1980:BRS

[52] J. D. Biggins. Book reviews: Stochastic Abundance Models with Emphasis on Biological Communities and Special Diversity, by S. Engen. Applied Statistics, 29(1):99–100, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Watts:1980:BRF

[53] Graham Watts. Book reviews: Forecasting and Planning, by R. Fildes and D. Wood. Applied Statistics, 29(1):100, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Downham:1980:BRP

[54] D. Y. Downham. Book reviews: Principles of Discrete Event Simulation, by George S. Fishman. Applied Statistics, 29(1):100–101, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitkin:1980:BRA

[55] M. Aitkin. Book reviews: Analyzing Qualitative/Categorical Data: Log-Linear Models and Latent Structure Analysis, by L. A. Goodman, J. A. Davis and J. Magidson. Applied Statistics, 29 (1):101–102, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1980:BRS

[56] Tony Greenfield. Book reviews: Statistics for Engineers, by A. Greer. Applied Statistics, 29(1):102–103, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reese:1980:BRI

[57] R. A. Reese. Book reviews: An Introduction to Probability and Statistics Using BASIC, by R. A. Groeneveld. Applied Statistics, 29(1):103–104, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1980:BRI

[58] Derek Bissell. Book reviews: Introductory Statistics, by A. M. Plews. Applied Statistics, 29(1):104–105, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1980:BRC

[59] T. M. F. Smith. Book reviews: Case Studies in Sample Design, by A. C. Rosander. Applied Statistics, 29(1):105, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chedzoy:1980:BRS

[60] O. B. Chedzoy. Book reviews: The Social Index. A New Technique for Measuring Social Trends, by P. J. Wallis. Applied Statistics, 29(1):105–106, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Green:1980:BRG

[61] J. R. Green. Book reviews: General Applied Statistics. 3rd Edition, by F. H. Zuwaylif. Applied Statistics, 29(1):106, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hidiroglou:1980:SAA

[62] M. A. Hidiroglou and G. B. Gray. Statistical algorithms: Algorithm AS 146: Construction of joint probability of selection for systematic P.P.S. sampling. Applied Statistics, 29(1):107– 112, March 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/146.

Lau:1980:SAA

[63] Chi-Leung L. Lau. Statistical algorithms: Algorithm AS 147: A simple series for the incomplete gamma integral. Applied Statistics, 29(1):113– 114, March 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/147. See corrigenda [94] and remarks [484, 741].

Matloff:1980:SAA

[64] Norman S. Matloff. Statistical algorithms: Algorithm AS 148: The jackknife. Applied Statistics, 29(1):115– 117, March 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/148. See remark [507].

Casagrande:1980:CPF

[65] J. T. Casagrande, M. C. Pike, and P. G. Smith. Correction to: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 29(1):118, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [19, 24].

Casagrande:1980:SAC

[66] J. T. Casagrande, M. C. Pike, and P. G. Smith. Statistical algorithms: Correction: Algorithm AS 129: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 29(1):118, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1980:BMPa

[67] Anonymous. Back matter. Applied Statistics, 29(1):[2], 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1980:FMPa

[68] Anonymous. Front matter. Applied Statistics, 29(1):[3], 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kass:1980:ETI

[69] G. V. Kass. An exploratory technique for investigating large quantities of categorical data. Applied Statistics, 29 (2):119–127, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Draper:1980:IOE

[70] Norman R. Draper and Irwin Guttman. Incorporating overlap effects from neighbouring units into response surface models. Applied Statistics, 29(2):128–134, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lwin:1980:NPS

[71] T. Lwin and J. S. Maritz. A note on the problem of statistical calibration. Applied Statistics, 29(2):135–141, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McCabe:1980:TCR

[72] B. P. M. McCabe and M. J. Harrison. Testing the constancy of regression relationships over time using least squares residuals. Applied Statistics, 29 (2):142–148, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Martin:1980:PTF

[73] K. J. Martin. A partition of a two-factor interaction, with an agricultural example. Applied Statistics, 29(2):149–155, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitkin:1980:FEW

[74] Murray Aitkin and David Clayton. The fitting of exponential, Weibull and extreme value distributions to complex censored survival data using GLIM. Applied Statistics, 29(2):156–163, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atkinson:1980:TPR

[75] A. C. Atkinson. Tests of pseudorandom numbers. Applied Statistics, 29 (2):164–171, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walter:1980:MCC

[76] S. D. Walter. Matched case-control studies with a variable number of controls per case. Applied Statistics, 29 (2):172–179, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1980:SLC

[77] A. F. M. Smith and D. G. Cook. Straight lines with a change-point: a Bayesian analysis of some renal transplant data. Applied Statistics, 29(2):180–189, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

John:1980:AFT

[78] J. A. John and N. R. Draper. An alternative family of transformations.

Applied Statistics, 29(2):190–197, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mabbett:1980:CVS

[79] A. Mabbett, M. Stone, and J. Washbrook. Cross-validatory selection of binary variables in differential diagnosis. Applied Statistics, 29(2):198–204, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1980:LE

[80] P. Prescott. Letter to the editors. Applied Statistics, 29(2):205, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sparks:1980:BRC

[81] David N. Sparks. Book reviews: Collecting Sensitive Data by Randomised Response: An Annotated Bibliography, by Wayne W. Daniel. Applied Statistics, 29(2):206, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ray:1980:BRP

[82] W. D. Ray. Book reviews: Practical Experiences with Modelling and Forecasting Time Series, by G. M. Jenkins. Applied Statistics, 29(2):206–207, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shotter:1980:BRS

[83] Margaret Shotter. Book reviews: Statistical Epidemiology in Veterinary Science, by F. B. Leech and K. C. Sellers. Applied Statistics, 29(2):207, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoinville:1980:BRR

[84] G. Hoinville. Book reviews: The Recall Method in Social Surveys, by L. Moss and H. Goldstein. Applied Statistics, 29 (2):207–208, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cran:1980:SAA

[85] G. W. Cran. Statistical algorithms: Algorithm AS 149: Amalgamation of means in the case of simple ordering. Applied Statistics, 29(2):209– 211, June 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/149.

Charnock:1980:SAAa

[86] David M. Charnock. Statistical algorithms: Algorithm AS 150: Spectrum estimate for a counting process. Applied Statistics, 29(2):211– 214, June 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/150.

Charnock:1980:SAAb

[87] David M. Charnock. Statistical algorithms: Algorithm AS 151: Spectral estimates for bivariate counting processes by sectioning the data. Applied Statistics, 29(2):214–220, June 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/151.

Lund:1980:SAA

[88] Richard E. Lund. Statistical algorithms: Algorithm AS 152: Cumulative hypergeometric probabilities. Applied Statistics, 29(2):221–223, June 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/152. See remark

[268, 712].

Farebrother:1980:SAA

[89] R. W. Farebrother. Statistical algorithms: Algorithm AS 153: Pan’s procedure for the tail probabilities of the Durbin-Watson statistic. Applied Statistics, 29(2):224–227, June 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/153. See

remark [416, 417].

Wolynetz:1980:SAR

[90] M. S. Wolynetz. Statistical algorithms: Remark AS R31: A remark on Algorithm AS 139. maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 29(2):228, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/139. See

[32, 137].

Anonymous:1980:CSE

[91] Anonymous. Corrigenda: “Shrunken Estimators in Discriminant and Canonical Variate Analysis”. Applied Statistics, 29(2):229, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [37].

Campbell:1980:CSE

[92] N. A. Campbell. Corrigenda: “Goodness of Link Tests for Generalized Linear Models”. Applied Statistics, 29(2):229, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [36].

Collett:1980:COC

[93] D. Collett. Corrigenda: Outliers in circular data. Applied Statistics, 29(2):229, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [42].

Lau:1980:CAS

[94] Chi-Leung Lau. Corrigenda: Algorithm AS 147. A simple series for the incomplete gamma integral. Applied Statistics, 29(2):229, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/147. See [63, 741].

Anonymous:1980:BMPb

[95] Anonymous. Back matter. Applied Statistics, 29(2):[2], 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1980:FMPb

[96] Anonymous. Front matter. Applied Statistics, 29(2):[3], 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Campbell:1980:RPM

[97] N. A. Campbell. Robust procedures in multivariate analysis. I. robust covariance estimation. Applied Statistics, 29 (3):231–237, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Banfield:1980:NGR

[98] C. F. Banfield and J. C. Gower. A note on the graphical representation of multivariate binary data. Applied Statistics, 29(3):238–245, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carroll:1980:RMF

[99] Raymond J. Carroll. Robust methods for factorial experiments with outliers. Applied Statistics, 29(3):246–251, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Munford:1980:CCB

[100] A. G. Munford. A control chart based on cumulative scores. Applied Statistics, 29(3):252–258, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1980:FTP

[101] G. H. Freeman. Fitting two-parameter discrete distributions to many data sets with one common parameter. Applied Statistics, 29(3):259–267, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whitehead:1980:FCR

[102] John Whitehead. Fitting Cox’s regression model to survival data using GLIM. Applied Statistics, 29(3):268–275, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gardner:1980:CET

[103] M. J. Gardner and A. G. Munford. The combined effect of two factors on disease in a case-control study. Applied Statistics, 29(3):276–281, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schagen:1980:SMO

[104] I. P. Schagen. A stochastic model for the occurrence of oil fields and its application to some North Sea data. Applied Statistics, 29(3):282–291, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawal:1980:TPP

[105] H. B. Lawal. Tables of percentage points of Pearson’s goodness-of-fit statistic for use with small expectations. Applied Statistics, 29(3):292–298, 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1980:PLM

[106] Martin J. Crowder. Proportional linear models. Applied Statistics, 29(3): 299–303, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1980:MAN

[107] B. J. R. Bailey. Miscellanea: Accurate normalizing transformations of a Student’s t variate. Applied Statistics, 29 (3):304–306, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Heiberger:1980:LE

[108] R. M. Heiberger and A. M. Kshirsagar. Letters to the editors. Applied Statistics, 29(3):307–308, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snook:1980:BRI

[109] Lionel I. Snook. Book reviews: Information Representation and Manipulation in a Computer. 2nd Edition, by E. S. Page and L. B. Wilson. Applied Statistics, 29(3):309, 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1980:BRI

[110] Richard Kay. Book reviews: Introduction to Statistics: A Non-Parametric Approach for the Social Sciences, by Chris Leach. Applied Statistics, 29 (3):309–310, 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gardner:1980:SAA

[111] G. Gardner, A. C. Harvey, and G. D. A.

Phillips. Statistical algorithms: Algorithm AS 154: An algorithm for exact maximum likelihood estimation of autoregressive-moving average models by means of Kalman filtering. Applied Statistics, 29(3):311–322, September 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/154.

Davies:1980:SAA

[112] Robert B. Davies. Statistical algorithms:

Algorithm AS 155: The distribution of a linear combination of χ^{2} random variables. Applied Statistics, 29(3):323–333, September 1980. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/155. See remark [417].

Jones:1980:SAA

[113] Byron Jones. Statistical algorithms: Algorithm AS 156: Combining two component designs to form a row-and-column design. Applied Statistics, 29(3):334– 337, September 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/156.

Anonymous:1980:BMPc

[114] Anonymous. Back matter. Applied Statistics, 29(3):[4], 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1980:FMPc

[115] Anonymous. Front matter. Applied Statistics, 29(3):[3], 1980. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1980:VIP

[116] Anonymous. Volume information. Applied Statistics, 29(3):[3], 1980. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1981:UAC

[117] Patricia L. Smith. The use of analysis of covariance to analyse data from designed experiments with missing or mixed-up values. Applied Statistics, 30(1):1–8, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1981:CSR

[118] R. J. Prescott. The comparison of success rates in cross-over trials in the presence of an order effect. Applied Statistics, 30(1):9–15, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schweder:1981:STS

[119] Tore Schweder. A simple test for a set of sums of squares. Applied Statistics, 30(1):16–21, 1981. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:1981:RDM

[120] J. A. Anderson and P. R. Philips. Regression, discrimination and measurement models for ordered categorical variables. Applied Statistics, 30(1):22– 31, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Raab:1981:EVF

[121] Gillian M. Raab. Estimation of a variance function, with application to immunoassay. Applied Statistics, 30(1):32– 40, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beamish:1981:SAW

[122] N. Beamish and M. B. Priestley. A study of autoregressive and window spectral estimation. Applied Statistics, 30(1):41– 58, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kempton:1981:UNP

[123] R. A. Kempton and C. W. Howes. The use of neighbouring plot values in the analysis of variety trials. Applied Statistics, 30(1):59–70, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1981:MEM

[124] E. R. Williams, D. Ratcliff, and T. P.

Speed. Miscellanea: Estimating missing values in multi-stratum experiments. Applied Statistics, 30(1):71–72, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mantel:1981:LE

[125] Nathan Mantel, Max Halperin, P. R. Mullins, and D. D. V. Morgan. Letters to the editors. Applied Statistics, 30(1):73–75, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1981:BRD

[126] G. H. Freeman. Book reviews: Design and Analysis of Experiments, by M. N. Das and N. C. Giri. Applied Statistics, 30(1):76, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

French:1981:BRO

[127] Simon French. Book reviews: Organization of Decision Making: A Systems-Theoretical Approach, by W. J. M. Kickert. Applied Statistics, 30(1):76–77, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1981:BRI

[128] Richard Kay. Book reviews: Introduction to Statistics: A Nonparametric Approach for the Social Sciences, by C. Leach. Applied Statistics, 30(1):77, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambert:1981:BRT

[129] Trevor Lambert. Book reviews: Tables of the F -Related Distributions with Algorithms, by K. V. Mardia and P. J. Zemroch. Applied Statistics, 30(1):78, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1981:BRG

[130] I. D. Hill. Book reviews: GPSSsimulation made simple, by Thomas M. O’Donovan. Applied Statistics, 30 (1):78–79, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snook:1981:BRI

[131] Lionel I. Snook. Book reviews: Information Representation and Manipulation in a Computer. 2nd Edition, by E. S. Page and L. B. Wilson. Applied Statistics, 30(1):79, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1981:BRR

[132] T. C. Brown. Book reviews: Random Processes: A First Look, by R. Syski. Applied Statistics, 30(1):80, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Grafton:1981:SAA

[133] R. G. T. Grafton. Statistical algorithms:

Algorithm AS 157: The runs-up and runs-down tests. Applied Statistics, 30 (1):81–85, March 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/157.

Cran:1981:SAA

[134] G. W. Cran. Statistical algorithms:

Algorithm AS 158: Calculation of the probabilities {P (l, k)} for the simply ordered alternative. Applied Statistics, 30 (1):85–91, March 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/158.

Patefield:1981:SAA

[135] W. M. Patefield. Statistical algorithms: Algorithm AS 159: An efficient method of generating random R × C tables with given row and column totals. Applied Statistics, 30(1):91–97, March 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/159.

Lustbader:1981:SAA

[136] Edward D. Lustbader and Robert K. Stodola. Statistical algorithms: Algorithm AS 160: Partial and marginal association in multidimensional contingency tables. Applied Statistics, 30(1): 97–105, March 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/160.

Wolynetz:1981:SAR

[137] M. S. Wolynetz. Statistical algorithms: Remark AS R32: A remark on Algorithm AS 139: Maximum likelihood estimation in a linear model from confined and censored normal data. Applied Statistics, 30(1):105, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/139. See

[32, 90].

Hill:1981:SAR

[138] I. D. Hill and R. E. Wheeler. Statistical algorithms: Remark AS R33: A remark on algorithms AS 99: Fitting Johnson curves by moments and AS 100: Normal-Johnson and Johnsonnormal transformations. Applied Statistics, 30(1):106, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [12, 13, 337].

Farley:1981:SAR

[139] T. M. M. Farley. Statistical algorithms:

Remark AS R34: A remark on Algorithm AS 112: Exact distributions derived from two-way tables. Applied Statistics, 30(1):106–107, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/112. See

[17, 504].

Casagrande:1981:SAR

[140] J. T. Casagrande. Statistical algorithms:

Remark AS R35: A remark on Algorithm AS 129: The power function of the “exact” test for comparing two binomial distributions. Applied Statistics, 30(1):108, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/129. See [24].

Anonymous:1981:BMPa

[141] Anonymous. Back matter. Applied Statistics, 30(1):[4], 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:FMPa

[142] Anonymous. Front matter. Applied Statistics, 30(1):[3], 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Meeker:1981:CST

[143] William Q. Meeker, Jr. A conditional sequential test for the equality of two binomial proportions. Applied Statistics, 30 (2):109–115, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tishler:1981:MLM

[144] Asher Tishler and Isreal Zang. A maximum likelihood method for piecewise regression models with a continuous dependent variable. Applied Statistics, 30 (2):116–124, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thompson:1981:CLF

[145] R. Thompson and R. J. Baker. Composite link functions in generalized linear models. Applied Statistics, 30(2): 125–131, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cunningham:1981:BDM

[146] R. B. Cunningham, T. Lewis, and A. G. L. Wilson. Biothermal development: A model for predicting the distribution of emergence times of diapausing Heliothis Armigera. Applied Statistics, 30(2):132–140, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Menzefricke:1981:BAC

[147] Ulrich Menzefricke. A Bayesian analysis of a change in the precision of a sequence of independent normal random variables at an unknown time point. Applied Statistics, 30(2):141–146, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1981:UED

[148] Martin J. Crowder and Jacki A. Tredger. The use of exponentially damped polynomials for biological recovery data. Applied Statistics, 30(2):147–152, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1981:NDT

[149] A. C. Kimber and H. J. Stevens. The null distribution of a test for two upper outliers in an exponential sample. Applied Statistics, 30(2):153–157, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shirley:1981:DFM

[150] Eryl A. C. Shirley. A distribution-free method for analysis of covariance based on ranked data. Applied Statistics, 30 (2):158–162, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jarque:1981:SPO

[151] Carlos M. Jarque. A solution to the problem of optimum stratification in multivariate sampling. Applied Statistics, 30(2):163–169, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Balasooriya:1981:MRE

[152] Uditha Balasooriya and Lai K. Chan.

Miscellanea: Robust estimation of Stigler’s data sets by cross validation. Applied Statistics, 30(2):170–172, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conradsen:1981:MSA

[153] Knut Conradsen and Henrik Spliid. Miscellanea: A seasonal adjustment filter for use in Box Jenkins analyses of seasonal time series. Applied Statistics, 30 (2):172–177, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farrow:1981:BRA

[154] M. Farrow. Book reviews: The Analysis of Time Series: An Introduction. 2nd Edition, by C. Chatfield. Applied Statistics, 30(2):178, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenfield:1981:BRM

[155] Tony Greenfield. Book reviews: Multivariate Analysis, by K. V. Mardia, J. T. Kent and J. M. Bibby. Applied Statistics, 30(2):178–179, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

MacFie:1981:BRT

[156] H. J. H. MacFie. Book reviews: Taxonomy and Behavioral Science: Comparative Performance of Grouping Methods, by J. E. Mezzich and H. Solomon. Applied Statistics, 30(2):179–180, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lambourne:1981:BRF

[157] R. D. Lambourne. Book reviews: Fundamentals of Behavioral Statistics, by R. P. Runyon and A. Haber. Applied Statistics, 30(2):180, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ray:1981:BRA

[158] W. D. Ray. Book reviews: Analysing Time Series: Proceedings of the International Conference Held in Guernsey, Channel Islands, in October 1979, by O. D. Anderson. Applied Statistics, 30 (2):180–181, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McDonald:1981:SAA

[159] Lyman L. McDonald, Henry R. Bauer III, Bruce M. Davis, and Betty Laby. Statistical algorithms: Algorithm AS 161: Critical regions of an unconditional non-randomized test of homogeneity in 2 × 2 contingency tables. Applied Statistics, 30(2):182–189, June 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/161.

Farebrother:1981:SAC

[160] R. W. Farebrother. Statistical algorithms: Correction: Algorithm AS 153: Pan’s procedure for the tail probabilities of the Durbin-Watson statistic. Applied Statistics, 30(2):189, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1981:SAA

[161] P. G. Smith, M. C. Pike, A. P. Hill, N. E. Breslow, and N. E. Day. Statistical algorithms: Algorithm AS 162: Multivariate conditional logistic analysis of stratum-matched case-control studies. Applied Statistics, 30(2):190– 197, June 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/162. See remark [361].

Clarke:1981:SAA

[162] M. R. B. Clarke. Statistical algorithms: Algorithm AS 163: A Givens algorithm for moving from one linear model to another without going back to the data. Applied Statistics, 30(2): 198–203, June 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/163.

Stirling:1981:SAAa

[163] W. Douglas Stirling. Statistical algorithms: Algorithm AS 164: Least squares subject to linear constraints. Applied Statistics, 30(2):204–212, June 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/164. See [690].

Anonymous:1981:BMPb

[164] Anonymous. Back matter. Applied Statistics, 30(2):[4], 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:FMPb

[165] Anonymous. Front matter. Applied Statistics, 30(2):[3], 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sturt:1981:CCF

[166] Elizabeth Sturt. Computerized construction in FORTRAN of a discriminant function for categorical data. Applied Statistics, 30(3):213–222, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunsmore:1981:GCT

[167] I. R. Dunsmore. Growth curves in twoperiod change over models. Applied Statistics, 30(3):223–229, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fisher:1981:TDS

[168] N. I. Fisher, T. Lewis, and M. E. Willcox. Tests of discordancy for samples from Fisher’s distribution on the sphere. Applied Statistics, 30(3):230–237, 1981.

CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ghaddar:1981:DTS

[169] D. K. Ghaddar and H. Tong. Data transformation and self-exciting threshold autoregression. Applied Statistics, 30 (3):238–248, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1981:EGL

[170] A. W. Kemp. Efficient generation of logarithmically distributed pseudorandom variables. Applied Statistics, 30 (3):249–253, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Perry:1981:TPL

[171] J. N. Perry. Taylor’s power law for dependence of variance on mean in animal populations. Applied Statistics, 30 (3):254–263, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hjorth:1981:MSB

[172] U. Hjorth and L. Holmqvist. On model selection based on validation with applications to pressure and temperature prognosis. Applied Statistics, 30(3): 264–274, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1981:AHA

[173] B. J. R. Bailey. Alternatives to Hastings’ approximation to the inverse of the normal cumulative distribution function. Applied Statistics, 30(3):275–276, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Esterby:1981:IAP

[174] S. R. Esterby and A. H. El-Shaarawi. Inference about the point of change in a regression model. Applied Statistics, 30 (3):277–285, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pocock:1981:RAM

[175] Stuart J. Pocock, Derek G. Cook, and Shirley A. A. Beresford. Regression of area mortality rates on explanatory variables: What weighting is appropriate? Applied Statistics, 30(3):286–295, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Boylan:1981:MFF

[176] T. A. Boylan and I. G. O’Muircheartaigh. Miscellanea: The functional form of the U.K. demand for money: A critique of a paper by mills. Applied Statistics, 30 (3):296–299, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lew:1981:MAC

[177] Robert A. Lew. Miscellanea: An approximation to the cumulative normal distribution with simple coefficients. Applied Statistics, 30(3):299–301, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Talwar:1981:MDS

[178] Prem P. Talwar and James E. Gentle. Miscellanea: Detecting a scale shift in a random sequence at an unknown time point. Applied Statistics, 30(3):301–304, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chapman:1981:BRB

[179] Myra B. Chapman. Book reviews: Business Statistics, by Robert L. Brite. Applied Statistics, 30(3):305, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pearce:1981:BRS

[180] S. C. Pearce. Book reviews: Statistics and Experimental Design. 2nd Edition, by Geoffrey M. Clarke. Applied Statistics, 30(3):305–306, 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bunday:1981:BRP

[181] Brian D. Bunday. Book reviews: Probability and Statistics in Engineering and Management Science. 2nd Edition, by William W. Hines and Douglas C. Montgomery. Applied Statistics, 30 (3):306–307, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1981:BRS

[182] I. D. Hill. Book reviews: Statistical Computing, by W. J. Kennedy, *Jr and J. E. Gentle. Applied Statistics, 30 (3):307–308, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holder:1981:BRA

[183] R. L. Holder. Book reviews: Applied Statistical Techniques, by K. D. C. Stoodley, T. Lewis and C. L. S. Stainton. Applied Statistics, 30(3):308–309, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thompson:1981:LE

[184] R. Thompson and Nathan Mantel. Letters to the editors. Applied Statistics, 30 (3):310–312, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sturt:1981:SAA

[185] E. Sturt. Statistical algorithms: Algorithm AS 165: An algorithm to construct a discriminant function in Fortran for categorical data. Applied Statistics, 30(3):313–325, September 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/165.

Zemroch:1981:SAAa

[186] P. J. Zemroch. Statistical algorithms: Algorithm AS 166: Generation of polynomial contrasts for incomplete factorial designs with quantitative levels. Applied Statistics, 30(3):325–333, September 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/166.

Zemroch:1981:SAAb

[187] P. J. Zemroch. Statistical algorithms: Algorithm AS 167: Screening algorithm for experimental designs with quantitative levels. Applied Statistics, 30(3): 334–339, September 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/167.

Stirling:1981:SAAb

[188] W. Douglas Stirling. Statistical algorithms: Algorithm AS 168: Scale selection and formatting. Applied Statistics, 30(3):339–344, September 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/168.

Stirling:1981:SAAc

[189] W. Douglas Stirling. Statistical algorithms: Algorithm AS 169: An improved algorithm for scatter plots. Applied Statistics, 30(3):345–349, September 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/169. See remarks [270, 418].

Narula:1981:SAA

[190] Subhash C. Narula and M. M. Desu.

Statistical algorithms: Algorithm AS 170: Computation of probability and non-centrality parameter of a noncentral chi-squared distribution. Applied Statistics, 30(3):349–352, September 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/170.

Bohrer:1981:SAR

[191] Robert Bohrer. Statistical algorithms:

Remark AS R36: A remark on AS 122: Weights for one-sided multivariable inference. Applied Statistics, 30 (3):352–353, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [18].

Frome:1981:SARa

[192] E. L. Frome. Statistical algorithms: Remark AS R37: A remark on AS 73: Cross-spectrum smoothing via the finite Fourier transform. Applied Statistics, 30(3):354, 1981. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic). See [9].

Frome:1981:SARb

[193] E. L. Frome. Statistical algorithms: Remark AS R38: A remark on AS 125: Maximum likelihood estimation for censored exponential survival data with covariates. Applied Statistics, 30(3):355, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [20].

England:1981:SAR

[194] R. England and D. Beynon. Statistical algorithms: Remark AS R39: A remark on AS 136: A K-means clustering algorithm. Applied Statistics, 30(3):355–356, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See remark [30].

Campbell:1981:SAR

[195] M. J. Campbell. Statistical algorithms: Remark AS R40: A remark on AS 141: Inversion of a symmetric matrix in regression models. Applied Statistics, 30(3):356, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [33].

Stirling:1981:SAC

[196] W. D. Stirling. Statistical algorithms: Correction: Algorithm AS 164: Least squares subject to linear constraints. Applied Statistics, 30(3):357, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:SAI

[197] Anonymous. Statistical algorithms: Index of statistical algorithms. Applied Statistics, 30(3):358–370, 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:SAL

[198] Anonymous. Statistical algorithms: List of routine names. Applied Statistics, 30 (3):371–373, 1981. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:BMPc

[199] Anonymous. Back matter. Applied Statistics, 30(3):[6], 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:FMPc

[200] Anonymous. Front matter. Applied Statistics, 30(3):[3], 1981. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1981:VIP

[201] Anonymous. Volume information. Applied Statistics, 30(3):[4], 1981. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Campbell:1982:RPM

[202] N. A. Campbell. Robust procedures in multivariate analysis II. Robust canonical variate analysis. Applied Statistics, 31(1):1–8, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosner:1982:GPT

[203] Bernard Rosner. A generalization of the paired t-test. Applied Statistics, 31(1):9– 13, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatterjee:1982:NLO

[204] Samprit Chatterjee and Sangit Chatterjee. New lamps for old: An exploratory analysis of running times in Olympic Games. Applied Statistics, 31(1):14–22, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vlachonikolis:1982:DMB

[205] I. G. Vlachonikolis and F. H. C. Marriott. Discrimination with mixed binary and continuous data. Applied Statistics, 31(1):23–31, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Patefield:1982:ETT

[206] W. M. Patefield. Exact tests for trends in ordered contingency tables. Applied Statistics, 31(1):32–43, 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anderson:1982:TSR

[207] J. A. Anderson and A. Senthilselvan. A two-step regression model for hazard functions. Applied Statistics, 31(1):44– 51, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wood:1982:BDS

[208] Andrew Wood. A bimodal distribution on the sphere. Applied Statistics, 31(1):52–58, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Korn:1982:CBI

[209] E. L. Korn. Confidence bands for isotonic dose-response curves. Applied Statistics, 31(1):59–63, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1982:CUE

[210] M. J. Crowder and J. A. Tredger.

Correction: The use of exponentially damped polynomials for biological recovery data. Applied Statistics, 31(1):63, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dyke:1982:MRE

[211] G. V. Dyke. Miscellanea: Residuals of equal magnitude in fractional replicate 2 n designs. Applied Statistics, 31(1):64– 65, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinneen:1982:DSF

[212] L. C. Dinneen and B. C. Blakesley. Definition of Spearman’s footrule. Applied Statistics, 31(1):66, 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinneen:1982:LE

[213] L. C. Dinneen and B. C. Blakesley.

Letter to the editors. Applied Statistics, 31(1):66, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Frome:1982:SAA

[214] E. L. Frome. Statistical algorithms: Algorithm AS 171: Fisher’s exact variance test for the Poisson distribution. Applied Statistics, 31(1):67–71, March 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/171.

OFlaherty:1982:SAA

[215] M. O’Flaherty and G. MacKenzie. Statistical algorithms: Algorithm AS 172: Direct simulation of nested Fortran DO-LOOPS. Applied Statistics, 31(1):71– 74, March 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/172.

MacKenzie:1982:SAA

[216] G. MacKenzie and M. O’Flaherty. Statistical algorithms: Algorithm AS 173: Direct design matrix generation for balanced factorial experiments. Applied Statistics, 31(1):74–80, March 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/173.

Schwertman:1982:SAA

[217] Neil C. Schwertman. Statistical algorithms: Algorithm AS 174: Multivariate multisample non-parametric tests. Applied Statistics, 31(1):80– 85, March 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/174.

Laurie:1982:SAA

[218] D. P. Laurie. Statistical algorithms: Algorithm AS 175: Cramér-Wold factorization. Applied Statistics, 31(1):86– 93, March 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/175.

Silverman:1982:SAA

[219] B. W. Silverman. Statistical algorithms: Algorithm AS 176: Kernel density estimation using the fast Fourier transform. Applied Statistics, 31(1):93– 99, March 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/176. See remark [360, 530].

elLozy:1982:SAR

[220] Mohamed el Lozy. Statistical algorithms: Remark AS R41: A remark on Algorithm AS 126: Probability integral of the normal range. Applied Statistics, 31(1):99, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/126. See [21].

Anonymous:1982:BMPa

[221] Anonymous. Back matter. Applied Statistics, 31(1):[4], 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1982:FMPa

[222] Anonymous. Front matter. Applied Statistics, 31(1):[3], 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnston:1982:UAB

[223] B. Johnston and J. F. Brewster. The use and analysis of a BIBD in a sample survey. Applied Statistics, 31(2):101–107, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shore:1982:SAI

[224] Haim Shore. Simple approximations for the inverse cumulative function, the density function and the loss integral of the normal distribution. Applied Statistics, 31(2):108–114, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Royston:1982:ESW

[225] J. P. Royston. An extension of Shapiro and Wilk’s W test for normality to large samples. Applied Statistics, 31 (2):115–124, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rothery:1982:UCV

[226] P. Rothery. The use of control variates in Monte Carlo estimation of power. Applied Statistics, 31(2):125–129, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shiue:1982:ESP

[227] Wei-Kei Shiue and Lee J. Bain. Experiment size and power comparisons for two-sample Poisson tests. Applied Statistics, 31(2):130–134, 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mountford:1982:EPF

[228] M. D. Mountford. Estimation of population fluctuations with application to the common bird census. Applied Statistics, 31(2):135–143, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1982:EBV

[229] D. A. Williams. Extra-binomial variation in logistic linear models. Applied Statistics, 31(2):144–148, 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carroll:1982:MTE

[230] R. J. Carroll. Miscellanea: Two examples of transformations when there are possible outliers. Applied Statistics, 31 (2):149–152, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Carroll:1982:TET

[231] R. J. Carroll. Two examples of transformations when there are possible outliers. Applied Statistics, 31(2):149–152, 1982.

Fisher:1982:MRE

[232] N. I. Fisher. Miscellanea: Robust estimation of the concentration parameter of Fisher’s distribution on the sphere. Applied Statistics, 31(2):152–154, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

North:1982:MQM

[233] W. R. S. North. Miscellanea: The quangle — A modification of the cusum chart. Applied Statistics, 31(2):155–158, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Scobey:1982:LE

[234] P. Scobey and D. G. Kabe. Letter to the editors. Applied Statistics, 31 (2):159–160, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Royston:1982:SAAa

[235] J. P. Royston. Statistical algorithms:

Algorithm AS 177: Expected normal order statistics (exact and approximate). Applied Statistics, 31(2):161– 165, June 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/177. See remark [310].

Clarke:1982:SAA

[236] M. R. B. Clarke. Statistical algorithms: Algorithm AS 178: The Gauss– Jordan sweep operator with detection of collinearity. Applied Statistics, 31(2): 166–168, June 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/178. See remark [740].

Berry:1982:SAA

[237] Kenneth J. Berry. Statistical algorithms: Algorithm AS 179: Enumeration of all permutations of multi-sets with fixed repetition numbers. Applied Statistics, 31(2):169–173, June 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/179.

Healy:1982:SAA

[238] M. J. R. Healy. Statistical algorithms: Algorithm AS 180: A linear estimator of standard deviation in symmetrically trimmed normal samples. Applied Statistics, 31(2):174– 175, June 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/180.

Royston:1982:SAAb

[239] J. P. Royston. Statistical algorithms: Algorithm AS 181: The W test for normality. Applied Statistics, 31(2):176– 180, June 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/181. See remark [529].

Harvey:1982:SAA

[240] A. C. Harvey and C. R. McKenzie. Statistical algorithms: Algorithm AS 182:

Finite sample prediction from ARIMA processes. Applied Statistics, 31(2):180– 187, June 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/182.

Wichmann:1982:SAA

[241] B. A. Wichmann and I. D. Hill. Statistical algorithms: Algorithm AS 183: An efficient and portable pseudo-random number generator. Applied Statistics, 31 (2):188–190, June 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/183. See correc-

tion [362] and remarks [459, 506].

Tanner:1982:SAR

[242] Martin A. Tanner and Ronald A.

Thisted. Statistical algorithms: Remark AS R42: A remark on AS 127. generation of random orthogonal matrices. Applied Statistics, 31(2):190–192, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [22].

Anonymous:1982:BMPb

[243] Anonymous. Back matter. Applied Statistics, 31(2):[4], 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1982:FMPb

[244] Anonymous. Front matter. Applied Statistics, 31(2):[3], 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sichel:1982:RBG

[245] H. S. Sichel. Repeat-buying and the generalized inverse Gaussian-Poisson distribution. Applied Statistics, 31(3): 193–204, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sweeting:1982:BAS

[246] T. J. Sweeting. A Bayesian analysis of some pharmacological data using a random coefficient regression model. Applied Statistics, 31(3):205–213, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Naylor:1982:AME

[247] J. C. Naylor and A. F. M. Smith. Applications of a method for the efficient computation of posterior distributions. Applied Statistics, 31(3):214–225, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

DenButter:1982:PSL

[248] F. A. G. Den Butter and R. L. Coenen. The process of soiling and the life of bank notes in the Netherlands. Applied Statistics, 31(3):226–237, 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gerontidis:1982:MCG

[249] I. Gerontidis and R. L. Smith. Monte Carlo generation of order statistics from general distributions. Applied Statistics, 31(3):238–243, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1982:PRR

[250] P. T. Davies and M. K.-S. Tso. Procedures for reduced-rank regression. Applied Statistics, 31(3):244–255, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Choi:1982:PTC

[251] S. C. Choi and D. M. Stablein. Practical tests for comparing two proportions with incomplete data. Applied Statistics, 31(3):256–262, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1982:TMO

[252] A. C. Kimber. Tests for many outliers in an exponential sample. Applied Statistics, 31(3):263–271, 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1982:TCA

[253] C. Chatfield. Teaching a course in applied statistics. Applied Statistics, 31 (3):272–289, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Butler:1982:BSA

[254] R. W. Butler. Bounds on the significance attained by the best-fitting regressor variable. Applied Statistics, 31 (3):290–292, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walker:1982:EAC

[255] Alexander M. Walker. Efficient assessment of confounder effects in matched follow-up studies. Applied Statistics, 31 (3):293–297, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Green:1982:MAS

[256] Jack Green. Miscellanea: Asymptotic sample size for given confidence interval length. Applied Statistics, 31(3):

298–300, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jolliffe:1982:MNU

[257] Ian T. Jolliffe. Miscellanea: A note on the use of principal components in regression. Applied Statistics, 31(3): 300–303, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Oxley:1982:LE

[258] L. T. Oxley and George A. Barnard. Letters to the editors. Applied Statistics, 31 (3):304–305, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chedzoy:1982:BRR

[259] O. B. Chedzoy. Book reviews: Reading and Understanding Applied Statistics, by Sidney M. Stahl and James D. Hennes. Applied Statistics, 31(3):306, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Armstrong:1982:BRT

[260] W. Armstrong. Book reviews: Tables for Normal Tolerance Limits, Sampling Plans, and Screening, by Robert E. Odeh and D. B. Owen. Applied Statistics, 31(3):306–307, 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1982:BRM

[261] I. D. Hill. Book reviews: Manual of Pharmacologic Calculations with Computer Programs, by R. J. Tallarida and R. B. Murray. Applied Statistics, 31 (3):307–308, 1982. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Chatfield:1982:BRR

[262] C. Chatfield. Book reviews: The Readable Maths and Statistics Book, by Barry Edwards. Applied Statistics, 31(3):308, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bohrer:1982:SAA

[263] Robert Bohrer, Mark Schervish, and Judith Sheft. Statistical algorithms: Algorithm AS 184: Non-central Studentized maximum and related multiple-t probabilities. Applied Statistics, 31(3): 309–317, September 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/184.

Lin:1982:SAA

[264] Shang P. Lin. Statistical algorithms:

Algorithm AS 185: Automatic model selection in contingency tables. Applied Statistics, 31(3):317–326, September 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/185.

Francik:1982:SAF

[265] Andrzej Francik and Janusz Kościelniak.

Statistical algorithms: Fast algorithm of data permutation in discrete fast Fourier transform. Applied Statistics, 31(3):327– 330, September 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moore:1982:SAA

[266] R. J. Moore. Statistical algorithms:

Algorithm AS 187: Derivatives of the incomplete gamma integral. Applied Statistics, 31(3):330–335, September 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/187.

Freeman:1982:SAR

[267] P. R. Freeman. Statistical algorithms: Remark AS R44: A remark on AS 6 and AS 7: Triangular decomposition of a symmetric matrix and inversion of a positive semi-definite symmetric matrix. Applied Statistics, 31(3):336–339, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [3, 4].

Lillestol:1982:SAR

[268] J. Lillestol. Statistical algorithms: Remark AS R45: Remark on AS 152. cumulative hypergeometric probabilities. Applied Statistics, 31(3):339–340, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [88, 712].

Gentleman:1982:SAC

[269] W. Morven Gentleman. Statistical algorithms: Correction: AS 75: Basic procedures for large, sparse or weighted linear squares problems. Applied Statistics, 31(3):340, 1982. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harris:1982:SAR

[270] R. G. Harris. Statistical algorithms: Remark AS R46: Remark on AS 169: An improved algorithm for scatter plots. Applied Statistics, 31(3):340, 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [189, 418].

Anonymous:1982:BMPc

[271] Anonymous. Back matter. Applied Statistics, 31(3):[4], 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1982:FMPc

[272] Anonymous. Front matter. Applied Statistics, 31(3):[3], 1982. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1982:VIP

[273] Anonymous. Volume information. Applied Statistics, 31(3):[3], 1982. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berman:1983:EPC

[274] Mark Berman. Estimating the parameters of a circle when angular differences are known. Applied Statistics, 32(1):1– 6, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1983:TGS

[275] A. C. Kimber. Trimming in gamma samples. Applied Statistics, 32(1):7–14, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Crowder:1983:GCA

[276] Martin Crowder. A growth curve analysis for EDP curves. Applied Statistics, 32(1):15–18, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Osanaiye:1983:CDI

[277] P. A. Osanaiye. Chain-deferred inspection plans. Applied Statistics, 32(1):19– 24, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Copas:1983:PA

[278] J. B. Copas. Plotting p against x. Applied Statistics, 32(1):25–31, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paulson:1983:IDE

[279] A. S. Paulson and E. H. Nicklin. Integrated distance estimators for linear models applied to some published data sets. Applied Statistics, 32(1):32–50, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Renshaw:1983:IPP

[280] E. Renshaw and E. D. Ford. Interpretation of process from pattern using twodimensional spectral analysis: Methods and problems of interpretation. Applied Statistics, 32(1):51–63, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clayton:1983:CTG

[281] D. G. Clayton. Comparing three groups. Applied Statistics, 32(1):64–68, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brown:1983:TSD

[282] P. J. Brown, J. Stone, and C. Ord-Smith. Toxaemic signs during pregnancy. Applied Statistics, 32(1):69–72, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fearn:1983:MRR

[283] T. Fearn. Misuse of ridge regression in the calibration of a near infrared reflectance instrument. Applied Statistics, 32(1):73–79, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1983:IMB

[284] J. Taylor and J. A. John. The incidence matrix of a binary incomplete block design with a given concurrence matrix. Applied Statistics, 32(1):80–83, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Norton:1983:MSA

[285] Vic Norton. Miscellanea: A simple algorithm for computing the non-central F distribution. Applied Statistics, 32 (1):84–85, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reid:1983:LE

[286] D. D. Reid, R. L. Sandland, R. Wootton, J. P. Royston, and David Griffiths. Letters to the editors. Applied Statistics, 32(1):86–90, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Reese:1983:BRP

[287] R. Allan Reese. Book reviews: Practical Nonparametric Statistics. 2nd Edition, by W. J. Conover. Applied Statistics, 32(1):91, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoinville:1983:BRM

[288] Gerald Hoinville. Book reviews: Measurement of Respondent Burden: Study Design and Early Findings, by Joanne Frankel. Applied Statistics, 32(1):92, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Baker:1983:BRS

[289] Anthony G. Baker. Book reviews: Sequencing and Scheduling, by Simon French. Applied Statistics, 32(1):92–93, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fuller:1983:BRS

[290] M. Fuller. Book reviews: Secondary Analysis in Social Research: A Guide to Data Sources and Methods with Examples, by Catherine Hakim. Applied Statistics, 32(1):93–94, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Healy:1983:BRB

[291] M. J. R. Healy. Book reviews: Biostatistics Casebook, by R. G. Miller Jr, B. Efron, B. W. Brown Jr and L. E. Moses. Applied Statistics, 32(1):94, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cass:1983:BRI

[292] J. M. Cass. Book reviews: Introduction to Linear Regression Analysis, by D. C. Montgomery and E. A. Peck. Applied Statistics, 32(1):94–95, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:BMPa

[293] Anonymous. Back matter. Applied Statistics, 32(1):[6], 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:FMPa

[294] Anonymous. Front matter. Applied Statistics, 32(1):[3], 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maritz:1983:USE

[295] J. S. Maritz and R. G. Jarrett. The use of statistics to examine the association between fluoride in drinking water and cancer death rates. Applied Statistics, 32(2):97–101, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clayton:1983:FGF

[296] D. G. Clayton. Fitting a general family of failure-time distributions using Glim. Applied Statistics, 32(2):102–109, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tallis:1983:AAA

[297] G. M. Tallis and B. R. Dansie. An alternative approach to the analysis of permutations. Applied Statistics, 32 (2):110–114, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Inoue:1983:RNG

[298] H. Inoue, H. Kumahora, Y. Yoshizawa, M. Ichimura, and O. Miyatake. Random numbers generated by a physical device. Applied Statistics, 32(2):115–120, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Royston:1983:STA

[299] J. P. Royston. Some techniques for assessing multivariate normality based on the Shapiro-Wilk W. Applied Statistics, 32(2):121–133, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snell:1983:RLT

[300] Martin C. Snell. Recent literature on testing for intergroup concordance. Applied Statistics, 32(2):134–140, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bhansali:1983:SSA

[301] R. J. Bhansali. Simulation study of autoregressive and window estimators of the inverse correlation function. Applied Statistics, 32(2):141–149, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Waterton:1983:ECS

[302] Jennifer J. Waterton. An exercise in controlled selection. Applied Statistics, 32 (2):150–164, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bennett:1983:LLR

[303] Steve Bennett. Log-logistic regression models for survival data. Applied Statistics, 32(2):165–171, 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Storer:1983:MLF

[304] Barry E. Storer, Sholom Wacholder, and Norman E. Breslow. Maximum likelihood fitting of general risk models to stratified data. Applied Statistics, 32 (2):172–181, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McLeod:1983:MCA

[305] A. I. McLeod and D. R. Bellhouse. Miscellanea: A convenient algorithm for drawing a simple random sample.

Applied Statistics, 32(2):182–184, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hopkins:1983:SAAa

[306] T. R. Hopkins, P. H. Welch, and J. Köllerström. Statistical algorithms: Algorithm AS 188: Estimation of the order of dependence in sequences. Applied Statistics, 32(2):185–196, June 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/188.

Smith:1983:SAA

[307] D. M. Smith. Statistical algorithms: Algorithm AS 189: Maximum likelihood estimation of the parameters of the beta binomial distribution. Applied Statistics, 32(2):196–204, June 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/189.

Lund:1983:SAA

[308] R. E. Lund and J. R. Lund. Statistical algorithms: Algorithm AS 190: Probabilities and upper quantiles for the Studentized range. Applied Statistics, 32(2): 204–210, June 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/190. See [563].

McLeod:1983:SAA

[309] A. I. McLeod and P. R. Holanda Sales.

Statistical algorithms: Algorithm AS 191: An algorithm for approximate likelihood calculation of ARMA and seasonal ARMA models. Applied Statistics, 32(2):211–223, June 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/191.

Koniger:1983:SAR

[310] W. Koniger. Statistical algorithms: Remark AS R47: A remark on AS 177. expected normal order statistics (exact and approximate). Applied Statistics, 32 (2):223–224, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [235].

Royston:1983:SAC

[311] J. P. Royston. Statistical algorithms: Correction: Algorithm AS 181: The W test for normality. Applied Statistics, 32(2):224, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:BMPb

[312] Anonymous. Back matter. Applied Statistics, 32(2):[3], 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:FMPb

[313] Anonymous. Front matter. Applied Statistics, 32(2):[3], 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:EGA

[314] Anonymous. Editorial: Guidelines for authors. Applied Statistics, 32(3): 225–227, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cohen:1983:SDE

[315] Ayala Cohen. Seasonal daily effect on the number of births in israel. Applied Statistics, 32(3):228–235, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinse:1983:RAT

[316] Gregg E. Dinse and S. W. Lagakos. Regression analysis of tumour prevalence data. Applied Statistics, 32(3):236–248, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunn:1983:IAN

[317] R. Dunn, S. Reader, and N. Wrigley. An investigation of the assumptions of the NBD model as applied to purchasing at individual stores. Applied Statistics, 32 (3):249–259, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krzanowski:1983:SLM

[318] W. J. Krzanowski. Stepwise location model choice in mixed-variable discrimination. Applied Statistics, 32(3): 260–266, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chang:1983:UPC

[319] Wei-Chien Chang. On using principal components before separating a mixture of two multivariate normal distributions. Applied Statistics, 32(3):267–275, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Peterson:1983:ACT

[320] Arthur V. Peterson, Jr. and Richard A.

Kronmal. Analytic comparison of three general-purpose methods for the computer generation of discrete random variables. Applied Statistics, 32(3): 276–286, 1983. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

OMuircheartaigh:1983:FFI

[321] I. G. O’Muircheartaigh and J. Sheil. Fore or five?-the indexing of a golf course. Applied Statistics, 32(3): 287–292, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harkness:1983:BSP

[322] R. D. Harkness and Valerie Isham. A bivariate spatial point pattern of ants’ nests. Applied Statistics, 32(3):293–303, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1983:DTG

[323] A. C. Kimber. Discordancy testing in gamma samples with both parameters unknown. Applied Statistics, 32 (3):304–310, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilkie:1983:MRT

[324] D. Wilkie. Miscellanea: Rayleigh test for randomness of circular data. Applied Statistics, 32(3):311–312, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Duong:1983:LE

[325] Q. P. Duong and M. J. Maher. Letters to the editors. Applied Statistics, 32 (3):313–314, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bryan-Jones:1983:BRI

[326] J. Bryan-Jones. Book reviews: An Introduction to Genstat, by N. Alvey, N.

Galwey and P. Lane. Applied Statistics, 32(3):315, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Slater:1983:BRB

[327] M. Slater. Book reviews: Basic Statistical Computing, by D. Cooke, A. H. Craven and G. M. Clarke. Applied Statistics, 32(3):315–316, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:1983:BRT

[328] P. H. Jackson. Book reviews: Test Equating, by P. W. Holland and D. R. Rubin. Applied Statistics, 32(3): 316–317, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aspinall:1983:BRE

[329] L. J. Aspinall. Book reviews: Encyclopedia of Statistical Sciences, by Samuel Kotz and Normal L. Johnson. Applied Statistics, 32(3):317–318, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McClean:1983:BRG

[330] Sally McClean. Book reviews: Growth and Structure of Human Population in the Presence of Migration, by M. Swamurthy. Applied Statistics, 32(3):318, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Daldry:1983:BRS

[331] S. J. Daldry. Book reviews: Statistics:

Discovering its Power, by T. H. Wonnacott and R. J. Wonnacott. Applied Statistics, 32(3):319–320, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whelehan:1983:BRA

[332] O. Whelehan. Book reviews: Applied Multivariate Statistical Analysis, by R. A. Johnson and D. W. Wichern. Applied Statistics, 32(3):320, 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ross:1983:BRA

[333] G. J. S. Ross. Book reviews: Applications, Basics and Computing of Exploratory Data Analysis, by Paul F. Vellenian and David C. Hoaglin. Applied Statistics, 32(3):320–321, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davis:1983:SAA

[334] C. S. Davis and M. A. Stephens. Statistical algorithms: Algorithm AS 192: Approximate percentage points using Pearson curves. Applied Statistics, 32(3): 322–327, September 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/192.

Hopkins:1983:SAAb

[335] T. R. Hopkins. Statistical algorithms: Algorithm AS 193: A revised algorithm for the spectral test. Applied Statistics, 32(3):328–335, September 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/193. See

remark [438].

Clarke:1983:SAA

[336] Brenton R. Clarke. Statistical algorithms: Algorithm AS 194: An algorithm for testing goodness of fit of ARMA(P,Q) models. Applied Statistics, 32(3):335–344, September 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/194.

Dodgson:1983:SAR

[337] J. H. Dodgson and I. D. Hill. Statistical algorithms: Remark AS R48: A remark on algorithm 100: Normal-Johnson and Johnson-normal transformations. Applied Statistics, 32(3):345, 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [13, 138].

Grant:1983:SAR

[338] P. M. Grant and T. R. Hopkins. Statistical algorithms: Remark AS R49: A remark on Algorithm AS 135: Minmax estimates for linear multiple regression problems. Applied Statistics, 32 (3):345–347, 1983. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/135. See [29].

Anonymous:1983:BMPc

[339] Anonymous. Back matter. Applied Statistics, 32(3):[5], 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:FMPc

[340] Anonymous. Front matter. Applied Statistics, 32(3):[3], 1983. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1983:VIP

[341] Anonymous. Volume information. Applied Statistics, 32(3):[3], 1983. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harding:1984:EMS

[342] E. F. Harding. Efficient, minimalstorage procedure for calculating the Mann–Whitney U, generalized U and similar distributions. Applied Statistics, 33(1):1–6, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stirling:1984:IRL

[343] W. Douglas Stirling. Iteratively reweighted least squares for models with a linear part. Applied Statistics, 33(1):7– 17, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1984:GLM

[344] Christopher Cox. Generalized linear models — the missing link. Applied Statistics, 33(1):18–24, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evett:1984:QTI

[345] I. W. Evett. A quantitative theory for interpreting transfer evidence in criminal cases. Applied Statistics, 33(1):25– 32, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Abeyasekera:1984:DCA

[346] S. Abeyasekera. The desirability of covariance adjustments. Applied Statistics, 33(1):33–37, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Breslow:1984:EPV

[347] N. E. Breslow. Extra-Poisson variation in log-linear models. Applied Statistics, 33(1):38–44, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hettmansperger:1984:TSI

[348] Thomas P. Hettmansperger. Twosample inference based on one-sample sign statistics. Applied Statistics, 33 (1):45–51, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittaker:1984:MIA

[349] Joe Whittaker. Model interpretation from the additive elements of the likelihood function. Applied Statistics, 33 (1):52–64, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bartlett:1984:ERB

[350] N. R. Bartlett and T. Lwin. Estimating a relationship between different destructive tests on timber. Applied Statistics, 33(1):65–72, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wright:1984:LE

[351] J. T. Wright, P. J. Brown, J. Stone, C. Ord-Smith, R. W. Farebrother, and M. E. Dewey. Letters to the editors. Applied Statistics, 33(1):73–76, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1984:BRS

[352] Cedric A. B. Smith. Book reviews: Statistical Analysis of DNA Sequence Data, by B. S. Weir. Applied Statistics, 33(1): 77, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Muir:1984:BRD

[353] W. W. Muir. Book reviews: Data Models and Statistical Analysis, by R. A. Cooper and A. J. Weekes. Applied Statistics, 33(1):77–78, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1984:BRA

[354] R. Kay. Book reviews: Applied Statistics. A Handbook of Techniques. 5th ed, by Lothar Sachs. Applied Statistics, 33(1):78, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinse:1984:CRA

[355] Gregg E. Dinse and S. W. Lagakos. Correction: Regression analysis of tumour prevalence data. Applied Statistics, 33 (1):79–80, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Schervish:1984:SAA

[356] Mark J. Schervish. Statistical algorithms: Algorithm AS 195: Multivariate normal probabilities with error bound. Applied Statistics, 33(1):81– 94, March 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.

edu/apstat/195. See corrections [439].

Krailo:1984:SAA

[357] M. D. Krailo and M. C. Pike. Statistical algorithms: Algorithm AS 196: Conditional multivariate logistic analysis of stratified case-control studies. Applied Statistics, 33(1):95–103, March 1984. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/196.

Melard:1984:SAA

[358] G. Mélard. Statistical algorithms: Algorithm AS 197: A fast algorithm for the exact likelihood of autoregressivemoving average models. Applied Statistics, 33(1):104–114, March 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/197.

Pillers:1984:SAA

[359] Carolyn Pillers, Tim Robertson, and F. T. Wright. Statistical algorithms: Algorithm AS 198: The level probabilities of order restricted inference. Applied Statistics, 33(1):115–119, March 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/198.

Jones:1984:SAR

[360] M. C. Jones and H. W. Lotwick. Statistical algorithms: Remark AS R50: A remark on Algorithm AS 176. kernal density estimation using the fast Fourier transform. Applied Statistics, 33 (1):120–122, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/176. See [219, 530].

Krailo:1984:SAR

[361] Mark D. Krailo. Statistical algorithms:

Remark AS R51: A remark on AS 162. multivariate conditional logistic analysis of stratum-matched case-control studies. Applied Statistics, 33(1):123, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [161].

Wichmann:1984:SAC

[362] B. A. Wichmann and I. D. Hill. Statistical algorithms: Correction: Algorithm AS 183: An efficient and portable pseudo-random number generator. Applied Statistics, 33(1):123, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1984:BMPa

[363] Anonymous. Back matter. Applied Statistics, 33(1):[4], 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1984:FMPa

[364] Anonymous. Front matter. Applied Statistics, 33(1):[3], 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Henery:1984:EVM

[365] R. J. Henery. An extreme-value model for predicting the results of horse races. Applied Statistics, 33(2):125–133, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Do:1984:EMP

[366] K. Do and G. J. McLachlan. Estimation of mixing proportions: A case study. Applied Statistics, 33(2):134–140, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Doran:1984:EQF

[367] Howard E. Doran. Evaluating quadratic forms of the matrix (x′x)^{−1} in a regression analysis, with applications. Applied Statistics, 33(2):141–144, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1984:PCC

[368] A. F. Bissell. Performance of control charts and cusums under linear trend. Applied Statistics, 33(2):145–151, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [518].

Bissell:1984:ELT

[369] A. F. Bissell. Estimation of linear trend from a cusum chart or tabulation. Applied Statistics, 33(2):152–157, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ulrich:1984:CGD

[370] Gary Ulrich. Computer generation of distributions on the m-sphere. Applied Statistics, 33(2):158–163, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krzanowski:1984:PCA

[371] W. J. Krzanowski. Principal component analysis in the presence of group structure. Applied Statistics, 33(2):164–168, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1984:POM

[372] A. N. Pettitt. Proportional odds models for survival data and estimates using ranks. Applied Statistics, 33(2):169–175, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gore:1984:RMN

[373] Sheila M. Gore, Stuart J. Pocock, and Gillian R. Kerr. Regression models and non-proportional hazards in the analysis of breast cancer survival. Applied Statistics, 33(2):176–195, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harding:1984:EAP

[374] C. M. Harding, A. W. Heathwood, R. G. Hunt, and K. L. Q. Read. The estimation of animal population size by the removal method. Applied Statistics, 33 (2):196–202, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dempster:1984:SCA

[375] Arthur P. Dempster, Chandu M. Patel, Murray R. Selwyn, and Arthur J. Roth. Statistical and computational aspects of mixed model analysis. Applied Statistics, 33(2):203–214, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

vWKotze:1984:IOT

[376] T. J. v W. Kotze and Douglas M. Hawkins. The identification of outliers in two-way contingency tables using 2 × 2 subtables. Applied Statistics, 33(2):215–223, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Huffman:1984:MIA

[377] Michael D. Huffman. Miscellanea: An improved approximate two-sample Poisson test. Applied Statistics, 33(2): 224–226, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Nigam:1984:MMS

[378] A. K. Nigam and V. K. Gupta. Miscellanea: A method of sampling with equal or unequal probabilities without replacement. Applied Statistics, 33(2):227–229, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Korn:1984:LE

[379] E. L. Korn. Letter to the editors. Applied Statistics, 33(2):230–231, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1984:BRF

[380] A. C. Kimber. Book reviews: Fixed-Point Estimation in Theory and Practice, by Reinhold Bergstrom and Herman Wold. Applied Statistics, 33(2):232, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lee:1984:BRR

[381] Peter N. Lee. Book reviews: Regression Analysis of Survival Data in Cancer Chemotherapy, by Walter H. Carter, *Jr, Galen L. Wampler and Donald M. Stablein. Applied Statistics, 33 (2):232–233, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ashworth:1984:BRA

[382] J. Ashworth. Book reviews: Applications of Econometrics, by Julia Hebden. Applied Statistics, 33(2):233–234, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Conolly:1984:BRA

[383] B. W. Conolly. Book reviews: Applications of Queueing Theory, by G. F.

Newell. Applied Statistics, 33(2):234, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1984:BRA

[384] J. Taylor. Book reviews: The Agricultural Field Experiment: A Statistical Examination of Theory and Practice, by S. C. Pearce. Applied Statistics, 33 (2):234–235, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Collings:1984:BRM

[385] Stanley Collings. Book reviews: Matrix Algebra Useful for Statistics, by Shayle R. Searle. Applied Statistics, 33(2):235, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Roberts:1984:SAA

[386] Sarah J. Roberts. Statistical algorithms: Algorithm AS 199: A branch and bound algorithm for determining the optimal feature subset of given size. Applied Statistics, 33(2):236– 241, June 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/199. See remark [544].

Balakrishnan:1984:SAA

[387] N. Balakrishnan. Statistical algorithms: Algorithm AS 200: Approximating the sum of squares of normal scores. Applied Statistics, 33(2):242– 245, June 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/200.

Macdonald:1984:SAA

[388] Ranald R. Macdonald. Statistical algorithms: Algorithm AS 201: Combined significance test of differences between conditions and ordinal predictions. Applied Statistics, 33(2):245– 248, June 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/201.

Tanner:1984:SAA

[389] Martin A. Tanner. Statistical algorithms: Algorithm AS 202: Databased non-parametric hazard estimation. Applied Statistics, 33(2):248– 258, June 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/202.

Anonymous:1984:BMPb

[390] Anonymous. Back matter. Applied Statistics, 33(2):[4], 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1984:FMPb

[391] Anonymous. Front matter. Applied Statistics, 33(2):[3], 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ncube:1984:CSC

[392] Matoteng M. Ncube and William H.

Woodall. A combined Shewhartcumulative score quality control chart. Applied Statistics, 33(3):259–265, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barnard:1984:CMT

[393] G. A. Barnard. Comparing the means of two independent samples. Applied Statistics, 33(3):266–271, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Matthews:1984:RMA

[394] J. N. S. Matthews. Robust methods in the assessment of multivariate normality. Applied Statistics, 33(3):272–277, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Donner:1984:CST

[395] Allan Donner and Shelley Bull. A comparison of significance-testing procedures for parent-child correlations computed from family data. Applied Statistics, 33(3):278–284, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brooks:1984:ALR

[396] R. J. Brooks. Approximate likelihood ratio tests in the analysis of beta-binomial data. Applied Statistics, 33(3):285–289, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bardsley:1984:MEU

[397] P. Bardsley and R. L. Chambers. Multipurpose estimation from unbalanced samples. Applied Statistics, 33(3): 290–299, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1984:EVA

[398] Stephen J. Taylor. Estimating the variances of autocorrelations calculated from financial time series. Applied Statistics, 33(3):300–308, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cheng:1984:GIG

[399] Russell C. H. Cheng. Generation of inverse Gaussian variates with given sample mean and dispersion. Applied Statistics, 33(3):309–316, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morton:1984:LE

[400] R. Hugh Morton. Letters to the editors. Applied Statistics, 33(3):317–318, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Marriott:1984:BRM

[401] F. H. C. Marriott. Book reviews: Multivariate Statistics: A Vector Space Approach, by Morris L. Eaton. Applied Statistics, 33(3):319, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seheult:1984:BRP

[402] A. H. Seheult. Book reviews: A primer in data reduction: An introductory statistics textbook, by A. S. C. Ehrenberg. Applied Statistics, 33(3):319–320, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morris:1984:BRG

[403] R. W. Morris. Book reviews: GLIM 82, by R. Gilchrist. Applied Statistics, 33(3):320, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cass:1984:BRE

[404] J. M. Cass. Book reviews: Educated Guessing: How to Cope in an Uncertain World, by S. Kotz and D. F. Stroup. Applied Statistics, 33(3):321, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Neave:1984:BRU

[405] H. R. Neave. Book reviews: Understanding Randomness, by D. S. Salsburg. Applied Statistics, 33(3):321–322, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Loynes:1984:BRR

[406] R. M. Loynes. Book reviews: Recent Trends in Statistics, by S. Heiler. Applied Statistics, 33(3):323, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lioukas:1984:BRS

[407] S. Lioukas. Book reviews: Statistical Sources and Techniques, by F. J. Randall and M. D. Wolf. Applied Statistics, 33(3):323–324, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pridmore:1984:BRC

[408] W. A. Pridmore. Book reviews: Contributions to Statistics: Essays in Honour of Norman L. Johnson, by P. K. Sen. Applied Statistics, 33(3):324–325, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ingham:1984:BRA

[409] A. Ingham. Book reviews: Aspiration Levels in Bargaining and Economic Decision Making, by R. Tietz. Applied Statistics, 33(3):325–326, 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Taylor:1984:BRC

[410] S. M. Taylor. Book reviews: Capture-Recapture and Removal Methods for Sampling Closed Populations, by G. C. White, D. R. Anderson, K. P. Burnham and D. L. Otis. Applied Statistics, 33(3):326, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Agha:1984:SAA

[411] M. Agha and M. T. Ibrahim. Statistical algorithms: Algorithm AS 203: Maximum likelihood estimation of mixtures of distributions. Applied Statistics, 33(3):327–332, September 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/203.

Farebrother:1984:SAA

[412] R. W. Farebrother. Statistical algorithms: Algorithm AS 204: The distribution of a positive linear combination of χ^{2} random variables. Applied Statistics, 33(3):332–339, September 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/204.

Saunders:1984:SAA

[413] Ian W. Saunders. Statistical algorithms:

Algorithm AS 205: Enumeration of R × C tables with repeated row totals. Applied Statistics, 33(3):340–352, September 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/205.

Bril:1984:SAA

[414] Gordon Bril, Richard Dykstra, Carolyn Pillers, and Tim Robertson. Statistical algorithms: Algorithm AS 206: Isotonic regression in two independent variables. Applied Statistics, 33(3):352–357, September 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/206. See [543].

Haber:1984:SAA

[415] Michael Haber. Statistical algorithms: Algorithm AS 207: Fitting a general loglinear model. Applied Statistics, 33(3): 358–362, September 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/207.

Farebrother:1984:SARa

[416] R. W. Farebrother. Statistical algorithms: Remark AS R52: The distribution of a linear combination of central χ^{2} random variables: A remark on AS 153: Pan’s procedure for the tail probabilities of the Durbin-Watson statistic. Applied Statistics, 33(3):363–366, 1984. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [89, 417].

Farebrother:1984:SARb

[417] R. W. Farebrother. Statistical algorithms: Remark AS R53: A remark on algorithms AS 106, AS 153 and AS 155: The distribution of a linear combination of χ^{2} random variables. Applied Statistics, 33(3):366–369, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [14, 89, 112, 416].

Dewey:1984:SAR

[418] Michael E. Dewey. Statistical algorithms: Remark AS R54: A remark on algorithms AS 169: An improved algorithm for scatter plots. Applied Statistics, 33(3):370–372, 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [189, 270, 769].

Anonymous:1984:BMPc

[419] Anonymous. Back matter. Applied Statistics, 33(3):[6], 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1984:FMPc

[420] Anonymous. Front matter. Applied Statistics, 33(3):[3], 1984. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1984:VIP

[421] Anonymous. Volume information. Applied Statistics, 33(3):[2], 1984. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunsmore:1985:DPA

[422] Ian R. Dunsmore and David E. Wright.

A decisive predictive approach to the construction of sequential acceptance sampling plans for lifetimes. Applied Statistics, 34(1):1–13, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cruz-Orive:1985:SAA

[423] L.-M. Cruz-Orive, H. Hoppeler, O. Mathieu, and E. R. Weibel. Stereological analysis of anisotropic structures using directional statistics. Applied Statistics, 34(1):14–32, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stirling:1985:HMA

[424] W. Douglas Stirling. Heteroscedastic models and an application to block designs. Applied Statistics, 34(1):33–41, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walker:1985:SSP

[425] Alexander M. Walker. Small sample properties of some estimators of a common hazard ratio. Applied Statistics, 34(1):42–48, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Russell:1985:EMP

[426] N. S. Russell, D. R. Farrier, and J. Howell. Evaluation of multinormal probabilities using Fourier series expansions. Applied Statistics, 34(1):49–53, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lagakos:1985:RAP

[427] S. W. Lagakos and Louise M. Ryan. On the representativeness assumption in prevalence tests of carcinogenicity. Applied Statistics, 34(1):54–62, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Osmond:1985:BMA

[428] C. Osmond. Biplot models applied to cancer mortality rates. Applied Statistics, 34(1):63–70, 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Snee:1985:GDR

[429] Ronald D. Snee. Graphical display of results of three-treatment randomized block experiments. Applied Statistics, 34(1):71–77, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hole:1985:BRS

[430] David J. Hole. Book reviews: Statistical Methods for Cancer Studies, by R. G. Cornell. Applied Statistics, 34(1):78, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Darby:1985:BRE

[431] Keith V. Darby. Book reviews: Experimental Design, Statistical Models, and Genetic Statistics, by K. Hinkelmann. Applied Statistics, 34(1):79, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cunningham:1985:BRT

[432] V. J. Cunningham. Book reviews:

Tracer Kinetics and Physiologic Modeling. Theory to Practice, by R. M. Lambrecht and A. Rescigno. Applied Statistics, 34(1):79–80, 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bacon-Shone:1985:SAAa

[433] J. Bacon-Shone. Statistical algorithms:

Algorithm AS 208: Fitting a multivariate logistic normal distribution by the method of moments. Applied Statistics, 34(1):81–86, March 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/208.

Royston:1985:SAA

[434] J. P. Royston. Statistical algorithms: Algorithm AS 209: The distribution function of skewness and kurtosis. Applied Statistics, 34(1):87– 94, March 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/209.

Bacon-Shone:1985:SAAb

[435] J. Bacon-Shone. Statistical algorithms: Algorithm AS 210: Fitting five parameter Johnson S_{B} curves by moments. Applied Statistics, 34(1):95– 100, March 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/210.

Chou:1985:SARa

[436] Youn-Min Chou. Statistical algorithms: Remark AS R55: A remark on Algorithm AS 76: An integral useful in calculating noncentral T and bivariate normal probabilities. Applied Statistics, 34 (1):100–101, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [10, 27, 31, 542, 769].

Chou:1985:SARb

[437] Youn-Min Chou. Statistical algorithms: Remark AS R56: A remark on Algorithm AS 5: The integral of the noncentral t-distribution. Applied Statistics, 34(1):102, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/5. See [2].

Hill:1985:SAR

[438] I. D. Hill. Statistical algorithms: Remark AS R57: A remark on Algorithm AS 193: A revised algorithm for the spectral test. Applied Statistics, 34 (1):102–103, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/193. See [335].

Schervish:1985:SAC

[439] Mark J. Schervish. Statistical algorithms: Corrections: Algorithm AS 195: Multivariate normal probabilities with error bound. Applied Statistics, 34 (1):103–104, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [356].

Lund:1985:SAC

[440] R. E. Lund and J. R. Lund. Statistical algorithms: Corrections: Algorithm AS 190: Probabilities and upper quantiles for the Studentized range. Applied Statistics, 34(1):104, 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1985:BMPa

[441] Anonymous. Back matter. Applied Statistics, 34(1):[4], 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1985:FMPa

[442] Anonymous. Front matter. Applied Statistics, 34(1):[3], 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1985:CLM

[443] Byron J. T. Morgan. The cubic logistic model for quantal assay data. Applied Statistics, 34(2):105–113, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hoerl:1985:PUR

[444] Arthur E. Hoerl, Robert W. Kennard, and Roger W. Hoerl. Practical use of ridge regression: A challenge met. Applied Statistics, 34(2):114–120, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farewell:1985:SRA

[445] V. T. Farewell. Some remarks on the analysis of crossover trials with a binary response. Applied Statistics, 34 (2):121–128, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitchison:1985:KDE

[446] J. Aitchison and I. J. Lauder. Kernel density estimation for compositional data. Applied Statistics, 34(2):129–137, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Diggle:1985:KMS

[447] Peter Diggle. A kernel method for smoothing point process data. Applied Statistics, 34(2):138–147, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clayton:1985:EAC

[448] David Clayton and Jack Cuzick. The EM algorithm for Cox’s regression model using GLIM. Applied Statistics, 34(2):148–156, 1985. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Neter:1985:RBB

[449] John Neter and James Godfrey. Robust Bayesian bounds for monetary-unit sampling in auditing. Applied Statistics, 34 (2):157–168, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1985:MND

[450] A. C. Kimber. Miscellanea: A note on the detection and accommodation of outliers relative to Fisher’s distribution on the sphere. Applied Statistics, 34 (2):169–172, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Janacek:1985:BRT

[451] G. J. Janacek. Book reviews: Time Series in the Frequency Domain, by D. R. Brillinger and P. R. Krishnaiah. Applied Statistics, 34(2):173, 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Latham:1985:BRS

[452] F. R. Latham. Book reviews: Statistical Methods in Research and Production with Special Reference to the Chemical Industry, by Owen L. Davies and Peter L. Goldsmith. Applied Statistics, 34 (2):173–174, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkins:1985:BRS

[453] Anne S. Hawkins. Book reviews: Statistics and Probability — An Introductory Course, by School Statistics Panel. Applied Statistics, 34(2):174, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Henery:1985:BRB

[454] R. J. Henery. Book reviews: Beat the Racetrack, by William T. Ziemba and Donald B. Hausch. Applied Statistics, 34(2):175, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pridmore:1985:BRA

[455] W. A. Pridmore. Book reviews: Applied Business Statistics. 2nd Edition, by Fadil H. Zuwaylif. Applied Statistics, 34(2):175–176, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Flury:1985:SAA

[456] Bernard N. Flury and Gregory Constantine. Statistical algorithms: Algorithm AS 211: The F -G diagonalization algorithm. Applied Statistics, 34(2):177– 183, June 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/211. See [635, 663].

Stirling:1985:SAAa

[457] W. Douglas Stirling. Statistical algorithms: Algorithm AS 212: Fitting the exponential curve by least squares. Applied Statistics, 34(2):183– 192, June 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/212.

Lin:1985:SAA

[458] Shang P. Lin and Robert B. Bendel. Statistical algorithms: Algorithm AS 213: Generation of population correlation matrices with specified eigenvalues. Applied Statistics, 34(2):193– 198, June 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/213.

McLeod:1985:SAR

[459] A. Ian McLeod. Statistical algorithms: Remark AS R58: A remark on Algorithm AS 183. an efficient and portable pseudo-random number generator. Applied Statistics, 34 (2):198–200, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/183. See [241, 506].

Anonymous:1985:BMPb

[460] Anonymous. Back matter. Applied Statistics, 34(2):[6], 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1985:FMPb

[461] Anonymous. Front matter. Applied Statistics, 34(2):[3], 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Larson:1985:MMR

[462] Martin G. Larson and Gregg E. Dinse.

A mixture model for the regression analysis of competing risks data. Applied Statistics, 34(3):201–211, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moreau:1985:GGF

[463] T. Moreau, J. O’Quigley, and M. Mesbah. A global goodness-of-fit statistic for the proportional hazards model. Applied Statistics, 34(3):212–218, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krzanowski:1985:UAM

[464] W. J. Krzanowski, R. Mead, and S. Thorne. On the use of average marks in degree examination data. Applied Statistics, 34(3):219–226, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lewis:1985:PCD

[465] S. M. Lewis and M. G. Tuck. Paired comparison designs for factorial experiments. Applied Statistics, 34(3): 227–234, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Campbell:1985:UFA

[466] N. A. Campbell. Updating formulae for allocation of individuals. Applied Statistics, 34(3):235–236, 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berman:1985:IAI

[467] Mark Berman and David Griffiths. Incorporating angular information into models for stone circle data. Applied Statistics, 34(3):237–245, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Barlow:1985:GRR

[468] William E. Barlow. General relative risk models in stratified epidemiologic studies. Applied Statistics, 34(3):246–257, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kelly:1985:USE

[469] Gabrielle E. Kelly. Use of the structural equations model in assessing the reliability of a new measurement technique. Applied Statistics, 34(3):258–263, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tam:1985:LFE

[470] S. M. Tam. On labour force estimators. Applied Statistics, 34(3):264–272, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hurley:1985:AGL

[471] M. A. Hurley. An application of generalized linear models to survival analysis with two types of failure. Applied Statistics, 34(3):273–281, 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Basford:1985:LEN

[472] K. E. Basford and G. J. McLachlan.

Likelihood estimation with normal mixture models. Applied Statistics, 34 (3):282–289, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Atherton:1985:BRM

[473] Judith Atherton. Book reviews: A Manager’s Guide to Quantitative Methods, by Michael Cuming. Applied Statistics, 34(3):290, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cranley:1985:BRM

[474] R. Cranley. Book reviews: Multivariate Analysis-Methods and Applications, by William R. Dilon and Matthew Goldstein. Applied Statistics, 34(3):290–291, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

House:1985:BRL

[475] Frank R. House. Book reviews: Linear Statistical Models and Related Methods, by J. Fox. Applied Statistics, 34 (3):291–292, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walker:1985:BRE

[476] K. Walker. Book reviews: Ensemble Modeling. Inference from Small-Scale Properties to Large-Scale Systems, by Alan E. Gelfand and Crayton C. Walker. Applied Statistics, 34(3):292–293, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Daldry:1985:BRL

[477] S. J. Daldry. Book reviews: Loss Distributions, by Robert V. Hogg and Stuart A. Klugman. Applied Statistics, 34 (3):293–294, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paterson:1985:BRA

[478] L. Paterson. Book reviews: Applied Factorial and Fractional Designs, by Robert A. McLean and Virgil L. Anderson. Applied Statistics, 34(3):294, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1985:LE

[479] M. C. Jones. Letter to the editors. Applied Statistics, 34(3):295, 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Buckland:1985:SAA

[480] S. T. Buckland. Statistical algorithms: Algorithm AS 214: Calculation of Monte Carlo confidence intervals. Applied Statistics, 34(3):296–301, September 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/214.

Hosking:1985:SAA

[481] J. R. M. Hosking. Statistical algorithms: Algorithm AS 215: Maximumlikelihood estimation of the parameters of the generalized extreme-value distribution. Applied Statistics, 34(3):301– 310, September 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/215. See remark

[711].

Stirling:1985:SAAb

[482] W. Douglas Stirling. Statistical algorithms: Algorithm AS 216: Fitting models with a linear part and auxiliary parameters. Applied Statistics, 34(3):310– 319, September 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/216.

Hartigan:1985:SAA

[483] P. M. Hartigan. Statistical algorithms:

Algorithm AS 217: Computation of the dip statistic to test for unimodality. Applied Statistics, 34(3):320–325, September 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/217.

Rice:1985:SAR

[484] Lory R. Rice and Rose E. Gaines Das. Statistical algorithms: A remark on algorithms AS 32 and AS 147. the incomplete gamma integral. Applied Statistics, 34(3):326, 1985. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [7, 63].

Anonymous:1985:BMPc

[485] Anonymous. Back matter. Applied Statistics, 34(3):[4], 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1985:FMPc

[486] Anonymous. Front matter. Applied Statistics, 34(3):[3], 1985. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1985:VIP

[487] Anonymous. Volume information. Applied Statistics, 34(3):[2], 1985. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cain:1986:OMC

[488] M. Cain and M. L. R. Price. Optimal mixture choice. Applied Statistics, 35(1):1–7, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whitmore:1986:IGR

[489] G. A. Whitmore. Inverse Gaussian ratio estimation. Applied Statistics, 35(1):8– 15, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kay:1986:AFL

[490] Richard Kay and Sarah Little. Assessing the fit of the logistic model: A case study of children with the haemolytic uraemic syndrome. Applied Statistics, 35(1):16– 30, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McClean:1986:EME

[491] S. McClean. Estimation for the mixed exponential distribution using grouped follow-up data. Applied Statistics, 35 (1):31–37, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Beach:1986:LCI

[492] C. M. Beach and S. F. Kaliski. Lorenz curve inference with sample weights: An application to the distribution of unemployment experience. Applied Statistics, 35(1):38–45, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cox:1986:PHS

[493] T. Cox and P. H. Gibson. The probability of hitting a synapse: A stereological problem in neuropathology. Applied Statistics, 35(1):46–48, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wright:1986:NCH

[494] D. E. Wright. A note on the construction of highest posterior density intervals. Applied Statistics, 35(1):49–53, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Berman:1986:TSA

[495] Mark Berman. Testing for spatial association between a point process and another stochastic process. Applied Statistics, 35(1):54–62, 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Johnson:1986:BRE

[496] Wesley Johnson, Jessica Utts, and Larry Pearson. Bayesian robust estimation of the mean. Applied Statistics, 35(1):63– 72, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1986:MOR

[497] A. F. Bissell. Miscellanea: Ordered random selection without replacement. Applied Statistics, 35(1):73–75, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Veevers:1986:BRR

[498] A. Veevers. Book reviews: Repairable Systems Reliability: Modeling, Inference, Misconceptions and Their Causes, by Harold Ascher and Harry Feingold. Applied Statistics, 35(1):76, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Brooks:1986:BRB

[499] R. J. Brooks. Book reviews: Bayesian Analysis of Linear Models, by Lyle D. Broemeling. Applied Statistics, 35 (1):76–77, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hunter:1986:BRA

[500] E. A. Hunter. Book reviews: Applied Regression Analysis and Experimental Design, by R. J. Brook and G. C. Arnold. Applied Statistics, 35(1):77–78, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sykes:1986:BRS

[501] Alan M. Sykes. Book reviews:

Self-Organising Methods in Modelling. GMDH Type Algorithms, by S. J. Farlow. Applied Statistics, 35(1):78–79, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Paige:1986:BRA

[502] W. S. Paige. Book reviews: Applied Statistics Algorithms, by P. Griffiths and I. D. Hill. Applied Statistics, 35(1):79, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Escobar:1986:SAA

[503] Luis A. Escobar and William Q. Meeker, Jr. Statistical algorithms: Algorithm AS 218: Elements of the Fisher information matrix for the smallest extreme value distribution and censored data. Applied Statistics, 35(1):80–86, March 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/218.

Dewey:1986:SAR

[504] M. E. Dewey. Statistical algorithms: Remark ASR 60: A remark on AS 112 exact distributions derived from two-way tables. Applied Statistics, 35(1):86–88, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [17, 139].

Dewey:1986:SAC

[505] M. E. Dewey. Statistical algorithms:

Correction: Algorithm AS 205: Enumeration of R × C tables with repeated row totals. Applied Statistics, 35(1):88, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zeisel:1986:SAR

[506] H. Zeisel. Statistical algorithms: Remark ASR 61: A remark on Algorithm AS 183. an efficient and portable pseudo-random number generator. Applied Statistics, 35(1):89, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

                                        /lib.stat.cmu.edu/apstat/183. See

[241, 459].

Paul:1986:SAR

[507] S. R. Paul and K. Yuen Fung. Statistical algorithms: Remark ASR 62: A remark on Algorithm AS 148: The jackknife. Applied Statistics, 35(1): 89–91, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/148. See [64].

Anonymous:1986:BMPa

[508] Anonymous. Back matter. Applied Statistics, 35(1):[2], 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1986:FMPa

[509] Anonymous. Front matter. Applied Statistics, 35(1):[3], 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Racine:1986:BMP

[510] A. Racine, A. P. Grieve, H. Flühler, and A. F. M. Smith. Bayesian methods in practice: Experiences in the pharmaceutical industry. Applied Statistics, 35 (2):93–150, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). With discussion.

Waldmann:1986:BDR

[511] K.-H. Waldmann. Bounds for the distribution of the run length of geometric moving average charts. Applied Statistics, 35(2):151–158, 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rubin:1986:ESC

[512] D. B. Rubin and N. Schenker. Efficiently simulating the coverage properties of interval estimates. Applied Statistics, 35 (2):159–167, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Broersen:1986:SRS

[513] P. M. T. Broersen. Subset regression with stepwise directed search. Applied Statistics, 35(2):168–177, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bairagi:1986:CVE

[514] R. Bairagi. On components of variation of estimated weight velocity of children. Applied Statistics, 35(2):178–182, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gilks:1986:RTS

[515] W. R. Gilks. Rapid two-stage modelling technique for exploring large data sets. Applied Statistics, 35(2):183–194, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Naes:1986:CLS

[516] T. Naes, C. Irgens, and H. Martens. Comparison of linear statistical methods for calibration of nir instruments. Applied Statistics, 35(2):195–206, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilson:1986:CSS

[517] S. R. Wilson and I. Gordon. Calculating sample sizes in the presence of confounding variables. Applied Statistics, 35 (2):207–213, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1986:CPC

[518] A. F. Bissell. Corrigendum: The performance of control charts and cusums under linear trend. Applied Statistics, 35(2):214, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [368].

Green:1986:BRP

[519] M. W. Green. Book reviews: Pareto Distributions, by B. C. Arnold. Applied Statistics, 35(2):215, 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moore:1986:BRN

[520] B. A. Moore. Book reviews: Nonparametric Statistical Inference. 2nd Edition, by Jean Dickinson Gibbons. Applied Statistics, 35(2):215–216, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chapman:1986:BRE

[521] Paul G. Chapman. Book reviews: Economic Statistics 1900–1983: United Kingdom, United States of America, France, Germany, Italy, Japan, by Thelma Liesner. Applied Statistics, 35(2):216, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bunday:1986:BRB

[522] B. D. Bunday. Book reviews: Basic Operational Research. 3rd Edition, by P. B. Moore. Applied Statistics, 35(2):217, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Skinner:1986:BRA

[523] C. J. Skinner. Book reviews: Analysing Change-Measurement and Explanation Using Longitudinal Data, by Ian Plewis. Applied Statistics, 35(2):217–218, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davies:1986:BRG

[524] Paul Davies. Book reviews: Genesis: An Authorship Study in Computer-Assisted Statistical Linguistics, by Y. T. Radday. Applied Statistics, 35(2):218, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hawkins:1986:BRE

[525] A. S. Hawkins. Book reviews: Elementary Statistics, by D. H. Skane. Applied Statistics, 35(2):218–219, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jowett:1986:BRD

[526] E. Jowett. Book reviews: Directory of Statistical Microcomputer Software, by W. A. Woodward, A. C. Elliott and H. L. Gray. Applied Statistics, 35(2):219, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dewey:1986:SAA

[527] M. E. Dewey. Statistical algorithms: Algorithm AS 219: Height balanced trees. Applied Statistics, 35(2):220– 226, June 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/219.

Bohrer:1986:SAA

[528] Robert Bohrer. Statistical algorithms: Algorithm AS 220: Operating characteristics of James–Stein and Efron-Morris estimators. Applied Statistics, 35(2): 226–232, June 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/220.

Royston:1986:SAR

[529] J. P. Royston. Statistical algorithms: Remark ASR 63: A remark on AS 181. the W test for normality. Applied Statistics, 35(2):232–234, 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [239].

Schiffelbein:1986:SAR

[530] Paul Schiffelbein. Statistical algorithms: Remark ASR 64: A remark on Algorithm AS 176: Kernel density estimation using the fast Fourier transform. Applied Statistics, 35(2):235–236, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/176. See

[219, 360].

Anonymous:1986:BMPb

[531] Anonymous. Back matter. Applied Statistics, 35(2):[4], 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1986:FMPb

[532] Anonymous. Front matter. Applied Statistics, 35(2):[3], 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ehrenberg:1986:RTE

[533] A. S. C. Ehrenberg. Reading a table: An example. Applied Statistics, 35 (3):237–244, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mau:1986:GMD

[534] J. Mau. On a graphical method for the detection of time-dependent effects of covariates in survival data. Applied Statistics, 35(3):245–255, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lwin:1986:CWS

[535] Thaung Lwin and Clifford H. Spiegelman. Calibration with working standards. Applied Statistics, 35(3):256–261, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Seeger:1986:DAE

[536] Paul Seeger. Design and analysis of experiments with sugar beet seeds. Applied Statistics, 35(3):262–268, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dansie:1986:NOS

[537] B. R. Dansie. Normal order statistics as permutation probability models. Applied Statistics, 35(3):269–275, 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dunstan:1986:MBC

[538] R. Dunstan and R. L. Chambers. Modelbased confidence intervals in multipurpose surveys. Applied Statistics, 35 (3):276–280, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Petersen:1986:FPS

[539] Trond Petersen. Fitting parametric survival models with time-dependent covariates. Applied Statistics, 35(3): 281–288, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ashby:1986:OPR

[540] Deborah Ashby, Stuart J. Pocock, and A. Gerald Shaper. Ordered polytomous regression: An example relating serum biochemistry and haematology to alcohol consumption. Applied Statistics, 35 (3):289–301, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

DerSimonian:1986:SAA

[541] R. DerSimonian. Statistical algorithms: Algorithm AS 221: Maximum likelihood estimation of a mixing distribution. Applied Statistics, 35(3):302–309, September 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/221.

Thomas:1986:SAR

[542] G. E. Thomas. Statistical algorithms:

Remark ASR 65: A remark on Algorithm AS 76: An integral useful in calculating non-central t and bivariate normal probabilities. Applied Statistics, 35 (3):310–312, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [10, 27, 31, 436, 769].

Murray:1986:SAR

[543] Gordon D. Murray and Alison J. Wilson.

Statistical algorithms: Remark ASR 66: Zero weighted cells. A remark on AS 206: Isotonic regression in two independent variates. Applied Statistics, 35 (3):312–314, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [414].

Cobby:1986:SAR

[544] John M. Cobby. Statistical algorithms:

Remark ASR 67: A remark on AS 199. A branch and bound algorithm for determining the optimal feature subset of given size. Applied Statistics, 35(3):314, 1986. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [386].

Anonymous:1986:BMPc

[545] Anonymous. Back matter. Applied Statistics, 35(3):[4], 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1986:FMPc

[546] Anonymous. Front matter. Applied Statistics, 35(3):[3], 1986. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1986:VIP

[547] Anonymous. Volume information. Applied Statistics, 35(3):[2], 1986. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1987:SVP

[548] D. J. Hand. Screening vs prevalence estimation. Applied Statistics, 36(1):1–7, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Moore:1987:MEV

[549] Dirk F. Moore. Modelling the extraneous variance in the presence of extrabinomial variation. Applied Statistics, 36(1):8–14, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Perry:1987:IIP

[550] Joe N. Perry. Iterative improvement of a power transformation to stabilise variance. Applied Statistics, 36(1):15–21, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Krzanowski:1987:SVP

[551] W. J. Krzanowski. Selection of variables to preserve multivariate data structure, using principal components. Applied Statistics, 36(1):22–33, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jupp:1987:FSP

[552] Peter E. Jupp and John T. Kent. Fitting smooth paths to spherical data. Applied Statistics, 36(1):34–46, 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cameron:1987:FMS

[553] Murray A. Cameron and T. Rolf Turner.

Fitting models to spectra using regression packages. Applied Statistics, 36 (1):47–57, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cowan:1987:HEG

[554] Richard Cowan. Hand evaluation in the game of contract bridge. Applied Statistics, 36(1):58–71, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wahrendorf:1987:BCN

[555] Jurgen Wahrendorf, Heiko Becher, and Charles C. Brown. Bootstrap comparison of non-nested generalized linear models: Applications in survival analysis and epidemiology. Applied Statistics, 36(1):72–81, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chang:1987:GTT

[556] Wei-Chien Chang. A graph for two training samples in a discriminant analysis. Applied Statistics, 36(1):82–91, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coleman:1987:MPH

[557] S. Y. Coleman and T. F. Cox. Miscellanea: The probability of hitting a synapse: An addendum. Applied Statistics, 36(1):92–93, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coleman:1987:PHS

[558] S. Y. Coleman and T. F. Cox. The probability of hitting a synapse: an addendum. Applied Statistics, 36(1):92–93, 1987.

Royston:1987:MCD

[559] J. P. Royston, J. B. Webb, P. Griffiths, and I. D. Hill. Miscellanea: The construction and description of algorithms. Applied Statistics, 36(1):94–103, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hardle:1987:SAA

[560] W. Härdle. Statistical algorithms: Algorithm AS 222: Resistant smoothing using the fast Fourier transform. Applied Statistics, 36(1):104–111, March 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/222. See [662].

Lee:1987:SAA

[561] Tze-San Lee. Statistical algorithms: Algorithm AS 223: Optimum ridge parameter selection. Applied Statistics, 36 (1):112–118, March 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/223.

Farebrother:1987:SAR

[562] R. W. Farebrother. Statistical algorithms: Remark AS R68: Constrained and unconstrained L_{1} estimation A remark on AS 108: Multiple linear regression with minimum sum of absolute errors. Applied Statistics, 36(1):118, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [15, 26].

Royston:1987:SAR

[563] J. P. Royston. Statistical algorithms:

Remark AS R69: A remark on Algorithm AS 190: Probabilities and upper quantiles for the Studentized range. Applied Statistics, 36(1):119, 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/190. See

[308].

Hill:1987:SAR

[564] I. D. Hill. Statistical algorithms: Remark AS R70: A remark on Algorithm AS 16. maximum likelihood estimation from grouped and censored normal data. Applied Statistics, 36 (1):119–120, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/16. See [5, 16].

Murray:1987:SAC

[565] Gordon D. Murray and Alison J. Wilson. Statistical algorithms: Correction: Algorithm AS 206: Istonic regression in two independent variates. Applied Statistics, 36(1):120, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1987:BMPa

[566] Anonymous. Back matter. Applied Statistics, 36(1):[4], 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1987:FMPa

[567] Anonymous. Front matter. Applied Statistics, 36(1):[3], 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Finkelstein:1987:ECP

[568] Dianne M. Finkelstein and Louise M. Ryan. Estimating carcinogenic potency from a rodent tumorigenicity experiment. Applied Statistics, 36(2):121–133, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1987:RCP

[569] M. C. Jones. Randomly choosing parameters from the stationarity and invertibility region of autoregressive-moving average models. Applied Statistics, 36 (2):134–138, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kent:1987:AEB

[570] John T. Kent. Asymptotic expansions for the Bingham distribution. Applied Statistics, 36(2):139–144, 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ridout:1987:RSS

[571] M. S. Ridout and J. M. Cobby. Ranked set sampling with nonrandom selection of sets and errors in ranking. Applied Statistics, 36(2):145–152, 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bacon-Shone:1987:NGM

[572] J. Bacon-Shone and W. K. Fung. A new graphical method for detecting single and multiple outliers in univariate and multivariate data. Applied Statistics, 36(2):153–162, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Strauss:1987:RPR

[573] David Strauss and Barry C. Arnold.

The rating of players in racquetball tournaments. Applied Statistics, 36 (2):163–173, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Evett:1987:ELR

[574] I. W. Evett, P. E. Cage, and C. G. G.

Aitken. Evaluation of the likelihood ratio for fibre transfer evidence in criminal cases. Applied Statistics, 36(2):174–180, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1987:GLM

[575] D. A. Williams. Generalized linear model diagnostics using the deviance and single case deletions. Applied Statistics, 36(2):181–191, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kenward:1987:LLM

[576] Michael G. Kenward and Byron Jones.

A log-linear model for binary cross-over data. Applied Statistics, 36(2):192–204, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shirley:1987:ARM

[577] Eryl A. C. Shirley. Applications of ranking methods to multiple comparison procedures and factorial experiments. Applied Statistics, 36(2):205–213, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Masarotto:1987:RIA

[578] G. Masarotto. Robust identification of autoregressive moving average models.

Applied Statistics, 36(2):214–220, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawrance:1987:MNV

[579] A. J. Lawrance. Miscellanea: A note on the variance of the Box–Cox regression transformation estimate. Applied Statistics, 36(2):221–223, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Altman:1987:LE

[580] Douglas G. Altman, J. Martin Bland, and Gabrielle E. Kelly. Letters to the editors. Applied Statistics, 36(2):224–227, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dean:1987:SAA

[581] A. M. Dean. Statistical algorithms: Algorithm AS 224: Combining component designs to form a design with several orthogonal blocking factors. Applied Statistics, 36(2):228–234, June 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/224.

Wollan:1987:SAA

[582] Peter C. Wollan and Richard L. Dykstra. Statistical algorithms: Algorithm AS 225: Minimizing linear inequality constrained Mahalanobis distances. Applied Statistics, 36(2):234– 240, June 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/225.

Lenth:1987:SAA

[583] Russell V. Lenth. Statistical algorithms:

Algorithm AS 226: Computing noncentral beta probabilities. Applied Statistics, 36(2):241–244, June 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/226. See

remarks [774, 775].

Marsh:1987:SAA

[584] N. W. A. Marsh. Statistical algorithms:

Algorithm AS 227: Efficient generation of all binary patterns by Gray code counting. Applied Statistics, 36(2):245– 249, June 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/227.

Anonymous:1987:BMPb

[585] Anonymous. Back matter. Applied Statistics, 36(2):[5], 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1987:FMPb

[586] Anonymous. Front matter. Applied Statistics, 36(2):[3], 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Glasbey:1987:TDF

[587] C. A. Glasbey. Tolerance-distribution-free analyses of quantal dose-response data. Applied Statistics, 36(3):251–259, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hastie:1987:NPL

[588] Trevor Hastie and Robert Tibshirani.

Non-parametric logistic and proportional odds regression. Applied Statistics, 36(3):260–276, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kemp:1987:RGF

[589] C. D. Kemp and Adrienne W. Kemp. Rapid generation of frequency tables. Applied Statistics, 36(3):277–282, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1987:MMC

[590] Richard L. Smith and Mark Corbett. Measuring marathon courses: an application of statistical calibration theory. Applied Statistics, 36(3):283–295, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kenward:1987:MCP

[591] Michael G. Kenward. Method for comparing profiles of repeated measurements. Applied Statistics, 36(3): 296–308, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hakulinen:1987:RAR

[592] T. Hakulinen and L. Tenkanen. Regression analysis of relative survival rates. Applied Statistics, 36(3):309–317, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

McLachlan:1987:BLR

[593] G. J. McLachlan. On bootstrapping the likelihood ratio test statistic for the number of components in a normal mixture. Applied Statistics, 36(3):318–324, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prentice:1987:FSP

[594] Michael J. Prentice. Fitting smooth paths to rotation data. Applied Statistics, 36(3):325–331, 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Aitkin:1987:MVH

[595] Murray Aitkin. Modelling variance heterogeneity in normal regression using Glim. Applied Statistics, 36(3):332–339, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Gaynor:1987:UTD

[596] Jeffrey J. Gaynor. The use of time dependent covariates in modelling data from an occupational cohort study. Applied Statistics, 36(3):340–351, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kimber:1987:ATT

[597] A. C. Kimber and C. Jeynes. Application of the truncated two-piece normal distribution to the measurement of depths of arsenic implants in silicon. Applied Statistics, 36(3):352–357, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1987:CML

[598] Richard L. Smith and J. C. Naylor.

A comparison of maximum likelihood and Bayesian estimators for the threeparameter Weibull distribution. Applied Statistics, 36(3):358–369, 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pinkham:1987:EAD

[599] R. S. Pinkham. An efficient algorithm for drawing a simple random sample. Applied Statistics, 36(3):370–372, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Joliffe:1987:LE

[600] I. T. Joliffe. Letter to the editors. Applied Statistics, 36(3):373–374, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stone:1987:BRS

[601] M. Stone. Book reviews: The Statistical Analysis of Compositional Data, by John Aitchison. Applied Statistics, 36(3):375, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lovie:1987:BRE

[602] Pat Lovie. Book reviews: The Elements of Graphing Data, by William S. Cleveland. Applied Statistics, 36(3):375–376, 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davy:1987:BRI

[603] P. J. Davy. Book reviews: Introductory Probability Theory, by Janos Galambos. Applied Statistics, 36(3):376, 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dykstra:1987:SAA

[604] Richard L. Dykstra and Peter C. Wollan. Statistical algorithms: Algorithm AS 228: Finding I-projections subject to a finite set of linear inequality constraints. Applied Statistics, 36(3):377– 383, September 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/228.

Koenker:1987:SAA

[605] Roger W. Koenker and Vasco D’Orey.

Statistical algorithms: Algorithm AS 229: Computing regression quantiles. Applied Statistics, 36(3):383–393, September 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/229.

Smith:1987:SAA

[606] V. L. Smith. Statistical algorithms: Algorithm AS 230: Distribution of customers in M/E_{T} /m queues using Hokstad’s approximation. Applied Statistics, 36(3):394–401, September 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/230.

Farebrother:1987:SAA

[607] R. W. Farebrother. Statistical algorithms: Algorithm AS 231: The distribution of a noncentral X^{2} variable with nonnegative degrees of freedom. Applied Statistics, 36(3):402–405, September 1987. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/231.

Anonymous:1987:BMPc

[608] Anonymous. Back matter. Applied Statistics, 36(3):[4], 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1987:FMPc

[609] Anonymous. Front matter. Applied Statistics, 36(3):[3], 1987. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1987:VIP

[610] Anonymous. Volume information. Applied Statistics, 36(3):[4], 1987. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arjas:1988:TDB

[611] Elja Arjas and David Venzon. A test for discriminating between additive and multiplicative relative risks in survival analysis. Applied Statistics, 37(1):1–11, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Raftery:1988:ASD

[612] Adrian E. Raftery. Analysis of a simple debugging model. Applied Statistics, 37(1):12–22, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Little:1988:REM

[613] Roderick J. A. Little. Robust estimation of the mean and covariance matrix from data with missing values. Applied Statistics, 37(1):23–38, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Giltinan:1988:DTS

[614] D. M. Giltinan, T. P. Capizzi, and H. Malani. Diagnostic tests for similar action of two compounds. Applied Statistics, 37(1):39–50, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Campbell:1988:STI

[615] N. A. Campbell and H. T. Kiiveri.

Spectral-temporal indices for discrimination. Applied Statistics, 37(1):51–62, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Riddle:1988:PMT

[616] Lawrence H. Riddle. Probability models for tennis scoring systems. Applied Statistics, 37(1):63–75, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See corrigendum [691].

Puterman:1988:LIA

[617] Martin L. Puterman. Leverage and influence in autocorrelated regression models. Applied Statistics, 37(1):76–86, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Venzon:1988:MCP

[618] D. J. Venzon and S. H. Moolgavkar. Method for computing profilelikelihood-based confidence intervals. Applied Statistics, 37(1):87–94, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Haines:1988:CFA

[619] Paul D. Haines. A closed form approximation for calculating the percentage points of the F and t distributions. Applied Statistics, 37(1):95–100, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fligner:1988:NTJ

[620] M. A. Fligner and J. S. Verducci. A nonparametric test for judges’ bias in an athletic competition. Applied Statistics, 37(1):101–110, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fidler:1988:LE

[621] V. Fidler, M. G. Kenward, and B. Jones. Letters to the editors. Applied Statistics, 37(1):111–113, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hills:1988:BRV

[622] Mary G. Hills. Book reviews: Validity in Survey Research, by W. A. Belson. Applied Statistics, 37(1):114, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Breeze:1988:BRG

[623] Peter Breeze. Book reviews: Goodnessof-Fit Techniques, by R. B. D’Agostino and M. A. Stephens. Applied Statistics, 37(1):115, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mullins:1988:BRP

[624] Eamonn Mullins. Book reviews: Probability and Statistics. 2nd edn, by M. H. DeGroot. Applied Statistics, 37 (1):116–117, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

MacRae:1988:BRC

[625] K. D. MacRae. Book reviews: Clinical Epidemiology: The Architecture of Clinical Research, by Alvan R. Feinstein. Applied Statistics, 37(1):117, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lawrence:1988:BRE

[626] Clive J. Lawrence. Book reviews: Encyclopedia of Statistical Sciences. Volume 7: Plackett Family of Distributions-Regression, Wrong, by Samuel Kotz, Norman L. Johnson and Campbell B. Read. Applied Statistics, 37(1):118, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kiountouzis:1988:BRC

[627] E. Kiountouzis and S. Lioukas. Book reviews: Computer Modeling for Business and Industry, by Bruce Bowerman and Richard O’Connell. Applied Statistics, 37(1):118–119, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Steer:1988:BRT

[628] C. D. Steer. Book reviews: Truncated and Censored Samples for Normal Populations, by Helmut Schnieder. Applied Statistics, 37(1):119–120, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Green:1988:BRD

[629] P. J. Green and A. H. Seheult. Book reviews: Density Estimation for Statistics and Data Analysis, by B. W. Silverman. Applied Statistics, 37(1):120–121, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1988:BRR

[630] A. N. Pettitt. Book reviews: Robust Inference, by M. L. Tiku, W. Y. Tan and N. Balakrishnan. Applied Statistics, 37 (1):121–122, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fochtmann:1988:BRQ

[631] J. A. Fochtmann. Book reviews: Quality Management Handbook, by Loren Walsh, Ralph Wurster and Raymond J. Kimber. Applied Statistics, 37(1): 122–123, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hinde:1988:SSR

[632] John Hinde. Statistical software reviews: MacSpin. Version 1.1. Applied Statistics, 37(1):124–126, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pate:1988:SAA

[633] M. B. Pate and N. Davies. Statistical algorithms: Algorithm AS 232: Computation of population and sample correlation and partial correlation matrices in MARMA(P,Q) time series. Applied Statistics, 37(1):127– 138, March 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/232.

Ridout:1988:SAA

[634] Martin S. Ridout. Statistical algorithms: Algorithm AS 233: An improved branch and bound algorithm for feature subset selection. Applied Statistics, 37(1):139– 147, March 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/233.

Clarkson:1988:SARa

[635] Douglas B. Clarkson. Statistical algorithms: Remark AS R71: A remark on Algorithm AS 211. the F -G diagonalization algorithm. Applied Statistics, 37 (1):147–151, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/211. See [456, 663].

Shea:1988:SAR

[636] B. L. Shea and A. J. Scallon. Statistical algorithms: Remark AS R72: A remark on Algorithm AS 128. approximating the covariance matrix of normal order statistics. Applied Statistics, 37 (1):151–155, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/128. See [23].

Anonymous:1988:BMPa

[637] Anonymous. Back matter. Applied Statistics, 37(1):[5], 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1988:FMPa

[638] Anonymous. Front matter. Applied Statistics, 37(1):[3], 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cheuvart:1988:NMM

[639] B. Cheuvart. A nonparametric model for multiple recurrences. Applied Statistics, 37(2):157–168, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lagakos:1988:UTL

[640] Stephen W. Lagakos and Thomas A.

Louis. Use of tumour lethality to interpret tumorigenicity experiments lacking cause-of-death data. Applied Statistics, 37(2):169–179, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Linnet:1988:TNT

[641] Kristian Linnet. Testing normality of transformed data. Applied Statistics, 37 (2):180–186, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Koehler:1988:CAS

[642] Anne B. Koehler and Emily S. Murphree. A comparison of the Akaike and Schwarz criteria for selecting model order. Applied Statistics, 37(2):187–195, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Im:1988:MMB

[643] S. Im and D. Gianola. Mixed models for binomial data with an application to lamb mortality. Applied Statistics, 37 (2):196–204, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1988:ANP

[644] N. J. Hill, A. R. Padmanabhan, and Madan L. Puri. Adaptive nonparametric procedures and applications. Applied Statistics, 37(2):205–218, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hill:1988:CEM

[645] Joe R. Hill and Chih-Ling Tsai. Calculating the efficiency of maximum quasilikelihood estimation. Applied Statistics, 37(2):219–230, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tang:1988:EDT

[646] Kwei Tang. Economic design of a twosided screening procedure using a correlated variable. Applied Statistics, 37 (2):231–241, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

DeStavola:1988:TDT

[647] Bianca L. De Stavola. Testing departures from time homogeneity in multistate Markov processes. Applied Statistics, 37(2):242–250, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Vinod:1988:EIB

[648] H. D. Vinod and Baldev Raj. Economic issues in bell system divestiture: A bootstrap application. Applied Statistics, 37 (2):251–261, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Thomas:1988:LE

[649] G. E. Thomas, M. J. Phillips, D. Strauss, B. C. Arnold, and John S. Fry. Letters to the editors. Applied Statistics, 37(2):262–266, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Courtenay:1988:BRY

[650] Gill Courtenay. Book reviews: Youth Unemployment: Social and Psychological Perspectives, by Michael H. Banks and Philip Ullah. Applied Statistics, 37 (2):267–268, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Trewsdale:1988:BRP

[651] Janet M. Trewsdale. Book reviews:

Part-Time Employment in Great Britain — An Analysis Using Establishment Data, by David Blanchflower and Bernard Corry. Applied Statistics, 37 (2):268–269, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hey:1988:BRB

[652] John D. Hey. Book reviews: Bayesian Analysis and Uncertainty in Economic Theory, by Richard M. Cyert and Morris H. Degroot. Applied Statistics, 37(2):269, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Matthews:1988:BRE

[653] J. N. S. Matthews. Book reviews: Essentials of Biostatistics, by Robert C. Elston and William D. Johnson. Applied Statistics, 37(2):269–270, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bissell:1988:BRS

[654] Derek Bissell. Book reviews: The Statistical Consultant in Action, by D. J. Hand and B. S. Everitt. Applied Statistics, 37(2):270–271, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hope:1988:BRS

[655] R. P. Hope. Book reviews: Statistics for Business and Economics, by Paul Newbold. Applied Statistics, 37(2):271, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Whittlestone:1988:BRS

[656] David Whittlestone. Book reviews: SPSS-X^{(R)} User’s Guide. Edition 2. Applied Statistics, 37(2):271–272, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bailey:1988:SSR

[657] Trevor Bailey and R. D. Stern. Statistical software reviews: Instat. Applied Statistics, 37(2):273–277, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Koning:1988:SAA

[658] A. J. Koning and R. J. M. M.

Does. Statistical algorithms: Algorithm AS 234: Approximating the percentage points of simple linear rank statistics with Cornish-Fisher expansions. Applied Statistics, 37(2):278– 284, June 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/234.

Royston:1988:SAA

[659] J. P. Royston and D. G. Altman. Statistical algorithms: Algorithm AS 235: Number tally. Applied Statistics, 37(2): 285–290, June 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/235.

Balmer:1988:SAA

[660] D. W. Balmer. Statistical algorithms:

Algorithm AS 236: Recursive enumeration of r × c tables for exact likelihood evaluation. Applied Statistics, 37 (2):290–301, June 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/236.

Mareschal:1988:SAA

[661] B. Mareschal and G. Mélard. Statistical algorithms: Algorithm AS 237: The corner method for identifying autoregressive moving average models. Applied Statistics, 37(2):301– 316, June 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/237.

Dusoir:1988:SAR

[662] Tony Dusoir. Statistical algorithms: Remark AS R73: A remark on Algorithm AS 222: Resistant smoothing using the fast Fourier transform. Applied Statistics, 37(2):316–317, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

                                        stat.cmu.edu/apstat/222. See [560].

Clarkson:1988:SARb

[663] Douglas B. Clarkson. Statistical algorithms: Remark AS R74: A least squares version of Algorithm AS 211: The F -G diagonalization algorithm. Applied Statistics, 37(2):317– 321, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/211. See [635, 456].

Anonymous:1988:BMPb

[664] Anonymous. Back matter. Applied Statistics, 37(2):[3], 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1988:FMPb

[665] Anonymous. Front matter. Applied Statistics, 37(2):[3], 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chambers:1988:DAR

[666] R. L. Chambers. Design-adjusted regression with selectivity bias. Applied Statistics, 37(3):323–334, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Greenland:1988:CFD

[667] Sander Greenland and Ruth M. Mickey.

Closed form and dually consistent methods for inference on strict collapsibility in 2 × 2 × K and 2 × J × K tables. Applied Statistics, 37(3):335–343, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mak:1988:AIC

[668] Tak K. Mak. Analysing intraclass correlation for dichotomous variables. Applied Statistics, 37(3):344–352, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Tyssedal:1988:AMS

[669] John S. Tyssedal and Dag Tjøstheim.

An autoregressive model with suddenly changing parameters and an application to stock market prices. Applied Statistics, 37(3):353–369, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Smith:1988:SDO

[670] Philip J. Smith. Survey design optimization for estimating the exploitable biomass of a fishery accounting for nonsampling errors. Applied Statistics, 37 (3):370–384, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Manly:1988:CSS

[671] Bryan F. J. Manly. The comparison and scaling of student assessment marks in several subjects. Applied Statistics, 37 (3):385–395, 1988. CODEN APSTAG.

ISSN 0035-9254 (print), 1467-9876 (electronic).

Sengupta:1988:SEP

[672] Sukalyan Sengupta. Statistical evaluation of pharmacopoeia weight variation tests for tablets using a ratio statistics. Applied Statistics, 37(3):396–400, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rosenbaum:1988:PTM

[673] Paul R. Rosenbaum. Permutation tests for matched pairs with adjustments for covariates. Applied Statistics, 37(3): 401–411, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Sikkel:1988:RTR

[674] Dirk Sikkel and Gerard Jelierse. Renewal theory and retrospective questions. Applied Statistics, 37(3):412–420, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1988:TDB

[675] D. A. Williams. Tests for differences between several small proportions. Applied Statistics, 37(3):421–434, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Dinse:1988:PAA

[676] Gregg E. Dinse. A prevalence analysis that adjusts for survival and tumour lethality. Applied Statistics, 37 (3):435–445, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ahmad:1988:BRI

[677] R. Ahmad. Book reviews: Introduction to Stochastic Differential Equations, by T. C. Gard. Applied Statistics, 37(3):446, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Walters:1988:BRD

[678] J. R. Walters. Book reviews: Decision Analysis, by G. Gregory. Applied Statistics, 37(3):446–447, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clarke:1988:BRR

[679] G. M. Clarke. Book reviews: Response Surfaces: Designs and Analyses, by A. I. Khuri and J. A. Cornell. Applied Statistics, 37(3):447–448, 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Bibby:1988:BRT

[680] John Bibby. Book reviews: Think and Explain with Statistics, by L. E. Moses. Applied Statistics, 37(3):448–449, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Morgan:1988:BRN

[681] Byron J. T. Morgan. Book reviews:

Non-Linear Parameter Estimation — An Integrated System in BASIC, by J. C. Nash and M. Walker-Smith. Applied Statistics, 37(3):449–450, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Usher:1988:BRI

[682] M. B. Usher. Book reviews: Introduction to Biostatistics, 2nd Edn, by R. R.

Sokal and F. J. Rohlf. Applied Statistics, 37(3):451, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Manning:1988:BRB

[683] D. N. Manning. Book reviews: Balance of Payments Adjustment 1945 to 1986: The IMF Experience, by Margaret G. de Vries. Applied Statistics, 37(3):451–452, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Arsham:1988:BRS

[684] H. Arsham. Book reviews: Statistical Data Analysis Handbook, by F. J. Wall. Applied Statistics, 37(3):452–454, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Clayton:1988:SSR

[685] David Clayton and Peter A. Lachenbruch. Statistical software reviews: SURVCALC, by P. A. Lachenbruch. Applied Statistics, 37(3):455–456, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Farebrother:1988:SAA

[686] R. W. Farebrother. Statistical algorithms: Algorithm AS 238: A simple recursive procedure for the L_{1} norm fitting of a straight line. Applied Statistics, 37(3):457–465, September 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/238.

Shea:1988:SAA

[687] B. L. Shea. Statistical algorithms: Algorithm AS 239: Chi-squared and incomplete gamma integral. Applied Statistics, 37(3):466–473, September 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/239.

Hui:1988:SAA

[688] Yer Van Hui. Statistical algorithms: Algorithm AS 240: Updating the inverse of the dispersion matrix. Applied Statistics, 37(3):474–477, September 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://

lib.stat.cmu.edu/apstat/240.

Wichura:1988:SAA

[689] Michael J. Wichura. Statistical algorithms: Algorithm AS 241: The percentage points of the normal distribution. Applied Statistics, 37(3):477–484, September 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/241.

Sallas:1988:SAR

[690] William M. Sallas. Statistical algorithms: Remark AS R75: Some remarks on Algorithm AS 164: Least squares subject to linear constraints. Applied Statistics, 37(3):484–489, 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/164. See

[163].

Riddle:1988:CPM

[691] Lawrence H. Riddle. Corrigendum:

Probability models for tennis scoring systems. Applied Statistics, 37(3):490, 1988. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See

Anonymous:1988:BMPc

[692] Anonymous. Back matter. Applied Statistics, 37(3):[5], 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1988:FMPc

[693] Anonymous. Front matter. Applied Statistics, 37(3):[3], 1988. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1988:VIP

[694] Anonymous. Volume information. Applied Statistics, 37(3):[3], 1988. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Haslett:1989:STM

[695] John Haslett and Adrian E. Raftery. Space-time modelling with long-memory dependence: Assessing Ireland’s wind power resource. Applied Statistics, 38 (1):1–50, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pettitt:1989:CWM

[696] A. N. Pettitt and I. Bin Daud. Caseweighted measures of influence for proportional hazards regression. Applied Statistics, 38(1):51–67, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lin:1989:ANT

[697] Jinn Tyan Lin. Approximating the normal tail probability and its inverse for use on a pocket calculator. Applied Statistics, 38(1):69–70, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Copas:1989:USS

[698] J. B. Copas. Unweighted sum of squares test for proportions. Applied Statistics, 38(1):71–80, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Archer:1989:RCD

[699] Linda E. Archer and Louise M. Ryan.

On the role of cause-of-death data in the analysis of rodent tumorigenicity experiments. Applied Statistics, 38(1):81–93, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simmons:1989:PMS

[700] John Simmons. A probabilistic model of squash: strategies and applications. Applied Statistics, 38(1):95–110, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Kahn:1989:NRP

[701] Barbara E. Kahn and Donald G. Morrison. A note on ‘random’ purchasing: Additional insights from dunn, reader and wrigley. Applied Statistics, 38(1):111–114, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Simonoff:1989:UGR

[702] Jeffrey S. Simonoff and Chih-Ling Tsai.

The use of guided reformulations when collinearities are present in nonlinear regression. Applied Statistics, 38(1): 115–126, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hamerle:1989:MSR

[703] Alfred Hamerle. Multiple-spell regression models for duration data. Applied Statistics, 38(1):127–138, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jolliffe:1989:RID

[704] I. T. Jolliffe. Rotation of ill-defined principal components. Applied Statistics, 38 (1):139–147, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Williams:1989:CRC

[705] E. R. Williams and J. A. John. Construction of row and column designs with contiguous replicates. Applied Statistics, 38(1):149–154, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Elliot:1989:SSR

[706] David Elliot and M. Volk. Statistical software reviews: Clusters: Version 3, by Vijay Verma and Mick Pearce. Applied Statistics, 38(1):155–157, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hunt:1989:SSR

[707] Richard Hunt and J. Sookias. Statistical software reviews: Statgraphics: Version 2.6, by STSC Inc. Applied Statistics, 38 (1):158–160, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shea:1989:SAA

[708] B. L. Shea. Statistical algorithms: Algorithm AS 242: The exact likelihood of a vector autoregressive moving average model. Applied Statistics, 38(1):161– 184, March 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/242.

Lenth:1989:SAA

[709] Russell V. Lenth. Statistical algorithms:

Algorithm AS 243: Cumulative distribution function of the non-central t distribution. Applied Statistics, 38(1):185– 189, March 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/243.

Geng:1989:SAA

[710] Zhi Geng. Statistical algorithms:

Algorithm AS 244: Decomposability and collapsibility for log-linear models. Applied Statistics, 38(1):189– 197, March 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/244.

Macleod:1989:SAR

[711] Allan J. Macleod. Statistical algorithms:

Remark AS R76: A remark on Algorithm AS 215: Maximum-likelihood estimation of the parameters of the generalized extreme-value distribution. Applied Statistics, 38(1):198–199, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http:/

/lib.stat.cmu.edu/apstat/215. See

[481].

Shea:1989:SAR

[712] B. L. Shea. Statistical algorithms: Remark AS R77: A remark on Algorithm AS 152: Cumulative hypergeometric probabilities. Applied Statistics, 38(1):199–204, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/152. See [88, 268].

Farebrother:1989:SAC

[713] R. W. Farebrother. Statistical algorithms: Correction: Correction to Algorithm AS 231: The distribution of a noncentral χ^{2} variable with nonnegative degrees of freedom. Applied Statistics, 38(1):204, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/231.

Anonymous:1989:BMPa

[714] Anonymous. Back matter. Applied Statistics, 38(1):[4], 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1989:FMPa

[715] Anonymous. Front matter. Applied Statistics, 38(1):[3], 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ansell:1989:PPS

[716] J. I. Ansell and M. J. Phillips. Practical problems in the statistical analysis of reliability data. Applied Statistics, 38 (2):205–247, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). With discussion and a reply by the authors.

vanderHeijden:1989:CAC

[717] Peter G. M. van der Heijden, Antoine de Falguerolles, and Jan de Leeuw. A combined approach to contingency table analysis using correspondence analysis and log-linear analysis. Applied Statistics, 38(2):249–292, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). With discussion and a reply by the authors.

Hunt:1989:IMV

[718] D. N. Hunt and C. M. Triggs. Iterative missing value estimation. Applied Statistics, 38(2):293–300, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Osanaiye:1989:ECS

[719] P. A. Osanaiye. An economic choice of sampling inspection plans under varying process quality. Applied Statistics, 38 (2):301–308, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Zeger:1989:PRS

[720] Scott L. Zeger and Sharon L. Edelstein.

Poisson regression with a surrogate X; an analysis of vitamin A and Indonesian children’s mortality. Applied Statistics, 38(2):309–318, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Keiding:1989:NET

[721] Niels Keiding and Per Kragh Andersen.

Nonparametric estimation of transition intensities and transition probabilities: a case study of a two-state Markov process. Applied Statistics, 38(2):319–329, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Makridakis:1989:SDP

[722] Spyros Makridakis and Robert L. Winkler. Sampling distributions of postsample forecasting errors. Applied Statistics, 38(2):331–342, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Minkin:1989:FAI

[723] Salomon Minkin. Fit assessment and identification of functional form in logistic regression. Applied Statistics, 38 (2):343–350, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Maclaren:1989:GMI

[724] N. M. Maclaren. The generation of multiple independent sequences of pseudorandom numbers. Applied Statistics, 38 (2):351–359, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Cullis:1989:NPA

[725] Brian R. Cullis, Warwick J. Lill, John A. Fisher, Barbara J. Read, and Alan C. Gleeson. A new procedure for the analysis of early generation variety trials. Applied Statistics, 38(2):361–375, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jackson:1989:LE

[726] David A. Jackson and Lawrence Riddle. Letter to the editors. Applied Statistics, 38(2):377–379, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1989:BRP

[727] Anonymous. Book reviews: Properties of Estimators for the Gamma Distribution, by K. O. Bowman and L. R. Shenton. Applied Statistics, 38(2):381, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Fryer:1989:BRL

[728] M. J. Fryer. Book reviews: Lognormal Distributions: Theory and Applications, by E. L. Crow and K. Shimizu. Applied Statistics, 38(2):381–382, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Rao:1989:BRT

[729] T. Subba Rao. Book reviews: Tensor Methods in Statistics, by P. McCullagh. Applied Statistics, 38(2):382–383, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Weatherup:1989:BRM

[730] S. T. C. Weatherup. Book reviews:

Manual of Crop Experimentation, by S. C. Pearce, G. M. Clarke, G. V. Dyke and R. E. Kempson. Applied Statistics, 38 (2):383–384, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hand:1989:BRM

[731] D. J. Hand. Book reviews: Mixture Models: Inference and Applications to Clustering, by G. J. McLachlan and K. E. Basford. Applied Statistics, 38 (2):384–385, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Stirzaker:1989:BRF

[732] D. Stirzaker. Book reviews: Foundations of Statistics, by D. G. Rees. Applied Statistics, 38(2):385, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Harding:1989:BRE

[733] E. F. Harding. Book reviews: Exploiting Continuity: Maximum Entropy Estimation of Continuous Distributions, by H. Theil and D. G. Fiebig. Applied Statistics, 38(2):386–387, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chalmers:1989:BRM

[734] C. P. Chalmers. Book reviews: Multivariate Exploratory Data Analysis: A Perspective on Exploratory Factor Analysis, by A. Yates. Applied Statistics, 38 (2):387–388, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mellor:1989:SSR

[735] Chris Mellor and David P. Reilly. Statistical software reviews: AUTOBOX-PLUS: Version 2.0, by Automatic Forecasting Systems Inc. Applied Statistics, 38(2):389–392, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jones:1989:SSR

[736] Byron Jones. Statistical software reviews: X-STAT, by Softpower Incorporated. Applied Statistics, 38(2):392–395, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Macleod:1989:SAA

[737] Allan J. Macleod. Statistical algorithms: Algorithm AS 245: A robust and reliable algorithm for the logarithm of the gamma function. Applied Statistics, 38 (2):397–402, June 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/245.

Azzalini:1989:SAA

[738] A. Azzalini. Statistical algorithms: Algorithm AS 246: An analysis of variance table for repeated measurements with unknown autoregressive parameter. Applied Statistics, 38(2):402– 411, June 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/246.

Murphy:1989:SAA

[739] B. P. Murphy and J. S. Rohl. Statistical algorithms: Algorithm AS 247: Updating the sufficient configurations for fitting ZPA models to multidimensional contingency tables. Applied Statistics, 38(2):412–420, June 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/247.

Ridout:1989:SAR

[740] Martin S. Ridout and John M. Cobby.

Statistical algorithms: Remark AS R78: A remark on Algorithm AS 178: The Gauss–Jordan sweep operator with detection of collinearity. Applied Statistics, 38(2):420–422, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/178. See [236].

Cran:1989:SAR

[741] G. W. Cran. Statistical algorithms: Remark AS R79: Comment on: A remark on algorithms AS 32 and AS 147: The incomplete gamma integral. Applied Statistics, 38(2):423, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). See [7, 94, 63].

Anonymous:1989:BMPb

[742] Anonymous. Back matter. Applied Statistics, 38(2):[4], 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1989:FMPb

[743] Anonymous. Front matter. Applied Statistics, 38(2):[3], 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lesaffre:1989:MGL

[744] E. Lesaffre and A. Albert. Multiplegroup logistic regression diagnostics. Applied Statistics, 38(3):425–440, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Wilson:1989:CST

[745] Jeffrey R. Wilson. Chi-square tests for overdispersion with multiparameter estimates. Applied Statistics, 38(3):441–453, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Ganesalingam:1989:CMA

[746] S. Ganesalingam. Classification and mixture approaches to clustering via maximum likelihood. Applied Statistics, 38(3):455–466, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Shoukri:1989:URM

[747] M. M. Shoukri and R. H. Ward. Use of regression models to estimate genetic parameters and measures of familial resemblance in man. Applied Statistics, 38 (3):467–479, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Trajstman:1989:ICD

[748] A. C. Trajstman. Indices for comparing decontaminants when data come from dose-response survival and contamination experiments. Applied Statistics, 38 (3):481–494, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hinkley:1989:MPL

[749] D. V. Hinkley. Modified profile likelihood in transformed linear models. Applied Statistics, 38(3):495–506, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Freeman:1989:TLD

[750] G. H. Freeman. The tactics of liar dice. Applied Statistics, 38(3):507–516, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Richardson:1989:MOP

[751] S. C. Richardson. Miscellanea: One-pass selection of a sample with probability proportional to size. Applied Statistics, 38(3):517–520, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Jolliffe:1989:BRL

[752] Ian Jolliffe. Book reviews: Latent Variable Models and Factor Analysis, by D. J. Bartholomew. Applied Statistics, 38(3):521, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Goodwin:1989:BRA

[753] Paul Goodwin. Book reviews: Applied Statistics: A First Course, by M. L.

Berenson, D. M. Levine and D. Rindskopf. Applied Statistics, 38(3):522, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1989:BRM

[754] Anonymous. Book reviews: Multivariate Statistical Modeling and Data Analysis, by H. Bozdogan and A. K. Gupta. Applied Statistics, 38(3):522–523, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Preece:1989:BRT

[755] D. A. Preece. Book reviews: Theory of Block Designs, by A. Dey. Applied Statistics, 38(3):523–524, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Prescott:1989:BRC

[756] R. J. Prescott. Book reviews: Clinimetrics, by A. R. Feinstein. Applied Statistics, 38(3):524–525, 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Holt:1989:BRM

[757] D. Holt. Book reviews: Multilevel Models in Educational and Social Research, by H. Goldstein. Applied Statistics, 38(3):525, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mosurski:1989:BRE

[758] K. R. Mosurski. Book reviews: Encyclopedia of Statistical Sciences, Vol. 9: Strata Chart to Zyskind-Martin Models, by S. Kotz and N. L. Johnson. Applied Statistics, 38(3):525–526, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Coe:1989:BRR

[759] Richard Coe. Book reviews: Reviews of United Kingdom Statistical Sources: Vol. XVII, Weather and Water, by W. F. Maunder and M. C. Fleming. Applied Statistics, 38(3):526–527, 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Waterton:1989:BRR

[760] Jennifer Waterton. Book reviews:

Reviews of United Kingdom Statistical Sources: Vol. XVIII, Posts and Telecommunications, by W. F. Maunder and M. C. Fleming. Applied Statistics, 38(3):527–528, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Mitchell:1989:BRS

[761] J. M. O. Mitchell. Book reviews:

STATLAB: Microcomputer-based practical classes and demonstrations for teaching statistical concepts, by W. D. Stirling. Applied Statistics, 38(3):528, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Chanter:1989:SSR

[762] Dennis O. Chanter and Volker W.

Rahlfs. Statistical software reviews: N: Version 1.2. Applied Statistics, 38 (3):529–531, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Plewis:1989:SSRa

[763] Ian Plewis, W. A. Fuller, W. Kennedy, D. Schnell, G. Sullivan, H. J. Park, and Richard Coe. Statistical software reviews: PC CARP, by W. A. Fuller, W. Kennedy, D. Schnell, G. Sullivan and H. J. Park. Applied Statistics, 38 (3):531–534, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Plewis:1989:SSRb

[764] Ian Plewis, W. A. Fuller, W. Kennedy, D. Schnell, G. Sullivan, H. J. Park, and Richard Coe. Statistical software reviews: EV CARP, by W. A. Fuller, D. Schnell, W. Kennedy, D. Schnell, G. Sullivan and H. J. Park. Applied Statistics, 38(3):531–534, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Davis:1989:SAA

[765] Charles S. Davis and Michael A. Stephens. Statistical algorithms: Algorithm AS 248: Empirical distribution function goodness-of-fit tests. Applied Statistics, 38(3):535–543, September 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/248.

Leppard:1989:SAA

[766] P. Leppard and G. M. Tallis. Statistical algorithms: Algorithm AS 249: Evaluation of the mean and covariance of the truncated multinormal distribution. Applied Statistics, 38(3):543–553, September 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/249.

Parhizgari:1989:SAA

[767] Ali M. Parhizgari and Arun J. Prakash. Statistical algorithms: Algorithm AS 250: Tests of the equality of dispersion matrices. Applied Statistics, 38(3):

553–564, September 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL http://lib.

stat.cmu.edu/apstat/250.

Dunnett:1989:SAA

[768] Charles W. Dunnett. Statistical algorithms: Algorithm AS 251: Multivariate normal probability integrals with product correlation structure. Applied Statistics, 38(3):564–579, September 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/251.

Boys:1989:SAA

[769] Richard J. Boys. Statistical algorithms:

Algorithm AS R80: A remark on Algorithm AS 76: An integral useful in calculating noncentral t and bivariate normal probabilities. Applied Statistics, 38 (3):580–582, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.

edu/apstat/76. See [10, 27, 31, 418, 436, 542].

Dewey:1989:SAR

[770] Michael E. Dewey. Statistical algorithms: Remark AS R81: A remark on Algorithm AS 131: Tabulating frequency distributions for variables with structured code sets. Applied Statistics, 38(3):582, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

tronic). URL http://lib.stat.cmu.
edu/apstat/131. See [25].

Anonymous:1989:BMPc

[771] Anonymous. Back matter. Applied Statistics, 38(3):[7], 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1989:FMPc

[772] Anonymous. Front matter. Applied Statistics, 38(3):[3], 1989. CODEN AP-STAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Anonymous:1989:VIP

[773] Anonymous. Volume information. Applied Statistics, 38(3):[3], 1989. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Frick:1990:SAA

[774] H. Frick. Statistical algorithms: Algorithm AS R84: A remark on Algorithm AS 226: Computing non-central beta probabilities. Applied Statistics, 39 (2):311–312, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/226. See [583, 775].

Lam:1995:SAR

[775] Ming Long Lam. Statistical algorithms: Remark AS R95: A remark on Algorithm AS 226: Computing non-central beta probabilities. Applied Statistics, 44 (4):551–552, 1995. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (elec-

                                        tronic). URL http://lib.stat.cmu.
                                        edu/apstat/226. See [583, 774].